A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7450–7499

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

(1, λ) − σ [744]. 2.5 [123]. 2 × 2 [283]. 3 [459, 476, 251, 339, 256, 441, 455, 474]. 4 [754]. ^{∗} [51]. _{P B} [211]. C^{2} [740]. EL^{++} [145]. χ^{2} [644]. ϵ [769]. μ [886, 659]. N [243, 754]. O(1) [745]. ± [874, 812]. Σ_{3}^{0} [101].

-Calculus [886, 659]. -composite [740]. -Gram [243]. -MOEA [769]. -tuples [754].

0/1 [770]. 0x [393].

1 [634]. 1.1 [665]. 10th [954, 968]. 12th [944, 933, 964, 965, 932]. 13th [959]. 15th [937, 949, 971, 956, 930, 972]. 160 [585]. 17th [934]. 18th [966, 957, 953]. 19th [935, 931, 963].

2 [666]. 2.0 [579, 705]. 2008 [503]. 2009 [542]. 20th [941]. 23rd [929].

3 [313, 719, 715, 239]. 34th [950]. 36th [950]. 37th [939]. 3GPP [385].

5PM [638]. 5th [944, 925, 951].

6DOF [475]. 6th [970, 969].

7th [943, 952].

8.9/TC [940]. 802.11s [528]. 802.16 [542]. 802.16-2009 [542]. 802.16m [376]. 8th [936, 960, 958].

9th [955, 942].

Aardvarks [295]. ABC [206]. ABC-Miner [206]. Abduction [880]. ABE [845]. Abilities [264]. Ability [257]. ABS [845]. Abstract [537, 192, 194, 203, 191, 244, 245]. Abstraction [196, 68, 407, 83, 193]. Abstractions [78]. Abuses [179]. Academia [967]. Academic [695]. Acceleration [752]. Acceptance [36]. Access [231, 877, 961, 224, 164, 371, 853, 299, 784, 16, 731, 540, 574]. Accessibility [325, 520]. Accessibility-Based [325]. Accidents [157]. Achieve [85, 151]. Achievements [449]. ACO [208, 747, 774]. AcoSeeD [223]. Acquiring [354]. Acquisition [130, 923, 124, 127, 266, 12, 932]. ACRI [968]. across [416]. Action [658]. ActionScript [239]. Active [708]. Actively [639]. Actively/Covertly [639]. Activities [790, 793, 357]. Activity [138, 260]. Acyclic [826]. Ad [517, 233]. Ad-Hoc [517]. Adapt [258, 764, 407]. Adaptation [743, 314, 322, 327, 329, 403]. Adapting [317, 321]. Adaption [502]. Adaptive [383, 615, 529, 544, 94, 721, 218, 221]. Addition [489]. Additive [248]. Additively [598]. Address [161]. Addressing [126]. Adjacency [524]. Adjoining [102]. Adjunctions [391]. Adjustment [448, 384]. Administration [47]. Administrative [42, 34]. Admissible [103]. Admission [387]. Adoption [92]. Advanced [383, 960, 719, 245, 568]. Advances [286, 882, 963]. Advancing [927]. Adversaries [641]. Advertising [545]. AEGS [946]. AES [639]. Affect [366]. Affecting [31]. Affective [408]. affine [450]. against [231, 928, 174, 243, 238, 912, 426]. Agencies [580]. Agency [400]. Agendas [562]. Agent [137, 401, 404, 652, 959, 399, 92, 152, 649, 930, 397, 797, 89, 138, 83, 139, 648, 658]. Agent-Oriented [401]. Agents [650, 406, 398, 946]. Aggregate [632, 762]. Aggregation [56, 386, 143]. Agile [768]. Ahead [247]. AHNE [362]. AI [143]. Aided [312, 184]. Air [57]. Aires [931]. Airport [447]. AK [233]. AK-PPM [233]. Alarm [371]. ALERT [240]. ALERT-ID [240]. Algebra [890]. Algebraic [62]. Algorithm [428, 277, 89, 775, 865, 148, 337, 422, 872, 205, 766, 220, 565, 211, 620, 908, 742, 430, 771, 206]. Algorithm-Dependent [742]. Algorithmic [117, 761, 830]. Algorithmics [269]. Algorithms [208, 748, 677, 457, 764, 210, 741, 739, 193, 490, 538]. Aligned [914].

Aligning [372, 554]. Allocation [132, 725, 86, 85]. Alternating [899]. Alternating-Time [899]. Alternations [922]. Alternatives [303]. Always [270]. Amalfi [958]. Ambient [651]. Ambiguity [331, 913]. among [795, 580]. Amortized [628]. Amplifier [439]. Amsterdam [937]. Analogies [757]. Analogy [259]. Analyses [186]. Analysing [207]. Analysis [561, 296, 58, 748, 197, 675, 7, 790, 949, 181, 167, 224, 311, 188, 243, 92, 677, 189, 18, 102, 202, 384, 457, 759, 22, 127, 234, 187, 564, 570, 515, 232, 813, 935, 446, 445, 826, 728, 528, 306, 907, 885, 452, 741, 402, 746, 67, 429, 796]. Analytics [961]. Analyze [240, 503]. Android [165, 163]. Anisotropic [471]. Annotated [919]. Annotation [150, 918, 466, 914]. Announcement [82]. ANNPR [951]. Anomalous [414]. Anomaly [594]. Anonymous [629, 850, 843]. Answer [871, 658]. Answering [960, 869, 131, 663]. Answers [874]. Ant [223, 224, 785, 206, 217]. Ant-Based [224, 206]. Antenna [529]. Antennas [383]. Anticipatory [96]. ANTS [936, 221]. Anytime [464]. Apache [785]. Apartments [354]. Apiary [773]. Appearance [683]. APPification [303]. Appliances [364]. Application [132, 294, 720, 659, 302, 610, 683, 776, 97]. Applications [313, 806, 671, 850, 165, 971, 1, 380, 371, 948, 614, 103, 904, 4, 898, 6, 558, 731]. Applying [709]. Appraisal [315]. Approach [140, 111, 887, 656, 633, 602, 222, 557, 660, 713, 41, 179, 870, 18, 36, 610, 922, 291, 564, 23, 327, 813, 328, 655, 516, 831, 350, 520, 860, 624, 382, 498, 237, 889]. Approaches [671, 530, 19, 390, 93]. Approbation [45]. Appropriation [95]. Approximate [279, 125, 587, 739]. Approximation [282, 249, 767]. Approximations [618]. Apps [302]. Apron [447]. Arabia [31]. Arborescence [210]. Architecting [550]. Architecture [408, 553, 541, 724]. Architectures [547, 733]. Archives [697]. Archiving [764]. Area [613]. Areas [289]. ARES [940]. Argentina [931]. Argument [912]. Argumentation [878, 647, 660, 659, 655, 672, 21]. Argumentation-Based [655]. Arithmetic [105]. armed [214]. Arrays [824]. Arrows [110]. Art [711]. Articles [680]. Articulated [450]. Artifacts [562, 366]. Artificial [933, 500, 321, 291, 142, 495, 771, 951]. Artistic [355]. ASK [615]. ASLan [170]. ASP [879]. ASPDA [879]. Assembly [217]. Assess [521]. Assessing [263, 228]. Assessment [447, 881, 57, 461, 522]. Assessments [688]. Assignment [179, 909]. Assistance [266, 444]. Assisted [227, 256, 776]. Associated [494]. Associative [708]. Assumption [672]. Assumption-Based [672]. Assumptions [630, 856, 628]. Async [187]. Async-Finish [187]. Asynchronous [76, 104]. Athens [947]. ATL [653]. Atl* [75]. Atomic [727]. Attack [165, 176, 427]. Attackers [171]. Attacks [231, 937, 174, 643, 585, 229, 237]. Attitudes [649]. Attribute [629, 323, 853, 911]. Attribute-Based [853]. Attributes [140]. Attribution [233]. Auction [88]. Auction-Based [88]. Audio [364, 439, 699, 943, 501, 353, 358, 365, 362, 354]. Audio-Haptic [364, 353, 358]. Audio-Tactile [362]. Auditory [352]. Augmenting [361].

August [944, 959, 955, 957, 943, 950, 940, 939, 938, 953]. Austria [928, 967, 926, 960, 925, 927, 970, 950, 963]. Austrian [37]. Authenticated [596, 858, 233, 857]. Authentication [535, 584, 426, 588]. Authorization [169, 16]. Autocorrelation [516]. Automata [78, 820, 69, 66, 821, 71, 829, 823, 825, 831, 832, 836, 833, 968, 838]. Automata-Based [831]. Automated [508, 111, 885, 482]. Automatic [196, 561, 123, 208, 124, 918, 708, 402]. Automation [534]. Automaton [839]. autonomous [253]. Autonomy [324]. Auxiliary [626]. Auxiliary-Input [626]. Availability [940]. Avatars [402]. Avoiding [69]. Aware [405, 173, 290, 592, 510, 866, 50, 654, 913, 555]. Awareness [373, 343, 634]. Axis [261].

B [887, 889, 481]. B-Spline [481]. Back [10, 27, 59, 98, 367, 395, 409, 420, 462, 505, 582, 661, 673, 692, 798, 818, 861, 900]. Backward [200, 872]. Bad [108]. Balance [291]. Balancing [400, 617]. Band [531]. Bandit [214]. Bare [209]. Barrier [727]. Based [312, 307, 537, 310, 750, 563, 140, 68, 313, 592, 877, 289, 383, 866, 408, 890, 534, 850, 125, 458, 51, 315, 447, 176, 345, 224, 599, 147, 470, 342, 765, 86, 311, 535, 320, 148, 417, 829, 427, 677, 138, 791, 321, 300, 643, 881, 610, 322, 155, 323, 158, 325, 326, 133, 922, 169, 202, 552, 483, 578, 424, 53, 327, 328, 501, 655, 769, 802, 831, 122, 853, 753, 299, 363, 88, 430, 542, 206, 16, 451, 444, 844, 468, 539, 672, 558, 97, 751, 624, 524, 626, 724, 854, 94, 3, 851, 348, 435, 431]. Based [797, 889, 941]. Baseline [455]. Bases [863, 190, 657]. Battery [50]. Battery-Aware [50]. Bayesian [206]. Be [295, 711]. beam [383]. Beanplots [689]. Beats [747]. Bee [771]. Been [373, 707]. Behavior [835, 399, 424, 695, 773, 834, 520, 429]. Behavioral [580]. Behaviors [83, 443, 595]. Behaviour [744, 408, 414]. Behaviours [318]. Belgium [936]. Belief [112]. Benchmark [730, 729]. Benchmarking [677]. BETA [889]. Better [407, 688]. Between [143, 58, 57, 257, 524, 347]. Beyond [864, 715, 678]. BGV [627]. BGV-Style [627]. Bi [208, 114]. Bi-implication [114]. Bi-objective [208]. Bias [753]. Bidirectional [390]. Big [824]. Bigraph [523]. Bigtable [3]. Bigtable/MapReduce [3]. Bigtable/MapReduce-Based [3]. Bilateral [193]. Bilinear [841, 857]. Bilingual [914]. BIM [335]. Binary [243, 832]. Bio [926, 207]. Bio- [926]. Bio-PEPA [207]. Biobjective [748]. Bioimpedance [26]. Biological [223, 21]. Biology [142]. Bipolar [647]. Bisimulation [659]. Bit [846]. Bit-Commitments [846]. Blackbox [634]. Blends [291]. Blind [631, 600, 154]. Block [489]. Blocking [734]. Blocks [592]. Blood [17]. Bloom [855]. BOA [753]. Bones [209]. Books [699]. Boolean [747]. Boosting [166]. Bootstrapping [762]. Boston [947]. Botnets [245]. Bottom [93]. Bottom-Up [93]. Boundary [742, 915]. Bounded [277, 278, 91, 650]. Bounds [108, 587]. BPM [954, 550, 559]. Bratislava [939]. Brazil [971]. Brazilian [971]. Breach [426]. Bringing [677]. Brno [972]. Broad [807]. Broadcast [635]. Browsers [589]. Brussels [936].

Budapest [953]. Buenos [931]. Building [534, 335, 839]. Buildings [52, 347]. Built [520]. Bundle [448]. Burst [530]. Business [561, 563, 553, 878, 40, 302, 370, 564, 303, 814, 568, 549, 954].

C [393]. CA [822, 826]. CAD [496]. Calculus [886, 197, 659, 219, 899]. Call [115]. Call-by-Value [115]. Calls [550]. Camera [458, 452, 474]. Camera-Based [458]. Can [295, 133]. Cancer [20, 910]. Candidate [160]. cAnt [211]. cAnt-Miner [211]. Capabilities [552, 859]. Capability [539]. Capacities [912]. Capital [48]. Capture [565, 456]. Card [590, 426, 37]. Cardinality [160]. Cardiomyopathy [23]. Cardiovascular [25]. Care [17]. Cartesian [752]. Cartograms [509]. Cartoon [123]. Case [312, 278, 537, 310, 313, 780, 845, 32, 315, 531, 130, 311, 320, 321, 323, 326, 552, 53, 327, 328, 785, 742, 941]. Case-Based [312, 310, 313, 315, 311, 320, 321, 323, 327, 328, 941]. Cased [348]. Cased-Based [348]. Cases [317]. Cash [640]. Cast [406]. Castle [949]. Catalonia [531]. Categorical [896]. Categorization [249, 130, 248]. Caused [255]. Cavity [431]. CBR [322]. CBR-Based [322]. CCA [634]. CCBR [325]. CD [940]. CD-ARES [940]. CDN [8, 55]. CDVE [942]. Cellar [823]. Cellular [820, 821, 829, 388, 825, 831, 832, 836, 833, 968, 839, 968, 838]. Censored [135]. Center [580]. Centerline [508]. Centralised [545]. Centrality [640]. Centred [700]. Centric [125]. Certificate [433]. Cesáro [158]. Chain [46]. Chaining [872]. Challenges [449, 126, 715, 794, 676]. Change [112, 57]. Changes [255]. Channel [179, 245, 530]. Channels [792]. Character [494, 855]. Characteristics [6, 229]. Characterization [886, 615, 488]. Characterizations [103]. Characterizing [785, 650]. Characters [406]. Charging [541]. Checking [886, 110, 888, 74, 83, 300, 601, 554, 555]. Chengde [948]. Chengdu [969]. Children [137, 260]. China [948, 969]. Choice [143, 815]. Choreographies [77]. Chronic [13]. Chunking [920]. CIGA [398]. Circuits [637]. Citation [147]. Cities [881]. City [349, 451, 574]. City-Scale [451]. Civil [449]. Class [807, 827, 453, 439]. Classes [300, 742]. Classification [477, 781, 325, 488, 211, 832, 206, 708, 751, 454, 482, 498]. Classifier [502, 153]. Classifiers [133]. CLEF [961, 678]. CLEF-IP [678]. Click [234]. Client [586]. CLIMA [959]. Climate [57]. Clocks [69]. Cloned [165]. Clones [165]. Closed [345]. Closures [199]. Cloud [563, 581, 925, 677, 2, 578, 608, 4, 226, 580, 574, 3]. Clouds [562]. Cluster [417]. Clustered [475, 592]. Clustering [835, 373, 15, 148, 760, 18]. Clusters [730, 654, 624, 623]. Clutter [248]. Co [43, 502, 413, 899]. Co-Creation [43]. Co-Creative [43]. Co-evolution [413]. Co-training [502]. Coalitional [82]. Code [592, 885]. Codes [388, 860]. Coevolution [752]. Coevolutionary [831]. Cognition [938]. Cognitive [397, 522]. coherent [733]. Cold [161]. Collaborate [46]. Collaborating [258, 342]. Collaboration [795, 596, 35, 786, 517, 966]. Collaboration-Preserving [596]. Collaborative [778, 790, 533, 723, 346, 779, 332, 793, 138, 791, 47, 794, 711, 50, 350, 340, 577, 797]. Collection [691, 402, 524]. Collections [680, 695, 701]. Collective [349, 219, 789]. Collectives [609, 726]. College [432]. Colony [223, 771, 217]. Color [480, 839, 435]. Columbus [952]. Combined [845]. Combining [507, 312, 294, 18, 918, 93]. Come [298]. Comfort [51]. Commerce [301]. Commitments [846]. Common [277, 600, 95, 921]. Common-Pool [95]. Communication [722, 628, 723, 732, 621, 882, 728, 731, 717, 953, 347, 721]. Communications [342, 724]. Communities [412, 783, 92, 773]. Community [411]. commutative [645]. Compact [631]. Comparative [314]. Compare [399, 485]. Comparing [689, 365]. Comparison [72, 385, 156, 363, 540, 642]. Competition [437]. Competition-Predator [437]. Compiler [718]. Complementary [390]. Complete [101, 872]. Completeness [111]. Complex [780, 319, 365, 429]. Complexity [282, 287, 268, 297, 746]. Compliance [556, 601, 555]. Component [201, 683]. Components [689]. Composable [641, 34]. Composed [575, 56]. composite [740]. Composited [823]. Composition [896]. Compositionality [892]. Comprehensive [218]. Compression [504]. Compromised [237]. COMPSs [608]. Computable [101]. Computation [594, 637, 4, 931, 275]. Computational [767, 24, 921, 746, 959]. Computer [312, 184, 778, 949, 339, 934, 939, 717, 724, 934]. Computer-Aided [184]. Computer-Mediated [339]. Computer-Supported [778]. Computing [277, 948, 612]. Concept [145, 18, 769, 299]. Concept-Based [769]. Concepts [393]. Conceptual [682]. Conceptualization [522]. Concrete [640]. CONCUR [929, 929]. Concurrency [791, 929, 62, 64]. Concurrent [65, 67]. Conditional [118]. Conditionally [843]. Conditions [167]. Conference [941, 944, 933, 412, 928, 954, 926, 961, 964, 965, 936, 956, 925, 966, 957, 927, 929, 970, 948, 942, 943, 940, 930, 968, 972, 938, 953, 969, 958, 952, 962]. Conferences [549]. Confidence [835, 326, 329]. Configuration [539]. Configurations [384]. Conflict [779, 348]. Confliction [827]. Confliction-Like [827]. Confluence [813]. Conformance [300, 554]. Congestion [375, 156, 96, 543]. Congruences [73]. Conjecture [805]. Connect [754]. Connect- [754]. Connectivity [592]. Conquer [148]. Conservation [408]. Consistency [655]. Consistent [874]. Constrained [744, 601, 591]. Constraint [276, 696]. Constraints [121, 153, 274]. Constructing [837]. Construction [634]. Constructions [899]. Consumption [612, 55]. Contacts [412]. Containment [68]. Content [705, 158, 8, 413]. Content-Based [158]. Context [307, 290, 373, 510, 725, 464, 657, 881, 159, 794, 532, 382, 454, 555, 863]. Context-Aware [290, 510, 555]. Contexts [404, 653, 649]. Contextual [796]. Continual [849]. Continual-Leakage [849]. Continuous [828, 740, 469, 817]. Continuous-State [828]. Contours [321]. Contracting [119]. Contraction [875]. Contrary [647]. Control [722, 197, 534, 375, 51, 224, 387, 791, 156, 96, 164, 372, 50, 376, 299, 16, 543, 574]. Controlling [772, 839]. Controversial [234]. Convergence [740, 736]. Convertible [842]. Convex [512, 736]. Convolution [158]. Cooling [50]. Cooperating [782]. Cooperation [339, 784]. Cooperative [338, 334, 222, 349, 337, 213, 388, 788, 336, 348, 942]. Coordination [258, 257]. Core [724, 53, 624]. Coreference [904]. Corpora [914]. Corpus [903, 919, 907, 918]. Correctness [553]. Correlation [371]. Correlogram [516]. Corruption [593]. Cost [535, 91, 4, 577]. Countable [117]. Counter [284, 83]. Counting [821, 739]. Coupled [472, 543]. Cover [737]. CQI [386]. CRAW [616]. CRAW/P [616]. Cray [727]. Creating [699]. Creation [43, 703, 127]. Creative [43]. Credentials [601]. Criminals [244]. Criteria [760]. Critical [828, 297]. Criticality [772]. CRIWG [966]. Croatian [915, 923]. CRON [805]. Cross [675, 682, 794, 238, 940]. Cross-Domain [238, 940]. Cross-Language [682]. Cross-Lingual [675]. Crowd [43]. Crowdsourcing [515]. Cryptanalysis [644, 645]. Cryptographic [841, 184, 168]. Cryptography [958]. Cryptosystems [643]. CSCL [780]. CSCW [781]. CTL [74]. CUDA [774]. Cultural [701]. Cumulated [687]. Cumulative [743]. Curvature [468]. Curves [472]. Custom [325]. CVFH [475]. Cycle [58, 823]. Cycles [821, 524]. Cyclic [826, 834]. Cyclists [356]. Cyprus [962]. Czech [922, 940, 972].

D [123, 459, 439, 476, 251, 339, 256, 441, 455, 474]. DAGM [950]. Dagstuhl [949]. Dangerous [425, 595]. Data [877, 579, 492, 780, 14, 866, 20, 125, 511, 458, 712, 126, 341, 667, 880, 513, 311, 15, 293, 417, 870, 677, 760, 128, 366, 24, 573, 955, 164, 514, 755, 670, 127, 488, 669, 485, 52, 456, 12, 414, 430, 815, 13, 306, 731, 519, 411, 580, 491, 876, 574, 336, 817, 925]. Data-Aware [866]. Data-Driven [456]. Data-Streams [574]. Database [428, 597, 25, 570]. Databases [294, 664]. Datalog [804, 967, 806, 808, 809, 664, 810, 811, 803, 814, 815, 967, 801, 874, 812, 816]. Dataset [508]. Datasets [710]. Datatypes [731]. DCOPs [91]. Deadlock [40, 898]. Deauville [935]. Debugging [84, 811]. Decade [549]. Decentralized [89, 156, 96, 517, 635]. Decidable [75]. Deciding [170, 168]. Decision [337, 499, 207]. Decision-Making [207]. Decisions [219]. Declarative [806, 811, 895, 554]. Declaratively [863]. Decreasing [198]. Decryption [598]. Deduction [864]. Deep [185]. Deep-Heap [185]. Defeasible [112, 651]. Defense [238]. Defenses [937]. Definability [880]. Defined [71, 252]. Defining [369, 566]. Deformable [480, 359, 481]. Deformation [359]. Degradation [334]. Degree [277, 587]. Delegation [603, 849, 724]. Delegation-Based [724]. Deletion [489]. Delineation [508]. Deliver [30]. Delivery [8, 38]. Delta [97]. DEMACRO [238]. Demands [910]. Democracy [927]. Demographics [503]. Dense [455, 414, 474, 252]. Deontic [886]. Dependence [728]. Dependencies [807]. Dependency [58, 155, 564, 909]. Dependent [78, 121, 742, 394]. Deployments [376]. Depth [480]. Derived [264]. Deriving [873]. Descent [745]. Describing [415, 654]. Description [865, 870, 663]. Descriptions [258, 264]. Descriptor [249, 683]. Descriptors [248]. Design [841, 45, 702, 685, 338, 439, 757, 346, 349, 35, 380, 22, 942, 943, 706, 353, 752, 350, 717, 558, 541, 348]. Designing [783, 5, 347]. Designs [780, 400]. Detecting [165, 174, 175, 917]. Detection [594, 241, 467, 609, 534, 240, 533, 458, 40, 478, 453, 128, 133, 304, 518, 915, 898, 708, 411, 751, 239, 482, 593, 431]. Detectors [236]. Determining [828]. Deterministic [847, 626, 115]. Developers [785]. Developing [313, 549]. Development [941, 32, 355, 808, 302, 786, 31]. Developments [837]. Devices [846]. DFP [428]. DFP-Growth [428]. Dialog [400]. Dialogs [400]. Dialogue [404, 972]. Dialogues [402]. Did [30]. Difference [22]. Different [858, 302, 260, 48]. Differential [644, 585, 571]. Difficulty [560, 586]. Digital [702, 531, 703, 699, 700, 694, 711, 709, 697, 962]. Digitized [496]. Dimensional [70, 832, 839]. Dimensionality [205]. Dione [232]. Direct [726]. Directed [102]. Direction [402]. Direction-Giving [402]. DIRECTions [685]. Directory [878]. Dirty [293]. Disambiguation [912, 906]. Disaster [575, 122]. Discipline [549]. Disclosure [95]. Discovering [237]. Discovery [312, 126, 24, 235, 336]. Discrimination [365]. Discriminative [479]. Discriminatively [153]. Discussion [418]. Disease [19, 13]. Dishonest [639]. Dishonest-Majority [639]. Disjunctive [268]. Disk [232, 97]. Dispersion [83]. Displays [256]. Disproportionality [829]. Dissimilarities [485]. Dissimilarity [473]. Distal [438]. Distance [333, 753]. Distance-Based [753]. Distinctive [251]. Distributed [877, 9, 76, 104, 810, 621, 573, 651, 813, 91, 605, 620, 518, 306, 217]. Distributions [736]. Diverse [314]. Diversity [213]. Divide [148]. Dividend [531]. DL [868]. DL-Programs [868]. DLIN [633]. DNS [178]. Do [346, 470, 694, 48, 803, 697]. Document [597, 332, 696, 304]. Documents [292, 704, 703, 700, 576, 134]. Does [366, 904]. Domain [590, 192, 126, 191, 238, 235, 940, 859]. Domain-Specific [590]. Domains [315]. Double [598]. Doubly [856]. Doubly-Spatial [856]. Down [133, 134, 135, 93]. Drama [400]. Dramatization [52]. Drawing [285, 778]. Driven [892, 51, 557, 894, 324, 564, 456, 419]. Driver [444]. Drivers [228]. Driving [157]. Duality [113]. Duplicate [809, 304]. Duration [434]. during [696]. Duty [647]. DynaLearn [406]. Dynamic [428, 594, 166, 384, 747, 215, 620, 635, 381, 524]. Dynamical [835, 834]. Dynamics [412, 820, 829, 407, 827, 437].

e-Administration [47]. E-Cash [640]. E-Commerce [301]. e-Government [36, 31, 37, 41]. e-Procurement [33]. e-pucks [216]. EA [747]. Earth [433, 768]. Easy [254]. EBEST [46]. ebook [695]. EC4MACS [57]. ECA [827]. ECDLP [642]. eCo [318]. Economic [540]. Economical [144]. ECOS [52]. Ecosystem [369]. Ecosystems [370]. Editing [577]. Education [140, 778, 360, 440]. Educational [338, 946]. Effect [820, 96, 359]. Effective [167]. Effectively [3]. Effects [256, 260, 681, 62, 528]. Efficiency [51, 311, 53]. Efficient [428, 54, 249, 630, 725, 600, 90, 85, 576, 734, 616, 755, 187, 853, 177, 622, 739, 245, 843, 851]. Efficiently [3]. Effort [393]. EGOVIS [927]. EGOVIS/EDEM [927]. eHealth [181]. Eiffel [897]. Eigenvalues [494]. Eigenvariations [252]. Elastic [1]. Elasticity [563]. eLearning [340]. Election [503]. Elections [44]. Electricity [236]. Electronic [927]. Electrophysiological [888]. Element [240]. Elitist [764]. Else [174]. Email [332]. Embedding [205]. Embodied [362]. Emergence [139]. Emergencies [575]. Emergency [795, 517, 796]. Emergent [405, 773]. EMOA [770]. Emotion [128, 501]. Emotional [434]. Empirical [778, 492, 343, 681, 595, 745]. Empowering [878]. Enabled [38]. Enabling [338, 792, 608, 580]. Enclosures [116]. Encrypted [597, 576]. Encryption [856, 596, 627, 2, 636, 634, 853, 598, 635, 626, 849, 854]. End [589, 565]. End-User [565]. Energy [54, 51, 342, 53, 612, 622, 55]. Energy-Efficient [54]. Enforcement [602]. Engagement [701]. Engine [694, 291, 242]. Engineering [892, 757, 720, 942, 526]. Engines [7]. Enhance [358]. Enhanced [859, 701]. Enhancing [533, 163, 761, 24, 655, 303]. Enigma [63]. Enrich [508]. Ensemble [9, 133]. Enterprise [290]. Enterprises [425]. Entities [858, 372]. Entity [675]. Entropy [251]. Envelope [839]. Envelopment [127]. Environment [401, 779, 138, 651, 360, 520, 532, 432]. Environments [247, 124, 215, 250, 260, 606, 398]. eParticipation [30, 44]. Epistemic [653, 168]. Equational [167]. Equi [450]. Equi-affine [450]. Equivalence [848, 101, 300]. Equivalent [191]. Error [452]. Errors [917, 255, 230]. Escalating [229]. Escrow [851]. Escrow-Free [851]. ESORICS [934]. Espionage [229]. Established [57]. Estimation [475, 476, 480]. Estonia [954]. Ethanol [291]. Ethanol-Gasoline [291]. Euler [524]. EUNICE [953]. EUNICE/IFIP [953]. Euro [957, 957]. Euro-Par [957, 957]. EuroMPI [963]. European [934, 963]. Evaluate [124, 691]. Evaluating [792, 730, 127, 236, 122, 55, 912]. Evaluation [685, 364, 123, 687, 492, 780, 961, 686, 611, 136, 4, 445, 353, 444, 816, 33, 543, 745, 37]. Event [557, 319, 422, 518, 554]. Event-Driven [557]. Events [550]. Evidence [112, 343]. Evident [846]. Evolution [740, 509, 757, 822, 654, 836, 413]. Evolutionary [750, 775, 748, 767, 761, 212, 762, 764, 752, 741]. Exact [728]. Examples [24]. Exception [897]. Exchange [880, 858, 815, 857]. Exchanges [173]. Execution [7, 34]. Existence [436]. Existential [804, 867, 268, 872, 802]. Expanded [307]. Expander [825]. Expanding [280, 911]. Expensive [767]. Experience [316, 41, 366]. Experimental [399, 746]. Experiments [676, 920]. Expert [310]. Explaining [84, 328]. Exploiting [443, 8, 727, 758]. Exploration [355, 695, 490]. Exploration/Exploitation [490]. Explorative [415]. Exploratory [786]. Explore [57]. Explorer [705]. Exploring [180, 514]. Exposing [712]. Expression [497, 242]. Expressions [916, 191, 102]. Expressive [105, 271]. Expressiveness [74]. Extended [589, 266, 299]. Extenders [859]. Extending [700]. Extensible [849].

Extension [348]. Extensions [111, 664, 595]. Extent [904]. Extraction [377, 667, 131, 158, 134]. Eye [479].

Face [412, 149, 148]. Face-to-Face [412]. Facebook [787]. Faces [715]. Facial [483, 497]. Facilitate [259]. Facilitates [266]. Facilitating [574]. Factors [36, 31, 413]. Fails [198]. Failure [255]. Failures [893]. Fairness [152, 544]. Fake [172]. Fall [458]. False [64]. Familiar [73]. Farewell [190]. Fast [308, 451]. Faster [774]. Fate [775]. Feasibility [243, 169, 384]. Feasible [68]. Feature [475, 447, 149, 158, 326, 489]. Feature-Based [447]. Features [477, 132, 251, 719, 248, 435]. FedCohesion [39]. Federated [39]. Federation [665]. Feedback [356]. Feet [356]. Female [261]. Femto [384]. Fence [196]. Field [514, 469]. Fifth [261]. File [613, 164]. Filter [470, 752, 855]. Filter-Based [470]. Filtering [146]. Final [392]. Financial [414]. Finding [223, 700]. Findings [778]. Fine [593]. Fine-Grained [593]. Finish [187]. Finite [72]. Finite-State [72]. Firefox [595]. Firewall [532]. Firing [830]. First [807, 314]. First-Order [807]. First-Principles [314]. Fitness [556, 762]. Fitting [254]. Fixation [834]. Fixed [275]. FlashDetect [239]. Flexible [232, 547, 574]. Flow [561, 467, 722, 166, 780, 197, 534, 736, 457]. Flow-Based [534]. Flows [304]. Fluent [895]. FNet [568]. Focus [696]. Focused [690]. Fodor [114]. Forensic [580]. Forensics [175, 709]. Forests [464, 453, 466]. Forgery [482]. Form [152]. Formal [887, 652, 181, 300, 18, 788, 971]. Formalism [380]. Formalizing [100]. Formation [90, 789]. Formed [836]. Forms [869]. Formulation [214]. Forward [636]. Forward-Secure [636]. Foundation [296]. Foundations [949, 971, 939, 817]. Fountain [388]. Four [110]. Fractional [158]. Fracture [499]. Fragment [170, 75]. Frame [102]. Frames [647]. Framework [171, 563, 459, 20, 32, 138, 791, 35, 254, 232, 444, 896, 921, 532, 607]. France [941, 959, 935]. Free [632, 733, 843, 851]. French [540]. Frequency [122, 622]. Frequency-Based [122]. Frequent [428]. Friend [787]. Friendly [182, 698, 37]. Front [11, 28, 506, 527, 548, 569, 583, 604, 625, 646, 662, 674, 693, 714, 735, 756, 777, 799, 819, 840, 862, 883, 901, 902, 905, 924, 49, 60, 81, 99, 120, 141, 162, 183, 204, 225, 246, 267, 288, 309, 330, 351, 368, 389, 396, 410, 421, 442, 463, 484]. Full [659, 95]. Fully [629, 856, 853]. Function [747]. Functional [71]. Functions [740, 281, 716, 743, 686, 906]. Fundamental [491]. Funding [541]. Future [550, 382, 230]. Fuzzy [865, 114, 767, 761].

Game [828, 338, 656, 497, 754]. Games [82, 276, 946, 65, 284, 404, 640, 152, 151, 83, 403, 797]. Gaps [289]. Garbled [637]. Gasoline [291]. Gaussian [736]. GCD [587]. Gender [261]. Gender-Stereotypical [261]. General [151, 853, 242]. General-Sum [151]. Generalized [479, 448, 521]. Generated [705, 316]. Generating [281, 680, 890, 144, 511, 807, 825]. Generation [121, 944, 208, 690, 314, 176, 427, 195, 540]. Generator [280]. Generic [563, 15, 391, 394, 945]. Genetic [738, 538, 746]. Genome [272]. Geographic [519, 952]. Geographically [416, 173]. Geometric [253, 738]. Geometries [450]. Geospatial [670, 526]. German [590, 907]. Germany [949, 956, 966, 938]. Gesture [359, 402]. Get [507]. Getting [903]. GFS [164]. GFS-like [164]. Ghosts [482]. Giant [683]. GIRPR [951]. GIScience [952, 514]. Giving [402]. Global [928, 477, 132, 443, 490, 521]. Globe [925]. GLOW [928]. Goal [324, 655, 650]. Goal-Based [655]. Goal-Driven [324]. Goal-Processing [650]. Goal-Specific [324]. Goes [464]. Going [678]. Gonna [912]. Goods [425]. Google [694]. Governance [866, 927]. Governing [379]. Government [32, 41, 36, 927, 31, 29, 33, 37]. GP [752]. GPP [242]. GPP-Grep [242]. GPU [730, 751]. GPUs [774]. Graders [261]. Gradient [490]. Graduate [140]. Grain [822]. Grained [593]. Gram [243]. Grammars [891, 102]. Grammaticalized [922]. Granularity [256]. Graph [750, 279, 891, 427, 248, 802, 494, 259, 97]. Graph-Associated [494]. Graph-Based [750, 802]. Graphical [492, 491]. Graphics [331]. Graphs [277, 285, 683, 825, 654]. Grayscale [504]. Graz [950]. Greece [947, 957, 968]. Green [93, 56, 52]. Grep [242]. Grid [925, 236, 376]. Gröbner [190]. Ground [358]. Group [342, 789, 963]. Group-Based [342]. Groups [841, 280, 858, 418, 857, 854]. Groupware [781]. Growth [428]. Guarantee [440]. Guaranteed [542]. Guarantor [45]. Guest [227]. Guest-Assisted [227]. Guide [776]. Guided [763, 607, 237]. Guideline [290, 729].

Hadoop [816]. HAID [943]. Hand [474]. Hand-Held [474]. Handling [755, 897]. Hands [720]. Hands-on [720]. Happen [187]. Haptic [364, 355, 353, 358, 354, 943]. Haptic-Audio [354]. HapticPulse [357]. Haptics [365]. Hard [283]. Hardening [164]. Hardware [716, 893]. Harmony [121]. Harnessing [316]. Hash [633, 850, 535, 859]. Hash-and-Sign [633]. Hash-Based [535]. Hashing [248]. Hate [718]. Heading [255]. Health [581, 341]. Heap [185]. Heart [357]. Held [474]. Help [904]. Hemisphere [255]. Heritage [701]. Heterogeneous [384, 737, 221]. Heuristic [68]. Hidden [128, 684]. Hiding [430]. Hierarchical [146, 599, 460, 487, 753]. Hierarchies [407, 465]. Hierarchy [466]. High [537, 343, 682, 128, 423, 22, 545, 518, 469, 242]. High-Level [537, 343, 518]. High-Order [128, 469]. High-Speed [242]. Higher [140, 587]. Histogram [475]. Historical [704, 703]. History [78, 632, 776]. History-Free [632]. Hoc [517, 233]. Home [9, 378, 17]. Homogeneous [117, 737]. Homomorphic [627, 598]. hop [529]. Hospital [14]. Hoster [234]. Hough [479]. Household [364]. Houses [144]. How-to [131]. HP [765]. HPC [613, 623]. HSDPA [386]. HTTP [544]. Human [173, 757, 399, 700]. Human-Centred [700]. Humanities [704]. Hungary [953]. Hunting [295]. Hybrid [720, 149, 659, 161, 220, 612, 210, 724]. Hydrography [508]. Hypertrophic [23].

I/O [614]. IAPR [951]. ICCBR [941]. ICICA [948]. ICT [928, 928]. ICT-GLOW [928]. ID [240, 725, 844]. Idea [298]. Identification [289, 308, 599, 826, 742]. Identifying [710, 893]. Identity [590, 39, 850, 599, 483, 38, 626, 851]. Identity-Based [850, 599, 626, 851]. Idiomatic [916]. IEEE [376, 528, 542]. IFIP [940]. IFP [642]. IGO [736]. IGO-Flow [736]. II [835, 965, 827]. ILE [406]. ILU [54]. Image [750, 132, 677, 158, 455, 446, 451, 471, 752, 482, 465, 435]. Image-Based [451]. Imagery [515]. Images [478, 480, 470, 415, 504]. Imaging [452]. Immature [289]. Impact [722, 716, 551, 531, 824, 536, 525]. Impacts [48]. Impaired [154]. Imperative [190]. Imperfections [386]. Implementation [45, 722, 734, 717, 543]. Implementing [639, 5, 658]. implication [114]. Implicit [454]. Improve [42]. Improved [765, 637]. Improving [356, 723, 134, 495, 211, 8, 876, 544]. Impulsive [436, 437]. IMS [537]. Incentive [559]. Incidents [157]. Including [719]. Inclusive [560]. Incomplete [194]. Inconsistency [812]. Inconsistency-Tolerant [812]. Inconsistent [668]. Incorporate [255]. Incorporating [152]. Incorrect [691]. Increased [728]. Increasing [704, 406]. Incremental [487]. Independence [868, 118]. Independent [690, 609]. Index [829, 578, 568]. Indexed [945]. Indexing [132, 308, 576]. India [50]. Indicators [589, 56, 566]. Individuals [784]. Indoor [247, 256, 523, 525]. Inductive [899]. Industrial [887, 229]. Industry [967, 968]. INEX [676]. Inexact [602]. Infer [192]. Inference [190, 252]. Inferring [404]. Infinite [73]. Influence [261, 520]. Influencing [36, 413]. Informal [134]. Informatics [926]. Information [377, 561, 307, 334, 166, 14, 125, 640, 32, 335, 256, 621, 134, 927, 910, 637, 424, 696, 304, 42, 931, 940, 430, 106, 758, 525, 572, 952, 926, 961, 956, 948, 953]. Information-Flow [166]. Information-Theoretic [637]. Infrastructure [685, 449, 3]. Infrastructures [226]. Inhibition [65]. Inhibits [888]. Initial [116]. Initiality [109]. Initiated [530]. Initiative [961]. Injection [231, 172]. Inner [854]. Inner-Product [854]. Innovative [350]. INNS [951]. Inpainting [468]. Input [792, 626, 593]. Inquiry [294]. Insects [97]. Insertion [196]. Inspection [32, 481]. Inspired [221]. Instance [501]. Integer [196]. Integrated [144, 57]. Integrating [307, 882, 411]. Integration [290, 383, 126, 341, 41, 794, 380, 433, 260, 430, 471, 558, 336, 37]. Integrative [18]. Intelligence [241, 933, 936, 142]. Intelligent [878, 12, 932, 398]. Intensive [12]. Intention [36]. Inter [40]. Inter-organizational [40]. Interacting [359]. Interaction [364, 331, 863, 124, 443, 943, 418, 363, 402]. Interactions [152, 57, 362]. Interactive [748, 676, 360, 465]. Interconnect [716]. Interdental [500]. Interdependent [86]. Interest [373]. Interface [78, 732, 353, 963, 354]. Interfaces [722, 795, 201, 297, 359, 788]. Interfacing [397]. Interference [22]. Intergroup [342]. International [941, 944, 933, 947, 928, 937, 967, 954, 946, 926, 961, 964, 965, 949, 936, 960, 959, 945, 956, 925, 966, 957, 927, 929, 970, 948, 942, 943, 935, 931, 940, 930, 939, 968, 972, 938, 953, 969, 958, 952, 962]. Internet [369, 45, 944, 315, 432]. Internet-of-Things [369]. Interoperability [302, 608]. Interplay [803]. Interpolation [417]. Interpretability [761]. Interpretation [194, 203]. Interpreters [392]. Interrupt [894]. Interrupt-Driven [894]. Intra [386]. Intra-site [386]. Introducing [212]. Introduction [453, 666]. Introspection [227]. Intrusion [241, 240, 533]. Intrusions [937]. Invariance [659]. Invariant [450, 195, 105]. Invariants [190]. Invasion [97]. Inverse [587]. Inversion [438]. Investigating [894]. Investigation [888, 179, 836]. Investigations [522]. IOT [370, 371, 372]. IP [678]. IPsec [177]. IPv6 [536]. IR [685, 689, 681]. Iris [495]. Irreducibility [167]. ISC [956]. Island [957, 968, 737]. Isolated [746]. Isomorphism [279]. Isotropic [740, 736]. Issue [534]. Issues [5]. Istanbul [955]. IT-Based [552]. IT-Cooling [50]. IT-Systems [50]. Italy [961, 964, 965, 934, 951, 958]. ITBAM [926]. Iterative [464, 824, 766, 515].

Japan [942]. JavaScript [186]. Job [618, 624]. Jobs [623]. Join [197, 560]. Joined [41]. Joined-Up [41]. Joint [459, 927, 386, 950]. JPEG [482]. July [949]. Jumps [209]. June [949]. Justified [112].

Kademlia [533]. Keeping [787]. Kernel [486, 734]. Kernel-Level [734]. Key [928, 845, 176, 858, 785, 626, 857, 429]. Keys [591]. Keyword [576, 852, 855]. Kit [615]. Kloster [938]. Knapsack [208, 643, 770]. Knocking [364]. Know [903]. Knowing [648]. Knowledge [312, 140, 112, 863, 628, 126, 130, 24, 657, 810, 139, 127, 266, 565, 122, 932, 29, 336, 354, 932]. KPD [592]. Kriging [417]. Kuching [933, 930, 932].

L1 [467]. Labeling [465]. Landmarks [352]. Landscapes [745]. Language [307, 690, 800, 682, 931, 681, 697]. Languages [281, 71, 670, 666, 658, 817]. LANs [177]. Large [949, 337, 2, 448, 304, 654, 920, 816, 3]. Large-Scale [949, 448, 920, 816, 3]. Latencies [621]. Lattice [628, 822]. Lattice-Grain [822]. Lattices [865, 629, 626]. Layer [542]. Lazy [328, 129]. Leakage [177, 572, 849]. Leaping [199]. Learner [918, 406]. Learning [78, 416, 680, 792, 492, 401, 9, 438, 145, 349, 339, 201, 793, 256, 90, 138, 151, 486, 218, 324, 483, 333, 487, 328, 485, 360, 753, 259, 754, 491, 593, 797]. Least [275]. Lectures [945]. Left [255]. Left-Right [255]. Legal [147, 130]. Legibility [521]. Legislative [713]. Length [749]. Lessons [676]. Level [537, 343, 385, 734, 518]. Levels [552]. Leveraging [54, 732]. Lexical [921]. Lexicalized [102]. Lexicon [923]. Lexicons [921, 911]. Libraries [730, 703, 698, 711, 962]. Library [702, 318, 699, 700, 694, 695, 709, 717]. Lies [247]. Life [828, 589, 58, 157, 397]. Life-Logging [157]. Lightweight [666]. Like [827, 164]. Limit [823]. Limitations [271]. Line [640, 371]. Linear [744, 743, 618, 812, 73]. Linear-Time [73]. Linearity [115]. Linearizability [79]. Linearizing [108]. Lingual [675]. Linguistic [712, 265]. Link [19, 528]. Linkability [180]. Linked [712, 514, 670, 669, 876]. linkedscience [514].

Linking [537, 675, 713, 573]. Listen [699]. Literacy [910]. Little [824]. Live [876]. Livelock [40]. Liveness [202]. Liveness-Based [202]. LLR [644]. Load [621, 617]. Local [477, 132, 641, 149, 323, 77, 545, 490]. LocalAlert [517]. Localising [175]. Localization [171, 222, 479, 451]. Localized [466]. Location [592, 373, 335, 699, 374, 172, 525, 431]. Location-Triggered [699]. Locations [415]. Logging [157]. Logic [886, 112, 877, 959, 179, 610, 873, 75, 895, 440, 899, 650, 931]. Logical [106, 817]. LogicBlox [800]. Logics [865, 870, 649, 663, 119, 105, 271]. Logistic [381]. Logistics [425]. Logs [240, 554]. Lone [241]. Loop [199]. Losing [707]. Lost [707]. Love [718]. Love-Hate [718]. Low [722, 439, 628, 535, 577]. Low-Cost [535, 577]. LR [849]. LR-UESDE [849]. LTE [383, 385, 384, 388]. LTSs [73]. Lund [943]. Lung [321]. Lustre [613]. Lyon [941].

M [793]. M-Learning [793]. M3 [380]. MA [947]. MAC [530]. Machine [43, 243, 154, 485, 532]. Machines [100, 227, 771, 64]. Macro [384]. Macro-Femto [384]. MACVILLE [138]. Magic [804]. Magic-Sets [804]. Mainly [815]. Majoring [108]. Majority [639]. Makes [75]. Makespan [741]. Making [68, 194, 302, 337, 207]. Malaysia [933, 930, 932]. Male [261]. Malicious [238]. Malleability [624]. Malleable [634, 622]. Malware [239]. Mammograms [496]. Man [176, 129]. Man-In-The-Middle [176]. Manage [46]. Managed [16]. Management [334, 140, 954, 575, 39, 779, 557, 332, 925, 139, 955, 552, 425, 122, 932, 5, 29, 549]. Managing [562, 43, 318, 3]. MANETs [171]. Manifolds [456]. Manipulation [52]. Manual [609, 918]. Manually [919]. Manufacturing [557]. Many [770, 759, 707, 724]. Many-Core [724]. Many-Objective [770, 759]. Manycores [616]. Map [521, 252]. Mapping [253, 513, 250, 921]. MapReduce [337, 4, 623]. Maps [511, 157, 266, 503]. March [945]. Marche [39]. Marine [125]. Markers [916]. Market [540]. Markets [165]. Markov [128, 684]. Mashup [290]. Mashups [379]. Mask [158]. Massive [306]. Matchballs [797]. Matching [638, 422, 423, 776, 473]. Matchmaking [864]. Material [480, 359]. Mathematical [939]. Mathematically [191]. Matrices [283, 494]. Matter [10, 27, 59, 98, 367, 395, 409, 420, 462, 505, 582, 661, 673, 692, 798, 818, 861, 900, 11, 28, 506, 527, 548, 569, 583, 604, 625, 646, 662, 674, 693, 714, 735, 756, 777, 799, 819, 840, 862, 883, 901, 902, 905, 924, 49, 60, 81, 99, 120, 141, 162, 183, 204, 225, 246, 267, 288, 309, 330, 351, 368, 389, 396, 410, 421, 442, 463, 484, 784, 815]. Maturity [552]. Maximum [277]. May [946, 187]. McEliece [591]. Mean [22, 469]. Means [158]. Measured [295]. Measures [686]. Measuring [589, 556, 319, 213, 520]. Mechanism [598, 16, 402, 87]. Mechanisms [375, 86, 559]. Media [947, 361, 415, 707, 907, 413, 411, 419]. Mediated [339]. Mediation [380]. Medical [926, 341, 15, 677, 573, 340]. Medicine [23, 12]. Medium [224]. Meeting [963]. Meets [61]. Mekong [97]. Memory [231, 349, 258, 732, 487, 545, 591, 733, 593, 230].

Memory-Constrained [591]. Men [175]. Men-in-the-Middle [175]. Mental [694, 264, 261, 257]. Mental-Rotation [261]. Mere [811]. Merging [592, 660]. Mesh [528]. Message [173, 487, 963, 733]. Messages [303]. Meta [15, 871, 650]. Meta-Data-Models [15]. Meta-logic [650]. Metadata [713]. Metafora [791]. Metalinguistic [916]. Metaverse [402]. Meters [182]. Method [750, 467, 887, 427, 300, 616, 433, 91, 440, 94, 431]. Methodological [382]. Methods [533, 344, 971, 50, 106]. Metric [687, 578]. Metric-Based [578]. Metrics [556, 303]. MFCS [939]. Micro [730, 731]. Micro-applications [731]. Micro-Benchmark [730]. Microenvironment [20]. Microsoft [885]. Middle [176, 175]. Middleware [547, 398]. Migration [619, 92, 532, 538]. Milner [61]. MIMO [383]. Miner [206, 211]. Minimization [113]. Minimum [286, 210]. Minimum-Weight [210]. Mining [428, 947, 14, 20, 9, 916, 155, 755, 415, 414, 13, 907]. Missing [691]. MITM [174]. MKNF [657]. MLP [495]. Mobile [377, 589, 253, 792, 355, 256, 699, 302, 333, 698, 433, 17, 16, 588, 233]. Mobile-Certificate [433]. Mobile-Friendly [698]. modal [419]. Modality [366, 909]. Mode [822]. Model [886, 307, 537, 110, 492, 132, 888, 438, 618, 551, 781, 74, 892, 476, 772, 470, 765, 600, 83, 254, 128, 337, 169, 842, 813, 603, 547, 122, 299, 896, 844, 419, 491, 523, 572, 245, 336, 684]. Model-Based [765]. Model-Driven [892, 419]. Model-Theoretic [813]. Modeled [152]. Modeling [307, 551, 163, 194, 434, 612, 42, 57, 559, 472, 34, 97, 788, 947]. Modelling [123, 58, 387, 746]. Models [561, 835, 537, 137, 335, 567, 15, 829, 124, 694, 755, 670, 370, 515, 737, 782, 654, 834, 452, 273, 347, 94, 554]. Modular [841, 188, 216]. Modulation [439]. MOEA [766, 769, 763]. MOEA/D [766, 763]. Monitoring [227, 32, 104, 1, 26, 371, 232, 17, 97, 431]. Monotonic [740]. Montpellier [959]. Morgan [865]. Morphological [504]. Morphosyntactic [912]. Mortality [283]. Mortals [811]. Motion [459, 102, 262, 456]. Motivation [406]. Move [600, 842]. Movement [173, 551]. Moving [110]. MPC [639]. MPI [963, 716, 730, 609, 718, 723, 725, 719, 715, 732, 734, 731, 898, 717, 729, 724]. MPI-3 [719]. MPICH2 [721]. mpicroscope [729]. MRF [468]. MRF-Based [468]. MRFs [470, 469]. MSM [947]. MTL [111]. Multi [89, 331, 208, 383, 863, 530, 70, 66, 767, 222, 476, 959, 478, 453, 179, 138, 83, 657, 139, 762, 159, 460, 648, 53, 388, 764, 360, 214, 529, 930, 833, 768, 419, 624, 749, 758, 788, 658, 437, 498, 797]. Multi-Agent [959, 930, 89, 138, 83, 139, 648, 658]. Multi-Agent-System [797]. Multi-armed [214]. Multi-beam [383]. Multi-channel [530, 179]. Multi-class [453]. Multi-Context [657, 159, 863]. Multi-core [53, 624]. Multi-Dimensional [70]. Multi-hop [529]. Multi-modal [419]. Multi-net [498]. Multi-Objective [767, 762, 764, 208, 833, 768, 749, 758]. Multi-point [388]. Multi-Prioritized [476]. Multi-Pushdown [66]. Multi-resolution [460]. Multi-robot [222]. Multi-sensory [360]. Multi-species [437]. Multi-touch [331]. Multi-user [331, 788]. Multi-view [478]. Multicast [388]. Multicasting [546]. Multidimensional [51]. Multidisciplinary [940]. Multiflow [386]. Multilinguality [961]. Multimodal [331, 679, 402]. Multimodality [961]. Multiobjective [761, 760]. Multiobjectivization [765]. Multipath [543]. Multiple [644, 792, 146, 476, 341, 302, 36, 91, 774]. Multipurpose [244]. Multithreaded [725]. Multitype [835]. MUSE [947]. Music [308, 679]. MusiClef [679]. Mutual [535, 327]. My [707]. myCBR [313].

Naïve [816]. Named [511]. Names [235]. Nano [360]. Narrative [405]. Nasty [274]. Natal [971]. National [508]. Natural [470]. Nature [964, 965, 261]. Navigation [665, 352]. Near [280, 304]. Near-Optimal [280]. Needs [701]. Negation [869]. Negative [739]. Negotiation [85, 601]. Neighborhood [759]. Nested [80]. net [498]. Netherlands [937]. Network [241, 383, 656, 240, 533, 243, 47, 422, 763, 414, 814, 493, 860, 558, 97, 538, 429, 796, 504]. Networking [944]. Networks [377, 613, 173, 69, 656, 530, 534, 375, 500, 19, 660, 342, 179, 417, 387, 321, 337, 726, 423, 384, 291, 161, 424, 433, 786, 244, 951, 495, 236, 529, 542, 413, 958, 540, 233, 3]. Neural [500, 321, 291, 951, 495, 493, 504]. Neuroscience [24]. NEW2AN [944]. Newcastle [929]. Next [944, 540]. Next-Generation [540]. NK [745]. NK-Landscapes [745]. NLP [904]. No [241, 174]. Node [539]. Nodes [429]. Noised [2]. Noisy [222]. Nominative [842]. Non [848, 115, 621, 634, 452, 645, 733]. Non- [848]. Non-coherent [733]. Non-commutative [645]. Non-deterministic [115]. Non-Malleable [634]. Non-perspective [452]. Non-uniform [621]. Nonlinear [205]. Normal [152, 782]. Normal-Form [152]. Norms [93]. North [264]. North-Up [264]. Notes [107]. Notions [848, 586, 491]. Novel [459, 492, 148, 422, 493, 491]. Novelty [212]. NP [283, 920]. NP-Hard [283]. NSDG [38]. Nuclear [709]. Numerical [660, 321].

O [614]. OAGM [950]. Obfuscation [194]. Object [475, 459, 249, 476, 478, 453, 441, 248]. Objective [767, 770, 762, 759, 764, 91, 763, 363, 776, 208, 833, 768, 749, 758]. Objective-Guided [763]. Objectives [70, 189, 57]. Objects [476, 480, 450, 362, 454, 498]. Obligations [647]. Observation [461]. Observations [768]. Observer [884]. Observing [341, 768]. Obstructive [13]. Occlusion [467]. October [947]. Off [640]. Off-Line [640]. Offload [722]. Offs [490]. OH [952]. OMB [730]. OMB-GPU [730]. Once [846]. One [174, 732, 234, 842, 721]. One-Click [234]. One-Move [842]. One-Sided [732, 721]. Ongoing [325]. Online [424, 244, 250, 418, 413, 33]. Only [648]. Only-Knowing [648]. Ontological [268]. Ontologies [667, 874, 327, 668]. Ontology [140, 877, 289, 866, 150, 666, 875, 526, 797]. Ontology-Based [140, 289, 866, 797]. Opacity [839]. Open [7, 712, 134, 371, 220, 88]. Open-System [88]. Operation [45, 334, 144, 425]. Operational [596, 838]. Operations [7, 727]. operative [312]. Operators [763, 73]. Opinion [92, 907, 911]. Opportunistic [656].

Optical [467, 457, 494]. Optimal [280, 631, 656, 223, 85, 151, 587]. Optimistic [791]. Optimization [748, 767, 223, 772, 18, 762, 759, 443, 766, 220, 605, 776, 833, 350, 768, 538, 749, 217]. Optimizer [213, 218]. Optimizers [215]. Optimizing [509, 915, 816]. OR-Join [560]. Oracle [843]. Oracle-Free [843]. Oracles [599, 847, 857]. Order [806, 807, 128, 469]. Ordering [512]. orders [108]. organising [775]. Organismic [301]. Organization [652]. Organizational [603, 40]. Organized [772]. Organizing [503, 403]. Oriented [475, 401, 41, 293, 264]. Origin [255]. Osaka [942]. OSDM [603]. Other [57]. OUR [475]. OUR-CVFH [475]. Outdoor [154, 456, 949]. Outerplanar [277]. Output [806]. Outsourced [594]. Overarching [781]. Overlap [723, 728]. Overlay [533]. Overload [376]. Overview [377, 452]. OWL [882, 666, 668, 539]. OWL-Based [539]. Ownership [79]. Oxford [945].

P2P [925]. PA [40]. Pachinko [132]. Pacific [933, 932]. Packet [172, 233]. Pain [312]. Pairwise [153]. Palytoxin [888]. Papers [947, 946, 949]. Paphos [962]. Parallel [619, 720, 616, 957, 187, 515, 771, 724, 964, 965]. Parallelism [54, 187]. Parallelization [607]. Parameter [90, 539, 491]. Parameterized [282, 269, 741]. Parameters [772]. Parametric [189]. Parametrized [195]. Parents [770]. Parity [284]. Parliamentary [713]. Parser [155]. Parsimony [749]. Part [964, 965, 248, 73, 129]. Part-Graph [248]. Part-of-Speech [129]. Partial [640]. Partially [631, 783]. Partially-Blind [631]. Participating [580]. Participation [418]. Participatory [48]. Particle [835, 772, 18, 213, 218, 205, 215, 220, 773, 776, 209]. Parties [174]. Partition [616]. Partitioning [89, 214, 217]. Partners [258]. Parts [481]. Party [637, 425]. Passau [956]. Passing [487, 963, 733]. Passport [507]. Password [426]. Past [653, 230]. Patched [178]. Patent [678, 681]. Patents [919]. Path [200, 260, 529]. Path-Sensitive [200]. Patient [26, 17, 16]. Patients [910]. Patrolling [88]. Pattern [638, 609, 415, 951, 950]. Pattern-Independent [609]. Patterns [428, 891, 867, 341, 423, 493, 731, 836, 566]. Paying [234]. Payload [231]. Pedagogical [794, 406]. Pedestrian [443]. PEKS [852]. Penalty [686]. 1 [770]. Covertly [639]. D [766, 763]. EDEM [927]. Exploitation [490]. IFIP [953]. MapReduce-Based [3]. P [616]. Palate [500]. TC [940]. TLS [174, 175]. XK [727]. Penetration [536]. People [161, 796]. PEPA [207]. Perceptual [123]. Perfect [653, 852]. Performance [613, 615, 56, 720, 385, 219, 457, 614, 611, 545, 261, 386, 528, 6, 729, 566]. Periodic [436]. Permission [163]. Permissiveness [166]. Permutation [280]. Persistent [898]. Persistent-Sets [898]. Personal [139, 549]. Personalized [146]. Perspective [781, 258, 801, 36, 927, 452, 6]. Perspectives [671, 29]. Perturbations [820]. Petersburg [944]. Pharmaceutical [919]. Phishing [235]. Phone [588]. Phonetic [908]. Photography [441]. Physical [700, 352, 357, 347]. Pilot [399]. Pipeline [431]. Piracy [234]. Pisa [934]. PKAW [932]. Places [373]. Plagiarism [691]. Plaintext [634]. Plan [314]. Planar [285]. Planning [507, 137, 314, 323, 763, 758, 658]. Platform [581, 41, 800, 371, 397, 46]. Platforms [302, 6, 591]. Player [152]. Playing [401]. PLC [887]. Point [460, 830, 255, 541, 431, 388]. Point-to-Origin [255]. Pointer [202]. Points [285, 275]. PoisonAmplifier [237]. Poisoning [237]. Policies [575, 602, 324, 93]. Policy [300, 57]. Polish [703, 917, 920]. Polling [723]. Polygon [285]. Polyhedral [189]. Polynomial [277, 190, 879, 116]. Polynomial-Time [277, 116]. Polynomials [249]. Pool [95]. Population [759]. Portfolio [322]. Pose [475, 476, 480]. Position [687]. Positioning [525]. Positions [855]. Possibilistic [159]. Possible [782]. Post [312]. Post-operative [312]. Poster [244, 245]. Potassium [888]. Potency [194]. Potential [801, 157]. Power [439, 528, 105, 271]. Powered [142]. PPM [233]. PPSN [964, 965]. Practical [14, 176, 720]. Practice [940, 930, 571, 962]. Pragmatics [404]. Prague [940, 909]. Preamble [530]. Preconditioners [54, 448]. Predator [436, 437]. Predator-Prey [436]. Predicate [196, 873, 854]. Predicates [105]. Predicting [247]. Prediction [19, 765, 326, 291, 499, 611, 654]. Preference [143]. Preferences [570]. Preliminary [647, 156, 159, 136]. Presence [386]. Present [230]. Preservation [705]. Preserve [859]. Preserving [296, 630, 633, 596, 706, 584]. Presidential [503]. Pressure [17, 749]. PrEV [705]. Prey [436]. PRICAI [933]. PRIMA [930]. Primary [778]. Principles [45, 314, 757, 930, 265]. Prior [468, 758]. Prioritized [476]. Priorlets [252]. Priors [252]. Privacy [173, 556, 296, 181, 573, 182, 570, 852, 584, 172, 571, 419]. Privacy-Aware [173]. Privacy-Preserving [584]. Private [591]. PRIVATUS [182]. Proactive [235, 382]. Probabilistic [171, 492, 888, 83, 118, 874, 172, 491]. Probability [493]. Problem [89, 744, 282, 286, 492, 964, 965, 66, 880, 322, 830, 737, 220, 210, 214, 832, 742, 771, 591, 491]. Problems [14, 767, 116, 736, 770, 161, 587, 766, 769, 746]. Procedures [42, 34]. Proceedings [964, 965, 713, 941, 944, 933, 928, 937, 967, 954, 926, 961, 936, 960, 959, 934, 971, 956, 925, 966, 955, 957, 927, 929, 970, 948, 942, 943, 951, 935, 931, 950, 940, 930, 932, 939, 968, 972, 938, 953, 969, 963, 958, 952, 962]. Process [561, 563, 954, 562, 890, 32, 551, 553, 557, 567, 150, 564, 782, 558, 568, 554, 566, 549]. Process-Based [558]. Process-Driven [564]. Processes [702, 58, 866, 43, 40, 139, 706, 658]. Processing [957, 669, 545, 650, 242]. Processors [619, 242]. Procurement [33]. Procurements [87]. Product [108, 316, 336, 854]. Production [135]. Profiles [666]. Profitability [761]. Program [746]. Programming [877, 618, 390, 945, 391, 732, 337, 871, 738, 746, 394]. Programs [196, 887, 190, 609, 76, 868, 894, 53, 813, 67]. Project [345, 320, 52]. Projects [339]. Promised [30]. Pronoun [160]. Proof [850, 86]. Proofs [184, 628]. Propagation [424]. Propagative [822]. Properties [891, 76, 345, 117, 480, 219, 894, 895, 859, 168, 697]. Proposal [44].

Proposals [345]. Propositional [886, 118]. Protection [164, 182]. Protein [765]. Protocol [639, 181, 167, 535, 85, 610, 573, 733, 857]. Protocols [80, 176, 179, 243, 168]. Prototype [488, 545]. Provable [969]. Provenance [579, 5]. Provide [47]. Proving [271]. Provision [95]. ProvSec [969]. Prudent [136]. Pseudo [680, 747]. Pseudo-Boolean [747]. Pseudonymous [590]. PSO [221]. Public [82, 845, 42, 591, 626]. Public-Key [845]. Publications [710, 706]. pucks [216]. Pulmonary [13]. Pulse [439]. Pump [888]. Purified [432]. Purpose [242]. Pushdown [66]. Puzzles [586]. Pyramid [446].

QAPs [774]. QoS [793]. Quadratic [736]. Qualitative [263, 523]. Quality [364, 551, 57, 382, 481]. Quantified [74]. Quantifiers [804]. Quantifying [516, 572, 107]. Quantitative [70, 602, 71]. Quantitatively [485]. Quantization [486]. Quantum [106, 67]. Quasi [108, 577]. Quasi-orders [108]. Quasi-Realtime [577]. Quaternion [645]. Queries [684]. Query [307, 690, 867, 960, 809, 869, 131, 670, 812, 663, 875, 817]. Querying [20, 150, 870, 876, 568]. Question [131, 325]. Question-Answering [131]. Quick [254]. Quo [698, 645].

Races [269]. Radio [387]. Radiotherapy [323]. Raesfeld [966]. RAID [937]. Railway [425]. Rainbow [645]. Ramsey [117]. Random [278, 599, 453, 466, 847, 857, 843]. Randomness [117]. Range [471]. Rank [680]. Ranking [687, 864, 345]. RANSAC [476]. Raster [511]. Rate [542, 357]. Rating [473]. Ray [478]. RBAC [603]. RDM [136]. RDR [136]. Re [704, 652, 700]. Re-finding [700]. Re-use [704]. Reachability [66, 826, 67]. Reachable [891]. Real [241, 240, 335, 458, 949, 319, 371, 333, 445, 340, 751]. Real-Time [241, 335, 319, 333, 340, 751]. Real-World [949, 445]. Realism [358]. Reality [567, 678]. Realization [778]. Realtime [577]. Reason [21]. Reasoning [941, 312, 310, 313, 667, 311, 268, 321, 810, 323, 159, 326, 970, 651, 118, 649, 328, 663, 668, 143, 265, 672, 876, 348, 960]. Reassignment [86]. Reassignment-Based [86]. Recall [704, 653, 876]. Receiver [530, 172]. Receiver-Initiated [530]. Receiver-Location [172]. Recognition [475, 459, 476, 149, 154, 951, 495, 501, 950, 497, 493, 494]. Recognizing [449]. Recombination [770, 759]. Recommendation [125, 161, 565, 796]. Recommendation-Based [125]. reconfigurable [216]. Reconstruct [272]. Reconstructing [449]. Reconstruction [459, 460, 455, 814, 474]. Record [713]. Recoverable [2]. Rectangular [509]. Recursive [273]. Reduced [585]. Reducibility [101]. Reducing [160]. Reduction [879, 297, 205, 898]. Reductions [278]. Reference [600]. References [710, 691]. Refinement [675, 145, 892]. Reflect [567]. reflection [195]. Reflections [549]. Refuse [524]. Region [39, 97]. Regression [416]. Regular [242]. Regularized [153]. Rehabilitation [24]. Reinforcement [145, 754]. Related [25, 235]. Relational [294, 570, 265]. Relations [58, 512, 101, 263]. Relationship [718, 169, 257, 524]. Relationship-Based [169]. Relative [687]. Relevance [688]. Reliability [940, 688, 525]. Remember [697]. Reoptimizing [286]. Repairing [567]. Repeatable [475]. Replacing [560]. Report [647, 159]. Reports [130]. Represent [21]. Representation [366, 191, 264, 299, 908]. Representations [258, 321, 203, 749]. Representative [488]. Represented [72]. Representing [870]. Reproducibility [311]. Republic [940, 972]. Reputation [688, 94]. Reputation-Based [94]. Requests [238]. Requirements [296, 793, 564, 655]. Research [941, 289, 704, 438, 51, 934, 441, 514, 425, 940, 968, 435, 432, 937]. Residuated [865]. Resilience [586, 177]. Resilient [849]. Resolution [809, 160, 904, 516, 348, 460]. Resource [278, 335, 921, 440]. Resource-Bounded [278]. Resources [408, 824, 547, 95, 707]. Response [907]. Responsible [546]. Restricted [869]. Results [110, 780, 156]. Retrieval [307, 597, 132, 125, 864, 686, 158, 676, 435]. Reusability [780]. Reusable [318]. Reusing [324]. Reveal [357]. Revealing [179]. Reversible [821, 658]. Reversing [237]. Reviews [180, 316]. Revised [947, 946, 949, 945]. Revision [262]. Revisited [598, 908]. Revisiting [586, 860]. Revolution [707]. Reward [234]. Rewards [559]. Rewritable [807]. Rewriting [809, 155, 812, 875]. RFID [535]. RGB [504]. Rhodes [957]. Ridesharing [381]. RIF [803]. Right [255]. Rim [933, 932]. Ring [627, 843]. Rings [645]. RIPEMD [585]. RIPEMD-160 [585]. Ripple [133, 134, 135]. Ripple-Down [134, 135]. RNA [22]. Road [589, 511, 751]. Robot [207, 222]. Robotics [212, 214]. Robots [253]. Robust [282, 227, 486, 172, 245, 87]. Robustness [496]. Role [349, 90, 264, 299, 16]. Role-Based [299]. Roles [412, 139]. Rome [961]. Rooted [210]. Rotation [261, 257]. Rotational [261]. Roubens [114]. Rough [344]. Round [631]. Round-Optimal [631]. Rounding [441]. Route [758]. Routed [173]. Routing [224, 220]. RR [970]. Rule [881, 323, 922, 970, 827]. Rule-Based [881, 323, 922]. Rules [867, 268, 155, 133, 134, 135, 882, 872, 873, 103, 802, 832, 897]. Runtime [748, 741]. ruSMART [944]. Russia [944].

SA [744]. SA-ES [744]. Safe [374]. Safety [297]. Safety-Critical [297]. Sampled [472]. Samples [460]. Sampling [615, 755, 306, 739]. Santorini [968]. Sarawak [930]. SAS [935]. Satellite [433, 515]. Satellite-Earth [433]. Satellites [768]. Satisfaction [276]. Satisfiability [322, 169]. Saturated [118]. Saudi [31]. Save [528]. Saving [342]. SBMF [971]. Scalability [215, 5]. Scalable [341, 371, 545, 403]. Scale [949, 448, 446, 920, 451, 816, 3]. Scale-Space [446]. Scaling [619, 622]. Scan [252]. Scanning [363]. Scanning-Based [363]. Scenario [796]. Scenarios [792, 864, 794]. Scene [949, 454]. Scenes [522]. Scheduler [542]. Scheduling [618, 320, 620, 831, 386, 771, 622, 741, 768, 617, 624, 623]. Scheme [231, 630, 387, 323, 634, 386, 844, 426, 233]. Schemes [845, 234, 853, 529, 273]. School [778, 960, 945]. Science [939, 952].

Scientific [680, 311, 706]. SCN [958]. Scoring [12]. Scott [270]. SCR [890]. Screens [22]. Scripted [400]. Scripts [780, 427]. SDM [955]. Seal [846]. Seal-Once [846]. Seaport [126]. Search [121, 308, 223, 212, 682, 576, 694, 91, 620, 855, 913, 237]. Searchers [678]. Searching [832]. Second [928, 967, 397, 962]. Secret [2, 860]. Secure [638, 227, 856, 639, 599, 636, 637, 578, 853, 855, 577, 955]. Secured [532]. Security [641, 534, 848, 602, 934, 956, 300, 178, 295, 298, 433, 170, 940, 859, 426, 969, 532, 419, 958]. Seeds [223]. Seeker [696]. Seeon [938]. Segmentation [750, 500, 464, 468, 454, 473, 465]. Segmented [498]. Selected [947, 949]. Selecting [768]. Selection [762, 325, 683, 695, 501, 489, 122, 529, 739, 402, 94]. Selections [90]. Selective [597]. Self [775, 772, 218, 216, 221, 503, 195, 473]. Self- [473]. Self-adaptive [218, 221]. Self-organising [775]. Self-Organized [772]. Self-reconfigurable [216]. Self-reflection [195]. Selfish [605]. Semantic [459, 58, 671, 20, 864, 960, 868, 464, 664, 150, 906, 250, 738, 5, 921]. Semantics [109, 652, 102, 186, 838, 746]. Semi [253, 294, 923, 502, 153, 816]. Semi-autonomous [253]. Semi-Naïve [816]. Semi-structured [294]. Semi-Supervised [153, 923, 502]. Sense [906]. Sensemaking [24]. Sensitive [200]. Sensitivity [516]. Sensor [377, 530, 125, 375, 341, 342, 417, 306, 558, 336, 3]. Sensors [222, 305]. sensory [360]. Sentence [915, 909]. Sentences [131]. Sentiment [923]. Separating [610]. September [941, 933, 947, 928, 937, 967, 954, 926, 961, 964, 965, 936, 960, 934, 971, 956, 925, 966, 927, 929, 970, 948, 942, 951, 935, 931, 930, 932, 968, 972, 938, 969, 963, 958, 952, 962]. Sequence [223, 198]. Sequences [108, 319]. Sequential [492, 632, 491]. Series [790, 518]. Serious [403]. Server [545]. Servers [8, 55]. Service [537, 563, 41, 387, 36, 379, 38, 381]. Service-Based [563]. Service-Oriented [41]. Services [47, 577]. Session [80, 77]. Set [278, 89, 282, 737, 493, 658, 849]. Sets [804, 280, 345, 344, 131, 871, 488, 455, 898]. Setting [887, 115, 626]. Settings [858]. Seven [676]. Several [846]. Shape [468, 473, 435]. Shared [69, 477, 732, 707, 733]. Shared-Memory [732]. Sharing [374, 860]. Ship [350]. Short [530]. Short-Preamble [530]. Should [711]. SI [291]. Side [274]. Sided [90, 732, 721]. Sign [364, 633, 502, 751]. Signal [176, 26]. Signal-Based [176]. Signals [25]. Signature [630, 842, 843, 851]. Signatures [590, 631, 633, 632, 600]. Signcryption [844]. Silence [177]. Similar [770]. Similarities [263]. Similarity [510, 319, 682, 906, 578, 522, 473]. Simple [748, 856, 270, 741, 273]. Simulate [407]. Simulated [404]. Simulating [517]. Simulation [68, 616, 358, 418, 350]. Simulations [946, 436, 437]. Simulators [537, 385]. Simultaneous [480]. Single [541]. site [386]. Situated [650]. Situation [447, 154, 461]. Situations [517]. Size [743, 759, 824, 681]. Skeletal [614]. Skill [257]. SLAs [298]. Slicing [200]. SLoM [254]. Slovakia [939]. Small [587]. Smart [944, 144, 881, 871, 378, 182, 380, 372, 236, 376, 529, 379, 381, 382, 347, 426, 37]. Smart-Card [37]. Smart-M3 [380]. Smartphone [333]. Smartphones [302]. SMEs [46]. Smoothed [287]. SMS [908, 770, 303]. SMS-EMOA [770]. Snapshots [226]. Snippet [690]. SNS [48]. SoCCR [791]. Social [835, 947, 47, 161, 443, 415, 424, 786, 244, 48, 784, 143, 707, 834, 559, 413, 411, 419, 521, 796]. Socially [405, 85, 151]. Socially-Aware [405]. Socio [781]. Socio-technical [781]. Sociophysical [829]. Sodium [888]. Sodium-Potassium [888]. Soft [112, 486, 753]. Software [296, 793, 783, 295, 786, 440, 429]. Solution [492, 116, 491, 774, 436]. Solutions [14, 151, 328, 591, 37]. Solvable [280]. Solver [322]. Solving [284, 964, 965, 832]. Some [835, 114]. Sound [364, 872, 898]. Soundex [908]. Sounds [353]. Source [546]. Space [380, 446, 620, 379, 523]. Spaced [223]. Spaces [944, 531, 871]. SPARQL [665, 803]. Sparse [618, 766]. Sparsely [472]. Spatial [856, 317, 258, 256, 264, 516, 259, 938, 265, 522, 354, 514, 938]. Spatially [518]. Spatio [263, 830]. Spatio-temporal [263, 830]. Speaker [584]. species [437]. Specific [590, 324, 636]. Specification [685, 891, 399]. Specifications [890]. Specificity [22]. Specified [855]. Specifying [863, 76, 895]. Spectral [365]. Speculative [651]. Speech [438, 686, 434, 972, 129]. Speed [619, 551, 423, 242]. Spline [481]. Spread [92]. Spreadsheets [192]. Spring [945]. Squad [830]. Square [837]. SSGIP [945]. SSL [589, 174, 175]. SSL/TLS [174, 175]. St [944]. Stable [836]. Stage [436]. Standard [111, 630, 842, 844]. Standardized [22]. Start [161]. State [828, 72, 620, 503, 44]. Stateful [76]. States [891, 155, 826]. Static [188, 423, 935]. Statistical [306, 493]. Statistics [644, 873]. Status [698]. Stealing [621]. Steepest [745]. Steiner [286]. Step [743, 160]. Step-Size [743]. Stereo [445, 444]. Stereo-Analysis [445]. Stereo-Based [444]. Stereoscopic [340]. Stereotypical [261]. Stiffness [359]. Stigmergy [156, 96]. Stimulation [361]. Stimuli [261, 365]. Storage [570, 545, 4, 306, 591, 577]. Stories [137]. Strange [64]. Strange-but-True [64]. Strategic [649, 168]. Strategies [740, 509, 258, 156, 303, 88]. Strategy [121, 890, 70, 765, 86, 148, 653, 75, 524, 721]. Strategy-Proof [86]. Stream [7, 669]. Streaming [340, 544]. Streams [574, 817]. Strength [642]. String [600]. Strings [278]. Stroke [24]. Strokes [346]. Strong [588]. Strongly [857]. Structural [838]. Structure [459, 492, 630, 633, 765, 436]. Structure-Preserving [630, 633]. Structured [593, 294]. Structures [334, 117, 853, 414]. Structuring [551]. Students [688, 432]. Studies [53]. Study [537, 780, 32, 314, 767, 531, 399, 36, 215, 786, 515, 785, 681, 13, 595]. Style [627]. Subclasses [114]. Subgraph [277, 448]. Subjects [399]. Successful [715]. Successive [618]. Succinct [203]. Succinctly [72]. Suite [121, 730]. Sum [151]. Summarisation [147]. Summer [960]. Supervised [438, 153, 603, 923, 502]. Supplements [746]. Supply [46]. Support [704, 575, 316, 793, 41, 150, 783, 154, 499, 606]. Supported [778]. Supporting [241, 795, 343, 349, 576, 794, 353, 784, 46]. SURE [251]. Surface [249, 251, 460, 481]. Surface-Approximation [249]. Surfaces [358, 472, 481]. Surgeries [340]. Surveillance [447, 97]. Survey [375]. Sustainability [702]. Sustained [542]. SVMs [146]. Swarm [772, 212, 219, 18, 213, 218, 205, 215, 220, 207, 214, 776, 936]. Swarms [773, 209]. Sweden [943]. Swing [503]. Switching [627]. SWRL [292]. Symbolic [167, 201, 193]. Symmetric [834, 854]. Symposium [937, 934, 971, 935, 950, 939]. Synchronisation [830]. Synchronization [822]. Synchronizers [837]. Synchronous [884]. Synergies [57]. Synergy [483]. Synergy-Based [483]. Syntax [109]. Synthesis [70]. Synthesising [77]. System [45, 312, 14, 850, 32, 712, 56, 163, 496, 385, 761, 25, 323, 499, 689, 333, 388, 547, 17, 12, 88, 834, 33, 436, 525, 437, 797]. Systematically [893]. Systems [835, 613, 310, 301, 72, 863, 550, 439, 596, 757, 104, 959, 535, 925, 35, 297, 657, 441, 601, 810, 621, 159, 326, 164, 142, 927, 794, 970, 50, 852, 784, 940, 930, 932, 5, 195, 382, 727, 526, 21].

T [121]. T-Way [121]. Tableau [865]. Tables [331]. Tacit [565]. Tactile [356, 266, 362]. Tag [912]. Taggers [912]. Tagging [711, 679, 129]. Tagless [392]. Tags [343, 415, 430]. Tailoring [769]. Taipei [946]. Taiwan [946]. Tallinn [954]. Tamper [846]. Tamper-Evident [846]. Tangible [795]. Taormina [964, 965]. Target [495]. Targeted [229]. Targeting [50]. Task [507, 54, 86, 53, 831, 679, 214, 88, 217]. Task-Parallelism [54]. Tasks [841, 258, 47, 260, 784, 255, 622]. Taxonomic [327]. Taxonomy [293]. TC [951]. TCP [543]. TDD [384]. Teach [346]. Team [90]. Teamwork [407]. Technical [778, 882, 781]. Technique [122, 97]. Techniques [399, 161, 136, 363, 37]. Techno [540]. Techno-Economic [540]. Technologies [960, 709, 953]. Technology [928, 383, 926, 32, 355, 718, 47, 966, 441]. Teeth [500]. Teeth/Palate [500]. Telecom [537]. Telephony [387]. Temperature [828]. Templates [562, 566]. Temporal [886, 757, 317, 870, 487, 895, 899, 263, 830, 697]. Tender [33]. Term [684]. Termination [67]. Test [45, 121, 680, 890, 691]. Testbed [528]. Testing [166, 536, 889]. Text [704, 303, 908, 972]. Texts [147]. Textual [703, 128]. Texture [158]. Thailand [32]. Theft [236]. Their [820, 716, 258, 688, 494]. Theoretic [656, 637, 813]. Theoretical [949]. Theory [78, 287, 138, 929, 62, 497, 106, 348, 962]. Thermal [144, 361, 291]. Thermodynamic [222]. Thesaurus [682]. Things [944, 369]. Third [926, 961, 174, 425, 948]. Third-Parties [174]. Third-Party [425]. Thirteen [885]. Threat [229]. Three [390, 854]. Threshold [835]. Thresholding [766]. Thresholds [365]. throughput [22]. Tier [726]. Time [740, 277, 241, 416, 790, 240, 335, 822, 116, 464, 293, 319, 894, 298, 636, 371, 333, 696, 547, 650, 518, 340, 751, 73, 899]. Time-Oriented [293]. Time-Situated [650]. Time-Specific [636]. Time-Triggered [547]. Timed [281, 69]. Timestep [822]. TinyStream [305]. Tipple [699]. Tokens [629]. Tolerant [812]. Tool [914, 729, 913]. Toolbox [244, 57]. Toolkit [271]. Tools [841, 808, 57, 606]. Top [93]. Top-Down [93]. Topic [611, 681, 606, 617]. Topological [820, 512]. Topologies [656]. Topology [247, 773, 763, 654]. Topology-Aware [654]. Totalistic [820, 836]. touch [331]. TPDL [962]. Trabecular [499]. Tracing [562]. Tracking [166, 335, 17]. Trade [490]. Trade-Offs [490]. Tradeoff [570]. Trading [761]. Traffic [502, 156, 96, 461]. Training [356, 779, 146, 502]. Trait [301]. Trajectories [740, 510]. Transaction [1]. Transducers [917]. TransElas [1]. Transfer [79, 753]. Transferring [416]. Transform [479, 446]. Transformation [596]. Transformations [537, 896]. Transition [155]. Transition-Based [155]. Translational [23]. Transmission [529]. Transport [425]. Transportation [274]. Traversal [562]. Treatment [312, 323, 17]. Tree [286, 102, 917, 826]. Treebank [917, 909, 913]. Trees [416]. Trend [417]. Trends [933]. Trento [951]. Triage [696]. Triangular [839]. Triggered [699, 547]. Trip [507]. Triple [245]. Triple-Channel [245]. True [64]. Trust [301, 174, 601, 789, 580, 525, 94]. Trusted [226]. Trustworthiness [228]. Trustworthy [594, 581]. TSD [972]. Tubular [472]. Tulipse [607]. Tuning [755]. Tuple [807]. Tuple-Generating [807]. tuples [754]. Turing [100, 61, 63]. Turkey [955]. Tutorial [800, 254]. TV [467]. TV-L1 [467]. Two [600, 90, 152, 726, 160, 637, 136, 832, 542, 839]. Two-Color [839]. Two-Dimensional [832, 839]. Two-Move [600]. Two-Party [637]. Two-Player [152]. Two-Sided [90]. Two-Step [160]. Two-Tier [726]. Tying [551]. Tyne [929]. Type [610, 131, 359]. Type-Based [610]. Typed [109, 392]. Types [192, 80, 77, 48, 394]. Typology [262].

U.S [503]. Ubiquitous [947]. UC [848]. UESDE [849]. UHF [531]. UIs [419]. UK [945, 929]. UML [299]. Unbounded [849]. Uncertain [668]. Uncertainty [344]. Uncover [414]. Underground [431]. Understand [373]. Understanding [613, 614, 786, 352, 229]. Underwater [452]. uniform [621]. Unifying [870]. Unique [475, 38]. Unit [97]. Units [144]. Universal [219]. Universally [641]. Unordered [276]. Unrelated [771]. Unsegmented [686]. Unsupervised [485]. Untrusted [226]. Update [869]. Updates [787]. Updating [476]. Uplink [385]. upon [929]. Upper [108]. Urban [349, 250]. Uruguayan [41]. USA [947, 952]. Usability [399, 124, 37]. Usage [601, 299]. Usage-Constrained [601]. Use [553, 22, 299, 697, 704]. Useable [302]. User [307, 180, 795, 125, 705, 316, 878, 124, 366, 297, 325, 359, 424, 565, 418, 701, 963, 525, 607, 37, 252, 331, 788]. User-Centric [125]. User-Defined [252]. User-Experience [366]. User-Friendly [37]. User-Generated [705, 316]. User-Guided [607]. Users [782, 694]. Uses [105]. Using [121, 137, 313, 556, 887, 644, 846, 888, 9, 618, 533, 116, 476, 500, 342, 809, 15, 682, 479, 83, 47, 26, 321, 254, 189, 154, 337, 810, 131, 621, 291, 917, 689, 789, 415, 614, 248, 825, 122, 303, 503, 763, 654, 814, 497, 195, 493, 472, 752, 259, 855, 538, 658, 566, 373, 916]. Utility [8].

Vadis [645]. Vague [519]. Validation [231, 292]. Value [115, 116].

Vandalism [708]. Variable [749, 481]. Variance [755]. Variants [490]. Varied [858]. VASS [65]. Vault [705]. Vector [511, 513, 486, 154]. Vectors [890, 495, 91]. Vehicle [220]. Verbal [266]. Verbose [684]. Verbs [262]. Verification [887, 897, 729]. Verifying [895]. Versatile [884]. Versions [703]. versus [749]. via [196, 886, 877, 113, 639, 167, 637, 443, 217]. Victim [796]. Video [447, 148, 387, 466, 544, 504]. Vienna [928, 967, 926, 960, 925, 927, 970, 963]. VietNam [33]. View [830, 802, 478]. Viewpoint [475]. VIII [938]. Virtual [401, 227, 404, 779, 399, 783, 124, 138, 400, 266, 358, 362, 532, 624, 406, 347, 398]. Virtualized [6]. Vision [949]. Visual [961, 256, 696]. Visualization [20, 339, 942, 52, 607]. Visualizing [157]. Visually [154]. Visuo [264]. Visuo-spatial [264]. VLDB [955]. VM [226]. Voting [45, 878, 503]. Voyage [185]. VPNs [177]. VR [378, 255]. VR-Smart [378]. vs [530]. Vulnerability [427].

Wait [733]. Wait-Free [733]. Walking [358, 520]. Wallet [182]. Wallet-Friendly [182]. WAN [613]. Warehouses [519]. Warming [928]. Way [121, 342, 129]. Weak [846, 66]. Wearable [795]. Web [960, 338, 579, 705, 316, 146, 878, 513, 664, 916, 791, 157, 134, 186, 970, 672]. Web-Based [791]. Web2.0 [1]. Webpage [430]. Websites [237]. Weight [210]. Weighted [416, 68]. Weighting [56, 323, 326, 684]. Welfare [521]. Well [57, 803, 108, 470]. Well-Established [57]. WG [940, 953]. Wheelchair [157]. Where [373, 374]. Which [697]. White [531]. Who [711]. Whose [298]. Wide [613, 455, 44]. Width [439]. Wikipedia [708]. Wikis [671]. WiMAX [542]. Wireless [377, 530, 375, 224, 342, 179, 387, 529, 558]. within [139]. without [174, 599, 95, 857]. Witness [94]. Wolf [241]. Wolfram [827]. WoLLIC [931]. Word [906]. Words [697]. Work [621, 788]. Work-Stealing [621]. Worked [44]. Workers [795]. Workflow [332, 169, 329]. Workflows [311, 895]. Working [344]. Workload [616]. WorkMail [332]. Workplace [700]. Workshop [967, 946, 949, 959, 955, 951, 931, 940, 932]. Workshops [947]. World [458, 949, 445]. Would [694]. WSN [592, 172].

X [478]. X-Ray [478]. X.509 [175]. XE [727]. XE/XK [727]. XII [964, 965]. XIII [959]. XML [292].

Years [676, 885].

Zero [628, 160]. Zero-Knowledge [628]. Zone [499]. Zones [192]. Zonotopes [188].

References

Gueye:2012:TET

[1] Ibrahima Gueye, Idrissa Sarr, and Hubert Naacke. TransElas: Elastic transaction monitoring for Web2.0 applications. Lecture Notes in Computer Science, 7450:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32344-7_1/.

Jajodia:2012:RET

[2] Sushil Jajodia, Witold Litwin, and Thomas Schwarz. Recoverable encryption through noised secret over a large cloud. Lecture Notes in Computer Science, 7450:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32344-7_2/.

Yu:2012:BMB

[3] Byunggu Yu, Alfredo Cuzzocrea, Dong Jeong, and Sergey Maydebura. A Bigtable/MapReduce-based cloud infrastructure for effectively and efficiently managing large-scale sensor networks. Lecture Notes in Computer Science, 7450:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32344-7_3/.

Moise:2012:MAC

[4] Diana Moise and Alexandra Carpen-Amarie. MapReduce applications in the cloud: a cost evaluation of computation and storage. Lecture Notes in Computer Science, 7450:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32344-7_4/.

Sakka:2012:SID

[5] Mohamed Amin Sakka and Bruno Defude. Scalability issues in designing and implementing semantic provenance management systems. Lecture Notes in Computer Science, 7450:49–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32344-7_5/.

Sodhi:2012:PCV

[6] Balwinder Sodhi and T. V. Prabhakar. Performance characteristics of virtualized platforms from applications perspective. Lecture Notes in Computer Science, 7450:62–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32344-7_6/.

Chen:2012:OEE

[7] Qiming Chen and Meichun Hsu. Open execution engines of stream analysis operations. Lecture Notes in Computer Science, 7450:75–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32344-7_7/.

Pallis:2012:ICD

[8] George Pallis. Improving content delivery by exploiting the utility of CDN servers. Lecture Notes in Computer Science, 7450:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32344-7_8/.

Cesario:2012:UMD

[9] Eugenio Cesario, Carlo Mastroianni, and Domenico Talia. Using mining@home for distributed ensemble learning. Lecture Notes in Computer Science, 7450:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32344-7_9/.

Anonymous:2012:BMa

[10] Anonymous. Back matter. Lecture Notes in Computer Science, 7450: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32344-7/1.

Anonymous:2012:FMa

[11] Anonymous. Front matter. Lecture Notes in Computer Science, 7450: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32344-7/1.

Portela:2012:IDA

[12] Filipe Portela, Manuel Filipe Santos, José Machado, Álvaro Silva, and Fernando Rua. Intelligent data acquisition and scoring system for intensive medicine. Lecture Notes in Computer Science, 7451:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32395-9_1/.

Santos:2012:DMS

[13] Maribel Yasmina Santos, Jorge Cruz, and Artur Teles de Araújo. Data mining in the study of the chronic obstructive pulmonary disease. Lecture Notes in Computer Science, 7451:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32395-9_2/.

Bursa:2012:PPS

[14] Miroslav Bursa, Lenka Lhotska, Vaclav Chudacek, Jiri Spilka, Petr Janku, and Martin Huser. Practical problems and solutions in hospital information system data mining. Lecture Notes in Computer Science, 7451: 31–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32395-9_3/.

Girardi:2012:UGM

[15] Dominic Girardi, Michael Giretzlehner, and Josef Küng. Using generic meta-data-models for clustering medical data. Lecture Notes in Computer Science, 7451:40–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32395-9_4/.

Santos-Pereira:2012:MBA

[16] Cátia Santos-Pereira, Alexandre B. Augusto, Manuel E. Correia, and Ana Ferreira. A mobile based authorization mechanism for patient managed role based access control. Lecture Notes in Computer Science, 7451: 54–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32395-9_5/.

Paschou:2012:CMM

[17] Mersini Paschou, Efrosini Sourla, George Basagiannis, and Evangelos Sakkopoulos. Care@HOME: a mobile monitoring system for patient treatment and blood pressure tracking. Lecture Notes in Computer Science, 7451:69–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32395-9_6/.

Hristoskova:2012:ICA

[18] Anna Hristoskova, Veselka Boeva, and Elena Tsiporkova. An integrative clustering approach combining particle swarm optimization and formal concept analysis. Lecture Notes in Computer Science, 7451:84–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32395-9_7/.

Folino:2012:LPA

[19] Francesco Folino and Clara Pizzuti. Link prediction approaches for disease networks. Lecture Notes in Computer Science, 7451:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32395-9_8/.

Ceci:2012:TSF

[20] Michelangelo Ceci, Fabio Fumarola, Pietro Hiram Guzzi, and Federica Mandreoli. Toward a semantic framework for the querying, mining and visualization of cancer microenvironment data. Lecture Notes in Computer Science, 7451:109–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32395-9_9/.

Wyner:2012:ARR

[21] Adam Wyner, Luke Riley, Robert Hoehndorf, and Samuel Croset. Argumentation to represent and reason over biological systems. Lecture Notes in Computer Science, 7451:124–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32395-9_10/.

Kozak:2012:UDS

[22] Karol Kozak. The use of design specificity in standardized mean difference for analysis of high throughput RNA interference screens. Lecture Notes in Computer Science, 7451:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32395-9_11/.

Machado:2012:TTM

[23] Catia M. Machado, Francisco M. Couto, Alexandra R. Fernandes, and Susana Santos. Toward a translational medicine approach for hypertrophic cardiomyopathy. Lecture Notes in Computer Science, 7451: 151–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32395-9_12/.

Holzinger:2012:CSE

[24] Andreas Holzinger, Reinhold Scherer, Martin Seeber, Johanna Wagner, and Gernot Müller-Putz. Computational sensemaking on examples of knowledge discovery from neuroscience data: Towards enhancing stroke rehabilitation. Lecture Notes in Computer Science, 7451: 166–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32395-9_13/.

Havlik:2012:DCS

[25] Jan Havlík, Lucie Kučerová, Imrich Kohút, Jan Dvořák, and Vratislav Fabián. The database of the cardiovascular system related signals. Lecture Notes in Computer Science, 7451:169–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-32395-9_14.

Havlik:2012:PMU

[26] Jan Havlík, Ondřej Fousek, and Miroslav Ložek. Patient monitoring using bioimpedance signal. Lecture Notes in Computer Science, 7451:171–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32395-9_15.

Anonymous:2012:BMb

[27] Anonymous. Back matter. Lecture Notes in Computer Science, 7451: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32395-9/1.

Anonymous:2012:FMb

[28] Anonymous. Front matter. Lecture Notes in Computer Science, 7451: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32395-9/1.

Traunmuller:2012:KMG

[29] Roland Traunmüller. Knowledge management in government: New perspectives. Lecture Notes in Computer Science, 7452:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32701-8_1/.

Prosser:2012:EDW

[30] Alexander Prosser. eParticipation — did we deliver what we promised? Lecture Notes in Computer Science, 7452:10–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32701-8_2/.

Nyakwende:2012:FAD

[31] Emmanuel Nyakwende and Ali Al Mazari. Factors affecting the development of e-government in Saudi Arabia. Lecture Notes in Computer Science, 7452:19–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32701-8_3/.

Chutimaskul:2012:FMD

[32] Wichian Chutimaskul and Suree Funilkul. The framework for monitoring the development process and inspection of government information system and technology: a case study in Thailand. Lecture Notes in Computer Science, 7452:29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32701-8_4/.

Tuan:2012:OTE

[33] La Anh Tuan and John Debenham. Online tender evaluation: VietNam government e-Procurement system. Lecture Notes in Computer Science, 7452:44–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32701-8_5/.

Strykowski:2012:CME

[34] Sergiusz Strykowski and Rafał Wojciechowski. Composable modeling and execution of administrative procedures. Lecture Notes in Computer Science, 7452:52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32701-8_6/.

Hawryszkiewycz:2012:DFC

[35] Igor T. Hawryszkiewycz. A design framework for collaboration in systems of systems. Lecture Notes in Computer Science, 7452:67–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32701-8_7/.

Hsiao:2012:SFI

[36] Chih-Hui Hsiao, Hui-Chih Wang, and Her-Sen Doong. A study of factors influencing e-government service acceptance intention: a multiple perspective approach. Lecture Notes in Computer Science, 7452: 79–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32701-8_8/.

Zefferer:2012:TUF

[37] Thomas Zefferer and Vesna Krnjic. Towards user-friendly e-government solutions: Usability evaluation of Austrian Smart-Card integration techniques. Lecture Notes in Computer Science, 7452:88–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32701-8_9/.

Shrivastava:2012:UIE

[38] Swapnil Shrivastava, Zia Saquib, Gopinath P., and Peeyush Chomal. Unique identity enabled service delivery through NSDG. Lecture Notes in Computer Science, 7452:103–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32701-8_10/.

Carota:2012:FFI

[39] Serenella Carota, Flavio Corradini, Damiano Falcioni, and Maria Laura Maggiulli. FedCohesion: Federated identity management in the Marche region. Lecture Notes in Computer Science, 7452:112–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32701-8_11/.

Falcioni:2012:LDD

[40] Damiano Falcioni, Andrea Polini, Alberto Polzonetti, and Barbara Re. Livelock and deadlock detection for PA inter-organizational business processes. Lecture Notes in Computer Science, 7452:125–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32701-8_12/.

Gonzalez:2012:SOI

[41] Laura González, Raúl Ruggia, Jorge Abin, Guzmán Llambías, Raquel Sosa, and Bruno Rienzi. A service-oriented integration platform to support a joined-up E-government approach: The Uruguayan experience. Lecture Notes in Computer Science, 7452:140–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32701-8_13/.

Munoz:2012:MAP

[42] José Félix Muñoz, Carlos Bobed, and Francisco Serón. Modeling administrative procedures to improve information to the public. Lecture Notes in Computer Science, 7452:155–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32701-8_14/.

Debenham:2012:CCM

[43] John Debenham and Simeon Simoff. The co-creation machine: Managing co-creative processes for the crowd. Lecture Notes in Computer Science, 7452:170–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32701-8_15/.

Sousa:2012:WPE

[44] Artur Afonso Sousa, Pedro Agante, and Luís Borges Gouveia. A worked proposal on eParticipation for state wide elections. Lecture Notes in Computer Science, 7452:178–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32701-8_16/.

Ablameyko:2012:TIG

[45] Sergey Ablameyko, Nikolai Kalosha, Sergey Bratchenya, and Vitali Lipen. Test implementation of the “guarantor” Internet voting system: Design principles, operation, and approbation. Lecture Notes in Computer Science, 7452:191–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32701-8_17/.

Ternai:2012:EPS

[46] Katalin Ternai and Ildikó Szabó. EBEST platform supporting SMEs to manage supply chain and collaborate. Lecture Notes in Computer Science, 7452:202–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32701-8_18/.

Hatzi:2012:USN

[47] Ourania Hatzi, Mara Nikolaidou, Panagiotis Katsivelis-Perakis, and Valentino Hudhra. Using social network technology to provide eadministration services as collaborative tasks. Lecture Notes in Computer Science, 7452:216–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32701-8_19/.

Park:2012:DDT

[48] Sangin Park, Yeontai Choi, Dukyun Hwang, and Miyon Paek. Do different types of SNS have different impacts on participatory social capital? Lecture Notes in Computer Science, 7452:231–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32701-8_20/.

Anonymous:2012:FMc

[49] Anonymous. Front matter. Lecture Notes in Computer Science, 7452: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32701-8/1.

Matsumura:2012:ICC

[50] Tadayuki Matsumura and Tetsuya Yamada. IT-cooling collaborative control methods for battery-aware IT-systems targeting India. Lecture Notes in Computer Science, 7453:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32606-6_1/.

Denguir:2012:RID

[51] Afef Denguir, François Trousset, and Jacky Montmain. Research for IT driven energy efficiency based on a multidimensional comfort control^{∗}. Lecture Notes in Computer Science, 7453:18–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32606-6_2/.

Polson:2012:EGB

[52] Deb Polson and Cassandra Selin. The ECOS Green buildings project: Data dramatization, visualization and manipulation. Lecture Notes in Computer Science, 7453:33–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32606-6_3/.

Lien:2012:CSM

[53] Hallgeir Lien, Lasse Natvig, Abdullah Al Hasib, and Jan Christian Meyer. Case studies of multi-core energy efficiency in task based programs. Lecture Notes in Computer Science, 7453:44–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32606-6_4/.

Aliaga:2012:LTP

[54] José I. Aliaga, Manuel F. Dolz, Alberto F. Martín, and Rafael Mayo. Leveraging task-parallelism in energy-efficient ILU preconditioners. Lecture Notes in Computer Science, 7453:55–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32606-6_5/.

ulIslam:2012:EEC

[55] Saif ul Islam and Jean-Marc Pierson. Evaluating energy consumption in CDN servers. Lecture Notes in Computer Science, 7453:64–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32606-6_6/.

Ferreira:2012:GPI

[56] Alexandre Mello Ferreira, Barbara Pernici, and Pierluigi Plebani. Green performance indicators aggregation through composed weighting system. Lecture Notes in Computer Science, 7453:79–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32606-6_7/.

Nguyen:2012:EIA

[57] Thanh Binh Nguyen, Fabian Wagner, and Wolfgang Schoepp. EC4MACS — an integrated assessment toolbox of well-established modeling tools to explore the synergies and interactions between climate change, air quality and other policy objectives. Lecture Notes in Computer Science, 7453: 94–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32606-6_8/.

Bertin:2012:SMD

[58] Benjamin Bertin, Marian Scuturici, Jean-Marie Pinon, and Emmanuel Risler. Semantic modelling of dependency relations between life cycle analysis processes. Lecture Notes in Computer Science, 7453:109– 124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32606-6_9/.

Anonymous:2012:BMc

[59] Anonymous. Back matter. Lecture Notes in Computer Science, 7453: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32606-6/1.

Anonymous:2012:FMd

[60] Anonymous. Front matter. Lecture Notes in Computer Science, 7453: I–X, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-32606-6/1.

Baeten:2012:TMM

[61] Jos C. M. Baeten, Bas Luttik, and Paul van Tilburg. Turing meets Milner. Lecture Notes in Computer Science, 7454:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32940-1_1/.

Plotkin:2012:CAT

[62] Gordon D. Plotkin. Concurrency and the algebraic theory of effects. Lecture Notes in Computer Science, 7454:21–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-32940-1_2.

Randell:2012:TE

[63] Brian Randell. A Turing enigma. Lecture Notes in Computer Science, 7454:23–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32940-1_3/.

Sewell:2012:FCS

[64] Peter Sewell. False concurrency and strange-but-true machines. Lecture Notes in Computer Science, 7454:37–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-32940-1_4.

Berard:2012:CGV

[65] Béatrice Bérard, Serge Haddad, Mathieu Sassolas, and Nathalie Sznajder. Concurrent games on VASS with inhibition. Lecture Notes in Computer Science, 7454:39–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32940-1_5/.

Czerwinski:2012:RPW

[66] Wojciech Czerwiński, Piotr Hofman, and Sławomir Lasota. Reachability problem for weak multi-pushdown automata. Lecture Notes in Computer Science, 7454:53–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32940-1_6/.

Yu:2012:RTA

[67] Nengkun Yu and Mingsheng Ying. Reachability and termination analysis of concurrent quantum programs. Lecture Notes in Computer Science, 7454:69–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32940-1_7/.

Avni:2012:MWC

[68] Guy Avni and Orna Kupferman. Making weighted containment feasible: a heuristic based on simulation and abstraction. Lecture Notes in Computer Science, 7454:84–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32940-1_8/.

Balaguer:2012:ASC

[69] Sandie Balaguer and Thomas Chatain. Avoiding shared clocks in networks of timed automata. Lecture Notes in Computer Science, 7454: 100–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32940-1_9/.

Chatterjee:2012:SSM

[70] Krishnendu Chatterjee, Mickael Randour, and Jean-François Raskin. Strategy synthesis for multi-dimensional quantitative objectives. Lecture Notes in Computer Science, 7454:115–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32940-1_10/.

Filiot:2012:QLD

[71] Emmanuel Filiot, Raffaella Gentilini, and Jean-François Raskin. Quantitative languages defined by functional automata. Lecture Notes in Computer Science, 7454:132–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32940-1_11/.

Brenguier:2012:CSR

[72] Romain Brenguier, Stefan Göller, and Ocan Sankur. A comparison of succinctly represented finite-state systems. Lecture Notes in Computer Science, 7454:147–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32940-1_12/.

Valmari:2012:ALT

[73] Antti Valmari. All linear-time congruences for familiar operators part 2: Infinite LTSs. Lecture Notes in Computer Science, 7454:162–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32940-1_13/.

DaCosta:2012:QCE

[74] Arnaud Da Costa, François Laroussinie, and Nicolas Markey. Quantified CTL: Expressiveness and model checking. Lecture Notes in Computer Science, 7454:177–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32940-1_14/.

Mogavero:2012:WMA

[75] Fabio Mogavero, Aniello Murano, Giuseppe Perelli, and Moshe Y. Vardi. What makes Atl* decidable? A decidable fragment of strategy logic. Lecture Notes in Computer Science, 7454:193–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32940-1_15/.

Chen:2012:SSA

[76] Tzu-Chun Chen and Kohei Honda. Specifying stateful asynchronous properties for distributed programs. Lecture Notes in Computer Science, 7454: 209–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32940-1_16/.

Lange:2012:SCL

[77] Julien Lange and Emilio Tuosto. Synthesising choreographies from local session types. Lecture Notes in Computer Science, 7454:225–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32940-1_17/.

Aarts:2012:THD

[78] Fides Aarts, Faranak Heidarian, and Frits Vaandrager. A theory of history dependent abstractions for learning interface automata. Lecture Notes in Computer Science, 7454:240–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32940-1_18/.

Gotsman:2012:LOT

[79] Alexey Gotsman and Hongseok Yang. Linearizability with ownership transfer. Lecture Notes in Computer Science, 7454:256–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32940-1_19/.

Demangeon:2012:NPS

[80] Romain Demangeon and Kohei Honda. Nested protocols in session types. Lecture Notes in Computer Science, 7454:272–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32940-1_20/.

Anonymous:2012:FMe

[81] Anonymous. Front matter. Lecture Notes in Computer Science, 7454: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32940-1/1.

AAgotnes:2012:CPA

[82] Thomas Ågotnes and Hans van Ditmarsch. Coalitional public announcement games. Lecture Notes in Computer Science, 7455:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32729-2_1/.

Hao:2012:PMC

[83] Jianye Hao, Songzheng Song, Yang Liu, Jun Sun, Lin Gui, Jin Song Dong, and Ho fung Leung. Probabilistic model checking multi-agent behaviors in dispersion games using counter abstraction. Lecture Notes in Computer Science, 7455:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32729-2_2/.

Hindriks:2012:DE

[84] Koen V. Hindriks. Debugging is explaining. Lecture Notes in Computer Science, 7455:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32729-2_3/.

Hao:2012:ENP

[85] Jianye Hao and Ho fung Leung. An efficient negotiation protocol to achieve socially optimal allocation. Lecture Notes in Computer Science, 7455:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32729-2_4/.

Ghoneim:2012:RBS

[86] Ayman Ghoneim. Reassignment-based strategy-proof mechanisms for interdependent task allocation. Lecture Notes in Computer Science, 7455:

61–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32729-2_5/.

Zhang:2012:MRP

[87] Yingqian Zhang and Sicco Verwer. Mechanism for robust procurements. Lecture Notes in Computer Science, 7455:77–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32729-2_6/.

Poulet:2012:ABS

[88] Cyril Poulet, Vincent Corruble, and Amal El Fallah Seghrouchni. Auctionbased strategies for the open-system patrolling task. Lecture Notes in Computer Science, 7455:92–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32729-2_7/.

Anders:2012:DMA

[89] Gerrit Anders, Florian Siefert, Jan-Philipp Steghöfer, and Wolfgang Reif. A decentralized multi-agent algorithm for the set partitioning problem. Lecture Notes in Computer Science, 7455:107–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32729-2_8/.

Hamada:2012:TSP

[90] Dai Hamada and Toshiharu Sugawara. Two-sided parameter learning of role selections for efficient team formation. Lecture Notes in Computer Science, 7455:122–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32729-2_9/.

Matsui:2012:DSM

[91] Toshihiro Matsui, Marius Silaghi, Katsutoshi Hirayama, Makoto Yokoo, and Hiroshi Matsuo. Distributed search method with bounded cost vectors on multiple objective DCOPs. Lecture Notes in Computer Science, 7455: 137–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32729-2_10/.

Hafizoglu:2012:AOS

[92] Feyza Merve Hafızoğlu and Sandip Sen. Analysis of opinion spread through migration and adoption in agent communities. Lecture Notes in Computer Science, 7455:153–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32729-2_11/.

Savarimuthu:2012:GNP

[93] Bastin Tony Roy Savarimuthu, Lam-Son Lê, and Aditya Ghose. From green norms to policies — combining bottom-up and topdown approaches. Lecture Notes in Computer Science, 7455:168–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32729-2_12/.

Yu:2012:AWS

[94] Han Yu, Zhiqi Shen, and Bo An. An adaptive witness selection method for reputation-based trust models. Lecture Notes in Computer Science, 7455: 184–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32729-2_13/.

Pitt:2012:PAC

[95] Jeremy Pitt and Julia Schaumeier. Provision and appropriation of common-pool resources without full disclosure. Lecture Notes in Computer Science, 7455:199–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32729-2_14/.

Ito:2012:EAS

[96] Takayuki Ito, Ryo Kanamori, Jun Takahashi, Ivan Marsa Maestre, and Enrique de la Hoz. Effect of anticipatory stigmergy on decentralized traffic congestion control. Lecture Notes in Computer Science, 7455: 214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32729-2_15/.

Truong:2012:MSN

[97] Viet Xuan Truong, Hiep Xuan Huynh, Minh Ngoc Le, and Alexis Drogoul. Modeling a surveillance network based on unit disk graph technique — application for monitoring the invasion of insects in mekong delta region. Lecture Notes in Computer Science, 7455:228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32729-2_16/.

Anonymous:2012:BMd

[98] Anonymous. Back matter. Lecture Notes in Computer Science, 7455: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32729-2/1.

Anonymous:2012:FMf

[99] Anonymous. Front matter. Lecture Notes in Computer Science, 7455: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32729-2/1.

Asperti:2012:FTM

[100] Andrea Asperti and Wilmer Ricciotti. Formalizing Turing machines. Lecture Notes in Computer Science, 7456:1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32621-9_1/.

Fokina:2012:ERC

[101] Ekaterina Fokina, Sy Friedman, and André Nies. Equivalence relations that are Σ_{3}^{0} complete for computable reducibility. Lecture Notes in Computer Science, 7456:26–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32621-9_2/.

Kallmeyer:2012:ADM

[102] Laura Kallmeyer and Rainer Osswald. An analysis of directed motion expressions with lexicalized tree adjoining grammars and frame semantics. Lecture Notes in Computer Science, 7456:34–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32621-9_3/.

Metcalfe:2012:ARC

[103] George Metcalfe. Admissible rules: From characterizations to applications. Lecture Notes in Computer Science, 7456:56–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32621-9_4/.

Diekert:2012:DMA

[104] Volker Diekert and Anca Muscholl. On distributed monitoring of asynchronous systems. Lecture Notes in Computer Science, 7456:70–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32621-9_5/.

Schweikardt:2012:EPL

[105] Nicole Schweikardt. On the expressive power of logics with invariant uses of arithmetic predicates. Lecture Notes in Computer Science, 7456: 85–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32621-9_6/.

Selinger:2012:LMQ

[106] Peter Selinger. Logical methods in quantum information theory. Lecture Notes in Computer Science, 7456:88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-32621-9_7.

vanDitmarsch:2012:QN

[107] Hans van Ditmarsch. Quantifying notes. Lecture Notes in Computer Science, 7456:89–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32621-9_8/.

Abriola:2012:LBS

[108] Sergio Abriola, Santiago Figueira, and Gabriel Senno. Linearizing bad sequences: Upper bounds for the product and majoring Well quasiorders. Lecture Notes in Computer Science, 7456:110–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32621-9_9/.

Ahrens:2012:ITS

[109] Benedikt Ahrens. Initiality for typed syntax and semantics. Lecture Notes in Computer Science, 7456:127–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32621-9_10/.

Areces:2012:MAF

[110] Carlos Areces, Raul Fervari, and Guillaume Hoffmann. Moving arrows and four model checking results. Lecture Notes in Computer Science, 7456: 142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32621-9_11/.

Baldi:2012:SCE

[111] Paolo Baldi, Agata Ciabattoni, and Lara Spendier. Standard completeness for extensions of MTL: An automated approach. Lecture Notes in Computer Science, 7456:154–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32621-9_12/.

Baltag:2012:LJB

[112] Alexandru Baltag, Bryan Renne, and Sonja Smets. The logic of justified belief change, soft evidence and defeasible knowledge. Lecture Notes in Computer Science, 7456:168–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32621-9_13/.

Bezhanishvili:2012:MD

[113] Nick Bezhanishvili, Clemens Kupke, and Prakash Panangaden. Minimization via duality. Lecture Notes in Computer Science, 7456:191– 205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32621-9_14/.

Callejas:2012:SSF

[114] Claudio Callejas, João Marcos, and Benjamín René Callejas Bedregal. On some subclasses of the Fodor–Roubens fuzzy bi-implication. Lecture Notes in Computer Science, 7456:206–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32621-9_15/.

Diaz-Caro:2012:LND

[115] Alejandro Díaz-Caro and Barbara Petit. Linearity in the non-deterministic call-by-value setting. Lecture Notes in Computer Science, 7456:216– 231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32621-9_16/.

Farjudian:2012:PTS

[116] Amin Farjudian. Polynomial-time solution of initial value problems using polynomial enclosures. Lecture Notes in Computer Science, 7456: 232–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32621-9_17/.

Fouche:2012:ARR

[117] Willem L. Fouché. Algorithmic randomness and Ramsey properties of countable homogeneous structures. Lecture Notes in Computer Science, 7456:246–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32621-9_18/.

Link:2012:PRA

[118] Sebastian Link. Propositional reasoning about Saturated conditional probabilistic independence. Lecture Notes in Computer Science, 7456: 257–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32621-9_19/.

Ribeiro:2012:CL

[119] Márcio M. Ribeiro and Marcelo E. Coniglio. Contracting logics. Lecture Notes in Computer Science, 7456:268–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32621-9_20/.

Anonymous:2012:FMg

[120] Anonymous. Front matter. Lecture Notes in Computer Science, 7456: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32621-9/1.

Al-Sewari:2012:CDW

[121] AbdulRahman A. Al-Sewari and Kamal Z. Zamli. Constraints dependent T-way test suite generation using harmony search strategy. Lecture Notes in Computer Science, 7457:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32541-0_1/.

Othman:2012:EDM

[122] Siti Hajar Othman and Ghassan Beydoun. Evaluating disaster management knowledge model by using a frequency-based selection technique. Lecture Notes in Computer Science, 7457:12–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32541-0_2/.

An:2012:PEA

[123] Fengqi An, Xiongcai Cai, and Arcot Sowmya. Perceptual evaluation of automatic 2.5D cartoon modelling. Lecture Notes in Computer Science, 7457:28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32541-0_3/.

Hanna:2012:AAU

[124] Nader Hanna, Deborah Richards, and Michael J. Jacobson. Automatic acquisition of user models of interaction to evaluate the usability of virtual environments. Lecture Notes in Computer Science, 7457: 43–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32541-0_4/.

Chen:2012:UCR

[125] Zhao Chen, Md. Sumon Shahriar, and Byeong Ho Kang. Usercentric recommendation-based approximate information retrieval from marine sensor data. Lecture Notes in Computer Science, 7457:58–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32541-0_5/.

Echeverry:2012:ACK

[126] Ana Ximena Halabi Echeverry and Deborah Richards. Addressing challenges for knowledge discovery from data in the domain of seaport integration. Lecture Notes in Computer Science, 7457:73–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32541-0_6/.

Kuah:2012:DEA

[127] Chuen Tse Kuah and Kuan Yew Wong. Data envelopment analysis for evaluating knowledge acquisition and creation. Lecture Notes in Computer Science, 7457:86–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32541-0_7/.

Ho:2012:HOH

[128] Dung T. Ho and Tru H. Cao. A high-order hidden Markov model for emotion detection from textual data. Lecture Notes in Computer Science, 7457:94–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32541-0_8/.

Zamin:2012:LMW

[129] Norshuhani Zamin, Alan Oxley, Zainab Abu Bakar, and Syed Ahmad Farhan. A lazy man’s way to part-of-speech tagging. Lecture Notes in Computer Science, 7457:106–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32541-0_9/.

Galgani:2012:KAC

[130] Filippo Galgani, Paul Compton, and Achim Hoffmann. Knowledge acquisition for categorization of legal case reports. Lecture Notes in Computer Science, 7457:118–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32541-0_10/.

Ishikawa:2012:EHT

[131] Kyohei Ishikawa and Hayato Ohwada. Extraction of how-to type questionanswering sentences using query sets. Lecture Notes in Computer Science, 7457:133–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32541-0_11/.

Boulemden:2012:IIR

[132] Ahmed Boulemden and Yamina Tlili. Image indexing and retrieval with Pachinko allocation model: Application on local and global features. Lecture Notes in Computer Science, 7457:140–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32541-0_12/.

Kelarev:2012:DCE

[133] Andrei Kelarev, Richard Dazeley, Andrew Stranieri, John Yearwood, and Herbert Jelinek. Detection of CAN by ensemble classifiers based on ripple down rules. Lecture Notes in Computer Science, 7457:147– 159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32541-0_13/.

Kim:2012:IOI

[134] Myung Hee Kim and Paul Compton. Improving open information extraction for informal Web documents with ripple-down rules. Lecture Notes in Computer Science, 7457:160–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32541-0_14/.

Kim:2012:RRC

[135] Yang Sok Kim, Paul Compton, and Byeong Ho Kang. Ripple-down rules with censored production rules. Lecture Notes in Computer Science, 7457:

175–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32541-0_15/.

Maruatona:2012:RRP

[136] Omaru Maruatona, Peter Vamplew, and Richard Dazeley. RM and RDM, a preliminary evaluation of two prudent RDR techniques. Lecture Notes in Computer Science, 7457:188–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32541-0_16/.

Ang:2012:PCS

[137] Karen Ang and Ethel Ong. Planning children’s stories using agent models. Lecture Notes in Computer Science, 7457:195–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32541-0_17/.

Hanna:2012:FMA

[138] Nader Hanna and Deborah Richards. A framework for a multi-agent collaborative virtual learning environment (MACVILLE) based on activity theory. Lecture Notes in Computer Science, 7457:209–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32541-0_18/.

Ismail:2012:EPK

[139] Shahrinaz Ismail and Mohd Sharifuddin Ahmad. Emergence of personal knowledge management processes within multi-agent roles. Lecture Notes in Computer Science, 7457:221–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32541-0_19/.

Atif:2012:TOB

[140] Amara Atif, Peter Busch, and Deborah Richards. Towards an ontologybased approach to knowledge management of graduate attributes in higher education. Lecture Notes in Computer Science, 7457:229–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32541-0_20/.

Anonymous:2012:FMh

[141] Anonymous. Front matter. Lecture Notes in Computer Science, 7457: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32541-0/1.

Kitano:2012:SBP

[142] Hiroaki Kitano. Systems biology powered by artificial intelligence. Lecture Notes in Computer Science, 7458:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32695-0_1.

Rossi:2012:PRA

[143] Francesca Rossi. Preference reasoning and aggregation: Between AI and social choice. Lecture Notes in Computer Science, 7458:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32695-0_2.

Chakraborty:2012:EOT

[144] Shantanu Chakraborty, Takayuki Ito, and Tomonobu Senjyu. Economical operation of thermal generating units integrated with smart houses. Lecture Notes in Computer Science, 7458:3–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32695-0_3/.

Chitsaz:2012:CLR

[145] Mahsa Chitsaz, Kewen Wang, Michael Blumenstein, and Guilin Qi. Concept learning for EL^{++} by refinement and reinforcement. Lecture Notes in Computer Science, 7458:15–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32695-0_4/.

Erdmann:2012:HTM

[146] Maike Erdmann, Duc Dung Nguyen, Tomoya Takeyoshi, Gen Hattori, and Kazunori Matsumoto. Hierarchical training of multiple SVMs for personalized Web filtering. Lecture Notes in Computer Science, 7458: 27–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32695-0_5/.

Galgani:2012:CBS

[147] Filippo Galgani, Paul Compton, and Achim Hoffmann. Citation based summarisation of legal texts. Lecture Notes in Computer Science, 7458: 40–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32695-0_6/.

Gou:2012:NVF

[148] Gaopeng Gou, Di Huang, and Yunhong Wang. A novel video face clustering algorithm based on divide and conquer strategy. Lecture Notes in Computer Science, 7458:53–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32695-0_7/.

Gou:2012:HLF

[149] Gaopeng Gou, Di Huang, and Yunhong Wang. A hybrid local feature for face recognition. Lecture Notes in Computer Science, 7458: 64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32695-0_8/.

Grcar:2012:OQS

[150] Miha Grčar, Vid Podpečan, Borut Sluban, and Igor Mozetič. Ontology querying support in semantic annotation process. Lecture Notes in Computer Science, 7458:76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32695-0_9/.

Hao:2012:LAS

[151] Jianye Hao and Ho fung Leung. Learning to achieve socially optimal solutions in general-sum games. Lecture Notes in Computer Science, 7458: 88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32695-0_10/.

Hao:2012:IFA

[152] Jianye Hao and Ho fung Leung. Incorporating fairness into agent interactions modeled as two-player normal-form games. Lecture Notes in Computer Science, 7458:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32695-0_11/.

Huang:2012:SSD

[153] Jijian Huang, Hui Xue, and Yuqing Zhai. Semi-supervised discriminatively regularized classifier with pairwise constraints. Lecture Notes in Computer Science, 7458:112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32695-0_12/.

Hwang:2012:OSR

[154] Jihye Hwang, Kyung tai Kim, and Eun Yi Kim. Outdoor situation recognition using support vector machine for the blind and the visually impaired. Lecture Notes in Computer Science, 7458:124–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32695-0_13/.

Inokuchi:2012:MRR

[155] Akihiro Inokuchi, Ayumu Yamaoka, Takashi Washio, Yuji Matsumoto, and Masayuki Asahara. Mining rules for rewriting states in a transitionbased dependency parser. Lecture Notes in Computer Science, 7458: 133–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32695-0_14/.

Ito:2012:CSS

[156] Takayuki Ito, Ryo Kanamori, Jun Takahashi, Iván Marsa Maestre, and Enrique de la Hoz. The comparison of stigmergy strategies for decentralized traffic congestion control: Preliminary results. Lecture Notes in Computer Science, 7458:146–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32695-0_15/.

Iwasawa:2012:LLW

[157] Yusuke Iwasawa and Ikuko Eguchi Yairi. Life-logging of wheelchair driving on Web maps for visualizing potential accidents and incidents. Lecture Notes in Computer Science, 7458:157–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32695-0_16/.

Jalab:2012:TFE

[158] Hamid A. Jalab and Rabha W. Ibrahim. Texture feature extraction based on fractional mask convolution with Cesáro means for contentbased image retrieval. Lecture Notes in Computer Science, 7458:170– 179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32695-0_17/.

Jin:2012:PRM

[159] Yifan Jin, Kewen Wang, and Lian Wen. Possibilistic reasoning in multicontext systems: Preliminary report. Lecture Notes in Computer Science, 7458:180–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32695-0_18/.

Kim:2012:TSZ

[160] Kye-Sung Kim, Su-Jeong Choi, Seong-Bae Park, and Sang-Jo Lee. A twostep zero pronoun resolution by reducing candidate cardinality. Lecture Notes in Computer Science, 7458:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32695-0_19/.

Kim:2012:HTA

[161] Yang Sok Kim, Alfred Krzywicki, Wayne Wobcke, Ashesh Mahidadia, and Paul Compton. Hybrid techniques to address Cold start problems for people to people recommendation in social networks. Lecture Notes in Computer Science, 7458:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32695-0_20/.

Anonymous:2012:FMi

[162] Anonymous. Front matter. Lecture Notes in Computer Science, 7458: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32695-0/1.

Fragkaki:2012:MEA

[163] Elli Fragkaki, Lujo Bauer, Limin Jia, and David Swasey. Modeling and enhancing Android’s permission system. Lecture Notes in Computer Science, 7459:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33167-1_1/.

Kelley:2012:HAC

[164] James Kelley, Roberto Tamassia, and Nikos Triandopoulos. Hardening access control and data protection in GFS-like file systems. Lecture Notes in Computer Science, 7459:19–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33167-1_2/.

Crussell:2012:ACD

[165] Jonathan Crussell, Clint Gibler, and Hao Chen. Attack of the clones: Detecting cloned applications on Android markets. Lecture Notes in Computer Science, 7459:37–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33167-1_3/.

Birgisson:2012:BPD

[166] Arnar Birgisson, Daniel Hedin, and Andrei Sabelfeld. Boosting the permissiveness of dynamic information-flow tracking by testing. Lecture Notes in Computer Science, 7459:55–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33167-1_4/.

Erbatur:2012:ESP

[167] Serdar Erbatur, Santiago Escobar, Deepak Kapur, Zhiqiang Liu, and Christopher Lynch. Effective symbolic protocol analysis via equational irreducibility conditions. Lecture Notes in Computer Science, 7459: 73–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33167-1_5/.

Schnoor:2012:DES

[168] Henning Schnoor. Deciding epistemic and strategic properties of cryptographic protocols. Lecture Notes in Computer Science, 7459:91–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33167-1_6/.

Khan:2012:SFR

[169] Arif Akram Khan and Philip W. L. Fong. Satisfiability and feasibility in a relationship-based workflow authorization model. Lecture Notes in Computer Science, 7459:109–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33167-1_7/.

Modersheim:2012:DSF

[170] Sebastian Mödersheim. Deciding security for a fragment of ASLan. Lecture Notes in Computer Science, 7459:127–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33167-1_8/.

Albanese:2012:PFL

[171] Massimiliano Albanese, Alessandra De Benedictis, Sushil Jajodia, and Paulo Shakarian. A probabilistic framework for localization of attackers in MANETs. Lecture Notes in Computer Science, 7459:145–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33167-1_9/.

Rios:2012:RPF

[172] Ruben Rios, Jorge Cuellar, and Javier Lopez. Robust probabilistic fake packet injection for receiver-location privacy in WSN. Lecture Notes in Computer Science, 7459:163–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33167-1_10/.

Aviv:2012:PAM

[173] Adam J. Aviv, Micah Sherr, Matt Blaze, and Jonathan M. Smith. Privacyaware message exchanges for geographically routed human movement networks. Lecture Notes in Computer Science, 7459:181–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33167-1_11/.

Dacosta:2012:TNO

[174] Italo Dacosta, Mustaque Ahamad, and Patrick Traynor. Trust no one else: Detecting MITM attacks against SSL/TLS without third-parties. Lecture Notes in Computer Science, 7459:199–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33167-1_12/.

Holz:2012:XFD

[175] Ralph Holz, Thomas Riedmaier, Nils Kammenhuber, and Georg Carle. X.509 forensics: Detecting and localising the SSL/TLS men-in-the-middle. Lecture Notes in Computer Science, 7459:217–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33167-1_13/.

Eberz:2012:PMM

[176] Simon Eberz, Martin Strohmeier, Matthias Wilhelm, and Ivan Martinovic. A practical man-in-the-middle attack on signal-based key generation protocols. Lecture Notes in Computer Science, 7459:235–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33167-1_14/.

Sadeghi:2012:SLE

[177] Ahmad-Reza Sadeghi, Steffen Schulz, and Vijay Varadharajan. The silence of the LANs: Efficient leakage resilience for IPsec VPNs. Lecture Notes in Computer Science, 7459:253–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33167-1_15/.

Herzberg:2012:SPD

[178] Amir Herzberg and Haya Shulman. Security of patched DNS. Lecture Notes in Computer Science, 7459:271–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33167-1_16/.

Gu:2012:RAC

[179] Qijun Gu, Kyle Jones, Wanyu Zang, Meng Yu, and Peng Liu. Revealing abuses of channel assignment protocols in multi-channel wireless networks: An investigation logic approach. Lecture Notes in Computer Science, 7459: 289–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33167-1_17/.

Almishari:2012:ELU

[180] Mishari Almishari and Gene Tsudik. Exploring linkability of user reviews. Lecture Notes in Computer Science, 7459:307–324, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33167-1_18/.

Dong:2012:FAP

[181] Naipeng Dong, Hugo Jonker, and Jun Pang. Formal analysis of privacy in an eHealth protocol. Lecture Notes in Computer Science, 7459: 325–342, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33167-1_19/.

Koo:2012:PWF

[182] Jinkyu Koo, Xiaojun Lin, and Saurabh Bagchi. PRIVATUS: Walletfriendly privacy protection for smart meters. Lecture Notes in Computer Science, 7459:343–360, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33167-1_20/.

Anonymous:2012:FMj

[183] Anonymous. Front matter. Lecture Notes in Computer Science, 7459: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33167-1/1.

Barthe:2012:CAC

[184] Gilles Barthe, Benjamin Grégoire, and Santiago Zanella Béguelin. Computer-aided cryptographic proofs. Lecture Notes in Computer Science, 7460:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33125-1_1.

Distefano:2012:VDH

[185] Dino Distefano. A voyage to the deep-heap. Lecture Notes in Computer Science, 7460:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33125-1_2.

Krishnamurthi:2012:SAJ

[186] Shriram Krishnamurthi. Semantics and analyses for JavaScript and the Web. Lecture Notes in Computer Science, 7460:4, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33125-1_3.

Lee:2012:EMH

[187] Jonathan K. Lee, Jens Palsberg, Rupak Majumdar, and Hong Hong. Efficient may happen in parallel analysis for async-finish parallelism. Lecture Notes in Computer Science, 7460:5–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33125-1_4/.

Goubault:2012:MSA

[188] Eric Goubault, Sylvie Putot, and Franck Védrine. Modular static analysis with zonotopes. Lecture Notes in Computer Science, 7460:24–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33125-1_5/.

Howe:2012:PAU

[189] Jacob M. Howe and Andy King. Polyhedral analysis using parametric objectives. Lecture Notes in Computer Science, 7460:41–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33125-1_6/.

Cachera:2012:IPI

[190] David Cachera, Thomas Jensen, Arnaud Jobin, and Florent Kirchner. Inference of polynomial invariants for imperative programs: a farewell to Gröbner bases. Lecture Notes in Computer Science, 7460: 58–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33125-1_7/.

Ioualalen:2012:NAD

[191] Arnault Ioualalen and Matthieu Martel. A new abstract domain for the representation of mathematically equivalent expressions. Lecture Notes in Computer Science, 7460:75–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33125-1_8/.

Cheng:2012:ADI

[192] Tie Cheng and Xavier Rival. An abstract domain to infer types over zones in spreadsheets. Lecture Notes in Computer Science, 7460:94– 110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33125-1_9/.

Thakur:2012:BAS

[193] Aditya Thakur, Matt Elder, and Thomas Reps. Bilateral algorithms for symbolic abstraction. Lecture Notes in Computer Science, 7460: 111–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33125-1_10/.

Giacobazzi:2012:MAI

[194] Roberto Giacobazzi and Isabella Mastroeni. Making abstract interpretation incomplete: Modeling the potency of obfuscation. Lecture Notes in Computer Science, 7460:129–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33125-1_11/.

Sanchez:2012:IGP

[195] Alejandro Sanchez, Sriram Sankaranarayanan, César Sánchez, and Bor-Yuh Evan Chang. Invariant generation for parametrized systems using self-reflection. Lecture Notes in Computer Science, 7460:146–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33125-1_12/.

Abdulla:2012:AFI

[196] Parosh Aziz Abdulla, Mohamed Faouzi Atig, Yu-Fang Chen, Carl Leonardsson, and Ahmed Rezine. Automatic fence insertion in integer programs via predicate abstraction. Lecture Notes in Computer Science, 7460:164–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33125-1_13/.

Calvert:2012:CFA

[197] Peter Calvert and Alan Mycroft. Control flow analysis for the join calculus. Lecture Notes in Computer Science, 7460:181–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33125-1_14/.

Halbwachs:2012:WDS

[198] Nicolas Halbwachs and Julien Henry. When the decreasing sequence fails. Lecture Notes in Computer Science, 7460:198–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33125-1_15/.

Biallas:2012:LLC

[199] Sebastian Biallas, Jörg Brauer, Andy King, and Stefan Kowalewski. Loop leaping with closures. Lecture Notes in Computer Science, 7460: 214–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33125-1_16/.

Jaffar:2012:PSB

[200] Joxan Jaffar, Vijayaraghavan Murali, Jorge A. Navas, and Andrew E. Santosa. Path-sensitive backward slicing. Lecture Notes in Computer Science, 7460:231–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33125-1_17/.

Giannakopoulou:2012:SLC

[201] Dimitra Giannakopoulou, Zvonimir Rakamarić, and Vishwanath Raman. Symbolic learning of component interfaces. Lecture Notes in Computer Science, 7460:248–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33125-1_18/.

Khedker:2012:LBP

[202] Uday P. Khedker, Alan Mycroft, and Prashant Singh Rawat. Livenessbased pointer analysis. Lecture Notes in Computer Science, 7460:265– 282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33125-1_19/.

Henry:2012:SRA

[203] Julien Henry, David Monniaux, and Matthieu Moy. Succinct representations for abstract interpretation. Lecture Notes in Computer Science, 7460:283–299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33125-1_20/.

Anonymous:2012:FMk

[204] Anonymous. Front matter. Lecture Notes in Computer Science, 7460: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33125-1/1.

Kramer:2012:PSE

[205] Oliver Kramer. A particle swarm embedding algorithm for nonlinear dimensionality reduction. Lecture Notes in Computer Science, 7461: 1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32650-9_1/.

Salama:2012:AMA

[206] Khalid M. Salama and Alex A. Freitas. ABC-miner: An ant-based Bayesian classification algorithm. Lecture Notes in Computer Science, 7461:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32650-9_2/.

Massink:2012:ARS

[207] Mieke Massink, Manuele Brambilla, Diego Latella, Marco Dorigo, and Mauro Birattari. Analysing robot swarm decision-making with bio-PEPA. Lecture Notes in Computer Science, 7461:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32650-9_3/.

Bezerra:2012:AGM

[208] Leonardo C. T. Bezerra, Manuel López-Ibáñez, and Thomas Stützle. Automatic generation of multi-objective ACO algorithms for the bi-objective knapsack. Lecture Notes in Computer Science, 7461:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32650-9_4/.

al-Rifaie:2012:BBP

[209] Mohammad Majid al Rifaie and Tim Blackwell. Bare bones particle swarms with jumps. Lecture Notes in Computer Science, 7461: 49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32650-9_5/.

Mateo:2012:HAM

[210] Sergi Mateo, Christian Blum, Pascal Fua, and Engin Türetgen. Hybrid algorithms for the minimum-weight rooted arborescence problem. Lecture Notes in Computer Science, 7461:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32650-9_6/.

Medland:2012:ICM

[211] Matthew Medland, Fernando E. B. Otero, and Alex A. Freitas. Improving the cAnt-Miner_{P B} classification algorithm. Lecture Notes in Computer Science, 7461:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32650-9_7/.

Gomes:2012:INS

[212] Jorge Gomes, Paulo Urbano, and Anders Lyhne Christensen. Introducing novelty search in evolutionary swarm robotics. Lecture Notes in Computer Science, 7461:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32650-9_8/.

Ismail:2012:MDC

[213] Adiel Ismail and Andries P. Engelbrecht. Measuring diversity in the cooperative particle swarm optimizer. Lecture Notes in Computer Science, 7461:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32650-9_9/.

Pini:2012:MAB

[214] Giovanni Pini, Arne Brutschy, Gianpiero Francesca, Marco Dorigo, and Mauro Birattari. Multi-armed bandit formulation of the task partitioning problem in swarm robotics. Lecture Notes in Computer Science, 7461: 109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32650-9_10/.

Leonard:2012:SSP

[215] Barend J. Leonard and Andries P. Engelbrecht. Scalability study of particle swarm optimizers in dynamic environments. Lecture Notes in Computer Science, 7461:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32650-9_11/.

Murray:2012:SRM

[216] Lachlan Murray, Jon Timmis, and Andy Tyrrell. Self-reconfigurable modular e-pucks. Lecture Notes in Computer Science, 7461:133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32650-9_12/.

Worcester:2012:TPA

[217] James Worcester and M. Ani Hsieh. Task partitioning via ant colony optimization for distributed assembly. Lecture Notes in Computer Science, 7461:145–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32650-9_13/.

Ismail:2012:SAC

[218] Adiel Ismail and Andries P. Engelbrecht. The self-adaptive comprehensive learning particle swarm optimizer. Lecture Notes in Computer Science, 7461:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32650-9_14/.

Hamann:2012:TSC

[219] Heiko Hamann. Towards swarm calculus: Universal properties of swarm performance and collective decisions. Lecture Notes in Computer Science, 7461:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32650-9_15/.

Marinakis:2012:HPS

[220] Yannis Marinakis and Magdalene Marinaki. A hybrid particle swarm optimization algorithm for the open vehicle routing problem. Lecture Notes in Computer Science, 7461:180–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32650-9_16/.

Nepomuceno:2012:SAH

[221] Filipe V. Nepomuceno and Andries P. Engelbrecht. A self-adaptive heterogeneous PSO inspired by ants. Lecture Notes in Computer Science, 7461: 188–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32650-9_17/.

Elor:2012:TAM

[222] Yotam Elor and Alfred M. Bruckstein. A “thermodynamic” approach to multi-robot cooperative localization with noisy sensors. Lecture Notes in Computer Science, 7461:196–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32650-9_18/.

Duc:2012:AAC

[223] Dong Do Duc, Huy Q. Dinh, Thanh Hai Dang, Kris Laukens, and Xuan Huan Hoang. AcoSeeD: An ant colony optimization for finding optimal spaced seeds in biological sequence search. Lecture Notes in Computer Science, 7461:204–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32650-9_19/.

Fang:2012:AAB

[224] Rui Fang, Zequn Huang, Louis Rossi, and Chien-Chung Shen. Analysis of ant-based routing with wireless medium access control. Lecture Notes in Computer Science, 7461:212–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32650-9_20/.

Anonymous:2012:FMl

[225] Anonymous. Front matter. Lecture Notes in Computer Science, 7461: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32650-9/1.

Srivastava:2012:TVS

[226] Abhinav Srivastava, Himanshu Raj, Jonathon Giffin, and Paul England. Trusted VM snapshots in untrusted cloud infrastructures. Lecture Notes in Computer Science, 7462:1–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33338-5_1/.

Carbone:2012:SRM

[227] Martim Carbone, Matthew Conover, Bruce Montague, and Wenke Lee. Secure and robust monitoring of virtual machines through guestassisted introspection. Lecture Notes in Computer Science, 7462:22– 41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33338-5_2/.

Zhang:2012:ATD

[228] Shengzhi Zhang and Peng Liu. Assessing the trustworthiness of drivers. Lecture Notes in Computer Science, 7462:42–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33338-5_3/.

Thonnard:2012:IET

[229] Olivier Thonnard, Leyla Bilge, Gavin O’Gorman, Seán Kiernan, and Martin Lee. Industrial espionage and targeted attacks: Understanding the characteristics of an escalating threat. Lecture Notes in Computer Science, 7462:64–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33338-5_4/.

vanderVeen:2012:MEP

[230] Victor van der Veen, Nitish dutt Sharma, Lorenzo Cavallaro, and Herbert Bos. Memory errors: The past, the present, and the future. Lecture Notes in Computer Science, 7462:86–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33338-5_5/.

Ahn:2012:MAV

[231] Dongkyun Ahn and Gyungho Lee. A memory access validation scheme against payload injection attacks. Lecture Notes in Computer Science, 7462:107–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33338-5_6/.

Mankin:2012:DFD

[232] Jennifer Mankin and David Kaeli. Dione: a flexible disk monitoring and analysis framework. Lecture Notes in Computer Science, 7462: 127–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33338-5_7/.

Xu:2012:APA

[233] Zhi Xu, Hungyuan Hsu, Xin Chen, Sencun Zhu, and Ali R. Hurson. AK-PPM: An authenticated packet attribution scheme for mobile ad hoc networks. Lecture Notes in Computer Science, 7462:147–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33338-5_8/.

Lauinger:2012:PPA

[234] Tobias Lauinger, Engin Kirda, and Pietro Michiardi. Paying for piracy? An analysis of one-click hosters’ controversial reward schemes. Lecture Notes in Computer Science, 7462:169–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33338-5_9/.

Marchal:2012:PDP

[235] Samuel Marchal, Jérôme François, Radu State, and Thomas Engel. Proactive discovery of phishing related domain names. Lecture Notes in Computer Science, 7462:190–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33338-5_10/.

Mashima:2012:EET

[236] Daisuke Mashima and Alvaro A. Cárdenas. Evaluating electricity theft detectors in smart grid networks. Lecture Notes in Computer Science, 7462:210–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33338-5_11/.

Zhang:2012:PGA

[237] Jialong Zhang, Chao Yang, Zhaoyan Xu, and Guofei Gu. PoisonAmplifier: a guided approach of discovering compromised Websites through reversing search poisoning attacks. Lecture Notes in Computer Science, 7462: 230–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33338-5_12/.

Lekies:2012:DDA

[238] Sebastian Lekies, Nick Nikiforakis, Walter Tighzert, Frank Piessens, and Martin Johns. DEMACRO: Defense against malicious cross-domain requests. Lecture Notes in Computer Science, 7462:254–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33338-5_13/.

VanOverveldt:2012:FAM

[239] Timon Van Overveldt, Christopher Kruegel, and Giovanni Vigna. FlashDetect: ActionScript 3 malware detection. Lecture Notes in Computer Science, 7462:274–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33338-5_14/.

Chu:2012:AIA

[240] Jie Chu, Zihui Ge, Richard Huber, Ping Ji, Jennifer Yates, and Yung-Chao Yu. ALERT-ID: Analyze logs of the network element in real time for intrusion detection. Lecture Notes in Computer Science, 7462: 294–313, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33338-5_15/.

Amann:2012:LWN

[241] Bernhard Amann, Robin Sommer, Aashish Sharma, and Seth Hall. A lone wolf no more: Supporting network intrusion detection with real-time intelligence. Lecture Notes in Computer Science, 7462:314–333, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33338-5_16/.

Valgenti:2012:GGH

[242] Victor C. Valgenti, Jatin Chhugani, Yan Sun, Nadathur Satish, Min Sik Kim, and Changkyu Kim. GPP-grep: High-speed regular expression processing engine on general purpose processors. Lecture Notes in Computer Science, 7462:334–353, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33338-5_17/.

Hadziosmanovic:2012:GAM

[243] Dina Hadžiosmanović, Lorenzo Simionato, Damiano Bolzoni, and Emmanuele Zambon. N-gram against the machine: On the feasibility of the N-gram network analysis for binary protocols. Lecture Notes in Computer Science, 7462:354–373, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33338-5_18/.

Mahmood:2012:OSN

[244] Shah Mahmood and Yvo Desmedt. Online social networks, a criminals multipurpose toolbox (poster abstract). Lecture Notes in Computer Science, 7462:374–375, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33338-5_19.

Xiang:2012:TCM

[245] Cui Xiang, Shi Jinqiao, Liao Peng, and Liu Chaoge. The triplechannel model: Toward robust and efficient advanced botnets (poster abstract). Lecture Notes in Computer Science, 7462:376–377, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33338-5_20.

Anonymous:2012:FMm

[246] Anonymous. Front matter. Lecture Notes in Computer Science, 7462: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33338-5/1.

Aydemir:2012:PWL

[247] Alper Aydemir, Erik Järleberg, Samuel Prentice, and Patric Jensfelt. Predicting what lies ahead in the topology of indoor environments. Lecture Notes in Computer Science, 7463:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32732-2_1/.

Marton:2012:OCC

[248] Zoltan-Csaba Marton, Ferenc Balint-Benczedi, and Florian Seidel. Object categorization in clutter using additive features and hashing of part-graph descriptors. Lecture Notes in Computer Science, 7463:17– 33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32732-2_2/.

Bormann:2012:EOC

[249] Richard Bormann, Jan Fischer, Georg Arbeiter, and Alexander Verl. Efficient object categorization with the surface-approximation polynomials descriptor. Lecture Notes in Computer Science, 7463:34–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32732-2_3/.

Mitsou:2012:OSM

[250] Nikos Mitsou, Roderick de Nijs, David Lenz, and Johannes Frimberger. Online semantic mapping of urban environments. Lecture Notes in Computer Science, 7463:54–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32732-2_4/.

Fiolka:2012:SSE

[251] Torsten Fiolka, Jörg Stückler, Dominik A. Klein, and Dirk Schulz. SURE: Surface entropy for distinctive 3D features. Lecture Notes in Computer Science, 7463:74–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32732-2_5/.

delaPuente:2012:DMI

[252] Paloma de la Puente and Andrea Censi. Dense map inference with userdefined priors: From priorlets to scan eigenvariations. Lecture Notes in Computer Science, 7463:94–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32732-2_6/.

Arbeiter:2012:TGM

[253] Georg Arbeiter, Richard Bormann, Jan Fischer, and Martin Hägele. Towards geometric mapping for semi-autonomous mobile robots. Lecture Notes in Computer Science, 7463:114–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32732-2_7/.

Hertzberg:2012:TQE

[254] Christoph Hertzberg, René Wagner, and Udo Frese. Tutorial on quick and easy model fitting using the SLoM framework. Lecture Notes in Computer Science, 7463:128–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32732-2_8/.

Riecke:2012:LRH

[255] Bernhard E. Riecke. Are left-right hemisphere errors in point-to-origin tasks in VR caused by failure to incorporate heading changes? Lecture Notes in Computer Science, 7463:143–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32732-2_9/.

Giudice:2012:EVG

[256] Nicholas A. Giudice and Hengshan Li. The effects of visual granularity on indoor spatial learning assisted by mobile 3D information displays. Lecture Notes in Computer Science, 7463:163–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32732-2_10/.

Pietsch:2012:RBC

[257] Stefanie Pietsch and Petra Jansen. The relationship between coordination skill and mental rotation ability. Lecture Notes in Computer Science, 7463: 173–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32732-2_11/.

Galati:2012:CST

[258] Alexia Galati and Marios N. Avraamides. Collaborating in spatial tasks: Partners adapt the perspective of their descriptions, coordination strategies, and memory representations. Lecture Notes in Computer Science, 7463:182–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32732-2_12/.

Smith:2012:USA

[259] Linsey A. Smith and Dedre Gentner. Using spatial analogy to facilitate graph learning. Lecture Notes in Computer Science, 7463:196– 209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32732-2_13/.

Neidhardt:2012:AEP

[260] Eva Neidhardt and Michael Popp. Activity effects on path integration tasks for children in different environments. Lecture Notes in Computer Science, 7463:210–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32732-2_14/.

Neuburger:2012:IRA

[261] Sarah Neuburger, Vera Heuser, Petra Jansen, and Claudia Quaiser-Pohl. Influence of rotational axis and gender-stereotypical nature of rotation stimuli on the mental-rotation performance of male and female fifth graders. Lecture Notes in Computer Science, 7463:220–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32732-2_15/.

Lestrade:2012:TRT

[262] Sander Lestrade and Nina Reshöft. Towards a revision of the typology of motion verbs. Lecture Notes in Computer Science, 7463:230– 241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32732-2_16/.

Klippel:2012:ASQ

[263] Alexander Klippel, Jinlong Yang, Jan Oliver Wallgrün, and Frank Dylla. Assessing similarities of qualitative spatio-temporal relations. Lecture Notes in Computer Science, 7463:242–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32732-2_17/.

Meneghetti:2012:MRD

[264] Chiara Meneghetti, Francesca Pazzaglia, and Rossana De Beni. The mental representation derived from spatial descriptions is north-up oriented: The role of visuo-spatial abilities. Lecture Notes in Computer Science, 7463:262–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32732-2_18/.

Tenbrink:2012:LPS

[265] Thora Tenbrink and Marco Ragni. Linguistic principles for spatial relational reasoning. Lecture Notes in Computer Science, 7463:279–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32732-2_19/.

Lohmann:2012:EVA

[266] Kris Lohmann and Christopher Habel. Extended verbal assistance facilitates knowledge acquisition of virtual tactile maps. Lecture Notes in Computer Science, 7463:299–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32732-2_20/.

Anonymous:2012:FMn

[267] Anonymous. Front matter. Lecture Notes in Computer Science, 7463: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32732-2/1.

Gottlob:2012:COR

[268] Georg Gottlob, Marco Manna, and Michael Morak. On the complexity of ontological reasoning under disjunctive existential rules. Lecture Notes in Computer Science, 7464:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32589-2_1/.

Komusiewicz:2012:NRP

[269] Christian Komusiewicz. New races in parameterized algorithmics. Lecture Notes in Computer Science, 7464:19–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32589-2_2/.

Salibra:2012:SAS

[270] Antonino Salibra. Scott is always simple. Lecture Notes in Computer Science, 7464:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32589-2_3/.

Schweikardt:2012:TPL

[271] Nicole Schweikardt. A toolkit for proving limitations of the expressive power of logics. Lecture Notes in Computer Science, 7464:46–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32589-2_4.

Ukkonen:2012:HRG

[272] Esko Ukkonen. How to reconstruct a genome. Lecture Notes in Computer Science, 7464:48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-32589-2_5.

Walukiewicz:2012:SMR

[273] Igor Walukiewicz. Simple models for recursive schemes. Lecture Notes in Computer Science, 7464:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32589-2_6/.

Woeginger:2012:TUN

[274] Gerhard J. Woeginger. Transportation under nasty side constraints. Lecture Notes in Computer Science, 7464:61–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-32589-2_7.

Yannakakis:2012:CLF

[275] Mihalis Yannakakis. Computation of least fixed points. Lecture Notes in Computer Science, 7464:63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32589-2_8.

Ahlroth:2012:UCS

[276] Lauri Ahlroth and Pekka Orponen. Unordered constraint satisfaction games. Lecture Notes in Computer Science, 7464:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32589-2_9/.

Akutsu:2012:PTA

[277] Tatsuya Akutsu and Takeyuki Tamura. A polynomial-time algorithm for computing the maximum common subgraph of outerplanar graphs of bounded degree. Lecture Notes in Computer Science, 7464:76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32589-2_10/.

Allender:2012:RSR

[278] Eric Allender, Harry Buhrman, and Luke Friedman. Reductions to the set of random strings: The resource-bounded case. Lecture Notes in Computer Science, 7464:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32589-2_11/.

Arvind:2012:AGI

[279] Vikraman Arvind and Johannes Köbler. Approximate graph isomorphism. Lecture Notes in Computer Science, 7464:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32589-2_12/.

Arvind:2012:NOE

[280] Vikraman Arvind and Partha Mukhopadhyay. Near-optimal expanding generator sets for solvable permutation groups. Lecture Notes in Computer Science, 7464:112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32589-2_13/.

Asarin:2012:GFT

[281] Eugene Asarin and Nicolas Basset. Generating functions of timed languages. Lecture Notes in Computer Science, 7464:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32589-2_14/.

Bazgan:2012:RSP

[282] Cristina Bazgan and Morgan Chopin. The robust set problem: Parameterized complexity and approximation. Lecture Notes in Computer Science, 7464:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32589-2_15/.

Bell:2012:MMN

[283] Paul C. Bell, Mika Hirvensalo, and Igor Potapov. Mortality for 2 × 2 matrices is NP-Hard. Lecture Notes in Computer Science, 7464:148– 159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32589-2_16/.

Berwanger:2012:SCP

[284] Dietmar Berwanger and Łukasz Kaiser. Solving counter parity games. Lecture Notes in Computer Science, 7464:160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32589-2_17/.

Biedl:2012:DPG

[285] Therese Biedl and Peter Floderus. Drawing planar graphs on points inside a polygon. Lecture Notes in Computer Science, 7464:172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32589-2_18/.

Bilo:2012:NAR

[286] Davide Bilò and Anna Zych. New advances in reoptimizing the minimum Steiner tree problem. Lecture Notes in Computer Science, 7464: 184–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32589-2_19/.

Blaser:2012:SCT

[287] Markus Bläser and Bodo Manthey. Smoothed complexity theory. Lecture Notes in Computer Science, 7464:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32589-2_20/.

Anonymous:2012:FMo

[288] Anonymous. Front matter. Lecture Notes in Computer Science, 7464: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32589-2/1.

Beckers:2012:OBI

[289] Kristian Beckers, Stefan Eicker, Stephan Faßbender, Maritta Heisel, and Holger Schmidt. Ontology-based identification of research gaps and immature research areas. Lecture Notes in Computer Science, 7465: 1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32498-7_1/.

Bader:2012:CAM

[290] Gerald Bader, Amin Anjomshoaa, and A. Min Tjoa. A context-aware mashup integration guideline for enterprise 2.0. Lecture Notes in Computer Science, 7465:17–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32498-7_2/.

Kiani:2012:ANN

[291] Mostafa Kiani Deh Kiani, Barat Ghobadian, Fathollah Ommi, and Gholamhassan Najafi. Artificial neural networks approach for the prediction of thermal balance of SI engine using ethanol-gasoline blends. Lecture Notes in Computer Science, 7465:31–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32498-7_3/.

Almendros-Jimenez:2012:VXD

[292] Jesús M. Almendros-Jiménez. Validation of XML documents with SWRL. Lecture Notes in Computer Science, 7465:44–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32498-7_4/.

Gschwandtner:2012:TDT

[293] Theresia Gschwandtner, Johannes Gärtner, Wolfgang Aigner, and Silvia Miksch. A taxonomy of dirty time-oriented data. Lecture Notes in Computer Science, 7465:58–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32498-7_5/.

Ebel:2012:CRS

[294] Marius Ebel and Martin Hulin. Combining relational and semi-structured databases for an inquiry application. Lecture Notes in Computer Science, 7465:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32498-7_6/.

Jaatun:2012:HAC

[295] Martin Gilje Jaatun. Hunting for aardvarks: Can software security be measured? Lecture Notes in Computer Science, 7465:85–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32498-7_7/.

Beckers:2012:FRA

[296] Kristian Beckers and Maritta Heisel. A foundation for requirements analysis of privacy preserving software. Lecture Notes in Computer Science, 7465:93–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32498-7_8/.

Holzinger:2012:CRU

[297] Andreas Holzinger, Evgenia Popova, Bernhard Peischl, and Martina Ziefle. On complexity reduction of user interfaces for safety-critical systems. Lecture Notes in Computer Science, 7465:108–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32498-7_9/.

Jaatun:2012:SSI

[298] Martin Gilje Jaatun, Karin Bernsmed, and Astrid Undheim. Security SLAs — an idea whose time has come? Lecture Notes in Computer Science, 7465:123–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32498-7_10/.

Poniszewska-Maranda:2012:URE

[299] Aneta Poniszewska-Maranda. UML representation of extended role-based access control model with the use of usage control concept. Lecture Notes in Computer Science, 7465:131–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32498-7_11/.

Hermann:2012:FEC

[300] Eckehard Hermann, Udo Litschauer, and Jürgen Fuß. A formal equivalence classes based method for security policy conformance checking. Lecture Notes in Computer Science, 7465:146–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32498-7_12/.

Azderska:2012:TOT

[301] Tanja Ažderska and Borka Jerman Blažič. Trust as an organismic trait of E-commerce systems. Lecture Notes in Computer Science, 7465: 161–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32498-7_13/.

Holzinger:2012:MAU

[302] Andreas Holzinger, Peter Treitler, and Wolfgang Slany. Making apps useable on multiple different mobile platforms: On interoperability for business application development on Smartphones. Lecture Notes in Computer Science, 7465:176–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32498-7_14/.

Paschou:2012:EBA

[303] Mersini Paschou, Evangelos Sakkopoulos, Efrosini Sourla, and Athanasios Tsakalidis. Enhancing business APPification using SMS text messages: Metrics, strategies and alternatives. Lecture Notes in Computer Science, 7465:190–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32498-7_15/.

Montanari:2012:NDD

[304] Daniele Montanari and Piera Laura Puglisi. Near duplicate document detection for large information flows. Lecture Notes in Computer Science, 7465:203–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32498-7_16/.

Furtado:2012:TS

[305] Pedro Furtado. TinyStream sensors. Lecture Notes in Computer Science, 7465:218–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32498-7_17/.

Sato:2012:DSS

[306] Hiroshi Sato, Hisashi Kurasawa, Takeru Inoue, Motonori Nakamura, and Hajime Matsumura. Distributed sampling storage for statistical analysis of massive sensor data. Lecture Notes in Computer Science, 7465: 233–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32498-7_18/.

Aknouche:2012:IQC

[307] Rachid Aknouche, Ounas Asfari, Fadila Bentayeb, and Omar Boussaid. Integrating query context and user context in an information retrieval model based on expanded language modeling. Lecture Notes in Computer Science, 7465:244–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32498-7_19/.

Cha:2012:ISF

[308] Guang-Ho Cha. Indexing and search for fast music identification. Lecture Notes in Computer Science, 7465:259–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32498-7_20/.

Anonymous:2012:FMp

[309] Anonymous. Front matter. Lecture Notes in Computer Science, 7465: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32498-7/1.

Althoff:2012:CBR

[310] Klaus-Dieter Althoff. Case-based reasoning and expert systems. Lecture Notes in Computer Science, 7466:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32986-9_1.

Gil:2012:RES

[311] Yolanda Gil. Reproducibility and efficiency of scientific data analysis: Scientific workflows and case-based reasoning. Lecture Notes in Computer Science, 7466:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-32986-9_2.

Ahmed:2012:CAS

[312] Mobyen Uddin Ahmed and Peter Funk. A computer aided system for postoperative pain treatment combining knowledge discovery and case-based reasoning. Lecture Notes in Computer Science, 7466:3–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32986-9_3/.

Bach:2012:DCB

[313] Kerstin Bach and Klaus-Dieter Althoff. Developing case-based reasoning applications using myCBR 3. Lecture Notes in Computer Science, 7466:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32986-9_4/.

Coman:2012:DPG

[314] Alexandra Coman and Héctor Muñoz-Avila. Diverse plan generation by plan adaptation and by first-principles planning: a comparative study. Lecture Notes in Computer Science, 7466:32–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32986-9_5/.

Dieterle:2012:CBA

[315] Sebastian Dieterle and Ralph Bergmann. Case-based appraisal of Internet domains. Lecture Notes in Computer Science, 7466:47–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32986-9_6/.

Dong:2012:HEW

[316] Ruihai Dong and Markus Schaal. Harnessing the experience Web to support user-generated product reviews. Lecture Notes in Computer Science, 7466:62–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32986-9_7/.

Dufour-Lussier:2012:AST

[317] Valmi Dufour-Lussier and Florence Le Ber. Adapting spatial and temporal cases. Lecture Notes in Computer Science, 7466:77–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32986-9_8/.

Florez-Puga:2012:EML

[318] Gonzalo Flórez-Puga and Guillermo Jiménez-Díaz. eCo: Managing a library of reusable behaviours. Lecture Notes in Computer Science, 7466: 92–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32986-9_9/.

Gundersen:2012:TMS

[319] Odd Erik Gundersen. Toward measuring the similarity of complex event sequences in real-time. Lecture Notes in Computer Science, 7466: 107–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32986-9_10/.

Gomez:2012:CBP

[320] Mario Gómez and Enric Plaza. Case-based project scheduling. Lecture Notes in Computer Science, 7466:122–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32986-9_11/.

Henriet:2012:ANR

[321] Julien Henriet and Pierre-Emmanuel Leni. Adapting numerical representations of lung contours using case-based reasoning and artificial neural networks. Lecture Notes in Computer Science, 7466:137–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32986-9_12/.

Hurley:2012:ACB

[322] Barry Hurley and Barry O’Sullivan. Adaptation in a CBR-based solver portfolio for the satisfiability problem. Lecture Notes in Computer Science, 7466:152–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32986-9_13/.

Jagannathan:2012:LRB

[323] Rupa Jagannathan and Sanja Petrovic. A local rule-based attribute weighting scheme for a case-based reasoning system for radiotherapy treatment planning. Lecture Notes in Computer Science, 7466:167–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32986-9_14/.

Jaidee:2012:LRG

[324] Ulit Jaidee and Héctor Muñoz-Avila. Learning and reusing goal-specific policies for goal-driven autonomy. Lecture Notes in Computer Science, 7466:182–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32986-9_15/.

Jalali:2012:CAB

[325] Vahid Jalali and David Leake. Custom accessibility-based CCBR question selection by ongoing user classification. Lecture Notes in Computer Science, 7466:196–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32986-9_16/.

Kar:2012:FWC

[326] Debarun Kar and Sutanu Chakraborti. Feature weighting and confidence based prediction for case based reasoning systems. Lecture Notes in Computer Science, 7466:211–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32986-9_17/.

Manzano:2012:CBA

[327] Sergio Manzano and Santiago Ontañón. A case-based approach to mutual adaptation of taxonomic ontologies. Lecture Notes in Computer Science, 7466:226–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32986-9_18/.

McSherry:2012:LLA

[328] David McSherry. A lazy learning approach to explaining case-based reasoning solutions. Lecture Notes in Computer Science, 7466:241–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32986-9_19/.

Minor:2012:CWA

[329] Mirjam Minor and Mohd. Siblee Islam. Confidence in workflow adaptation. Lecture Notes in Computer Science, 7466:255–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32986-9_20/.

Anonymous:2012:FMq

[330] Anonymous. Front matter. Lecture Notes in Computer Science, 7466: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32986-9/1.

Barthes:2012:AMI

[331] Jean-Paul A. Barthès and Alistair Jones. Ambiguity in multimodal interaction with multi-touch multi-user graphics tables. Lecture Notes in Computer Science, 7467:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_1/.

Gazze:2012:WCD

[332] Davide Gazzè and Mariantonietta N. La Polla. WorkMail: Collaborative document workflow management by email. Lecture Notes in Computer Science, 7467:14–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32609-7_2/.

Lee:2012:RTM

[333] Jang Ho Lee. Real-time mobile distance learning system for Smartphone. Lecture Notes in Computer Science, 7467:24–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32609-7_3/.

Aruga:2012:CIM

[334] Takashi Aruga and Nobuyoshi Yabuki. Cooperative information management of degradation of structures in operation and management. Lecture Notes in Computer Science, 7467:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32609-7_4/.

Costin:2012:RTR

[335] Aaron Costin and Nipesh Pradhananga. Real-time resource location tracking in building information models (BIM). Lecture Notes in Computer Science, 7467:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32609-7_5/.

Yabuki:2012:CIP

[336] Nobuyoshi Yabuki and Yuta Ashida. Cooperative integration of product model and sensor data model for knowledge discovery. Lecture Notes in Computer Science, 7467:49–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_6/.

Indyk:2012:CDM

[337] Wojciech Indyk and Tomasz Kajdanowicz. Cooperative decision making algorithm for large networks using MapReduce programming model. Lecture Notes in Computer Science, 7467:53–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32609-7_7/.

Ahmadi:2012:ECE

[338] Navid Ahmadi, Mehdi Jazayeri, and Monica Landoni. Enabling cooperative educational game design on the Web. Lecture Notes in Computer Science, 7467:57–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32609-7_8/.

Fominykh:2012:LCM

[339] Mikhail Fominykh. Learning computer-mediated cooperation in 3D visualization projects. Lecture Notes in Computer Science, 7467:65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32609-7_9/.

Ubik:2012:RTS

[340] Sven Ubik, Jiří Navrátil, and Petr Žejdl. Real-time stereoscopic streaming of medical surgeries for collaborative eLearning. Lecture Notes in Computer Science, 7467:73–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_10/.

Ferreira:2012:SIM

[341] Hugo Sereno Ferreira and Tiago Boldt Sousa. Scalable integration of multiple health sensor data for observing medical patterns. Lecture Notes in Computer Science, 7467:78–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_11/.

Garcia:2012:CUI

[342] Miguel Garcia, Diana Bri, and Jaime Lloret. Collaborating using intergroup communications in group-based wireless sensor networks: Another way for saving energy. Lecture Notes in Computer Science, 7467: 85–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32609-7_12/.

Chen:2012:EET

[343] Cong Chen, Kang Zhang, and Takayuki Itoh. Empirical evidence of tags supporting high-level awareness. Lecture Notes in Computer Science, 7467:94–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32609-7_13/.

Encheva:2012:RSM

[344] Sylvia Encheva. Rough sets methods for working with uncertainty. Lecture Notes in Computer Science, 7467:102–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32609-7_14/.

Encheva:2012:PPR

[345] Sylvia Encheva and Sharil Tumin. Project proposals ranking based on closed sets properties. Lecture Notes in Computer Science, 7467: 110–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32609-7_15/.

Elsen:2012:WDS

[346] Catherine Elsen and Françoise Darses. What do strokes teach us about collaborative design? Lecture Notes in Computer Science, 7467:114– 125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32609-7_16/.

Yi:2012:CBP

[347] Ingeun Yi, Seongki Lee, and Sunyoung Jang. Communication between physical and virtual models in designing smart buildings. Lecture Notes in Computer Science, 7467:126–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_17/.

Zhao:2012:CBR

[348] Yanwei Zhao, Huan Wang, and Huanhuan Hong. Cased-based reasoning based on extension theory for conflict resolution in cooperative design. Lecture Notes in Computer Science, 7467:134–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32609-7_18/.

Felasari:2012:RCC

[349] Sushardjanti Felasari and Chengzhi Peng. The role of a city’s collective memory in supporting cooperative urban design learning. Lecture Notes in Computer Science, 7467:143–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_19/.

Solesvik:2012:SOC

[350] Marina Z. Solesvik and Tatiana Iakovleva. Simulation and optimization in collaborative ship design: Innovative approach. Lecture Notes in Computer Science, 7467:151–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32609-7_20/.

Anonymous:2012:FMr

[351] Anonymous. Front matter. Lecture Notes in Computer Science, 7467: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32609-7/1.

McGookin:2012:UAN

[352] David McGookin and Stephen A. Brewster. Understanding auditory navigation to physical landmarks. Lecture Notes in Computer Science, 7468:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32796-4_1/.

Murphy:2012:SSD

[353] Emma Murphy, Camille Moussette, Charles Verron, and Catherine Guastavino. Supporting sounds: Design and evaluation of an audio-haptic interface. Lecture Notes in Computer Science, 7468:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32796-4_2/.

Yu:2012:HAI

[354] Junlei Yu and Christopher Habel. A haptic-audio interface for acquiring spatial knowledge about apartments. Lecture Notes in Computer Science, 7468:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32796-4_3/.

Cuartielles:2012:MHT

[355] David Cuartielles, Andreas Göransson, Tony Olsson, and Ståle Stenslie. Mobile haptic technology development through artistic exploration. Lecture Notes in Computer Science, 7468:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32796-4_4/.

Bial:2012:ICT

[356] Dominik Bial, Thorsten Appelmann, Enrico Rukzio, and Albrecht Schmidt. Improving cyclists training with tactile feedback on feet. Lecture Notes in Computer Science, 7468:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32796-4_5/.

Timmermann:2012:HRY

[357] Janko Timmermann, Benjamin Poppinga, Susanne Boll, and Wilko Heuten. HapticPulse — reveal your heart rate in physical activities. Lecture Notes in Computer Science, 7468:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32796-4_6/.

Nilsson:2012:AHS

[358] Niels C. Nilsson, Rolf Nordahl, Luca Turchet, and Stefania Serafin. Audiohaptic simulation of walking on virtual ground surfaces to enhance realism. Lecture Notes in Computer Science, 7468:61–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32796-4_7/.

Kildal:2012:IDU

[359] Johan Kildal. Interacting with deformable user interfaces: Effect of material stiffness and type of deformation gesture. Lecture Notes in Computer Science, 7468:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32796-4_8/.

Palmerius:2012:IMS

[360] Karljohan Lundin Palmerius, Gunnar Höst, and Konrad Schönborn. An interactive and multi-sensory learning environment for nano education. Lecture Notes in Computer Science, 7468:81–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32796-4_9/.

Halvey:2012:AMT

[361] Martin Halvey, Michael Henderson, Stephen A. Brewster, Graham Wilson, and Stephen A. Hughes. Augmenting media with thermal stimulation. Lecture Notes in Computer Science, 7468:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32796-4_10/.

Tahiroglu:2012:EIA

[362] Koray Tahiroğlu, Johan Kildal, Teemu Ahmaniemi, Simon Overstall, and Valtteri Wikström. Embodied interactions with audio-tactile virtual objects in AHNE. Lecture Notes in Computer Science, 7468:101– 110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32796-4_11/.

Poppinga:2012:TOC

[363] Benjamin Poppinga, Martin Pielot, Wilko Heuten, and Susanne Boll. Towards an objective comparison of scanning-based interaction techniques. Lecture Notes in Computer Science, 7468:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32796-4_12/.

Altinsoy:2012:KSQ

[364] M. Ercan Altinsoy. Knocking sound as quality sign for household appliances and the evaluation of the audio-haptic interaction. Lecture Notes in Computer Science, 7468:121–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32796-4_13/.

Picinali:2012:SDT

[365] Lorenzo Picinali, Christopher Feakes, Davide A. Mauro, and Brian F. G. Katz. Spectral discrimination thresholds comparing audio and haptics for complex stimuli. Lecture Notes in Computer Science, 7468: 131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32796-4_14/.

Hogan:2012:HDR

[366] Trevor Hogan and Eva Hornecker. How does representation modality affect user-experience of data artifacts? Lecture Notes in Computer Science, 7468:141–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32796-4_15/.

Anonymous:2012:BMe

[367] Anonymous. Back matter. Lecture Notes in Computer Science, 7468: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32796-4/1.

Anonymous:2012:FMs

[368] Anonymous. Front matter. Lecture Notes in Computer Science, 7468: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32796-4/1.

Mazhelis:2012:DIT

[369] Oleksiy Mazhelis, Eetu Luoma, and Henna Warma. Defining an internetof-things ecosystem. Lecture Notes in Computer Science, 7469:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32686-8_1/.

Leminen:2012:TIE

[370] Seppo Leminen, Mika Westerlund, Mervi Rajahonka, and Riikka Siuruainen. Towards IOT ecosystems and business models. Lecture Notes in Computer Science, 7469:15–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32686-8_2/.

Kos:2012:OSI

[371] Andrej Kos, Damijan Pristov, Urban Sedlar, Janez Sterle, Mojca Volk, and Tomaž Vidonja. Open and scalable IoT platform and its applications for real time access line monitoring and alarm correlation. Lecture Notes in Computer Science, 7469:27–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32686-8_3/.

Kotis:2012:ASC

[372] Konstantinos Kotis, Artem Katasonov, and Jarkko Leino. Aligning smart and control entities in the IoT. Lecture Notes in Computer Science, 7469:39–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32686-8_4/.

Boytsov:2012:WYB

[373] Andrey Boytsov, Arkady Zaslavsky, and Zahraa Abdallah. Where have you been? Using location clustering and context awareness to understand places of interest. Lecture Notes in Computer Science, 7469: 51–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32686-8_5/.

Namiot:2012:WTN

[374] Dmitry Namiot and Manfred Sneps-Sneppe. Where are they now — safe location sharing. Lecture Notes in Computer Science, 7469:63–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32686-8_6/.

Dashkova:2012:SCC

[375] Ekaterina Dashkova and Andrei Gurtov. Survey on congestion control mechanisms for wireless sensor networks. Lecture Notes in Computer Science, 7469:75–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32686-8_7/.

Petrov:2012:IOC

[376] Vitaly Petrov, Sergey Andreev, Andrey Turlikov, and Yevgeni Koucheryavy. On IEEE 802.16m overload control for smart grid deployments. Lecture Notes in Computer Science, 7469:86–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32686-8_8/.

Abuarqoub:2012:OIE

[377] Abdelrahman Abuarqoub, Mohammad Hammoudeh, and Tariq Alsboui. An overview of information extraction from mobile wireless sensor networks. Lecture Notes in Computer Science, 7469:95–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32686-8_9/.

Jozam:2012:VSH

[378] Mohammadali Heidari Jozam, Erfaneh Allameh, Bauke De Vries, and Harry Timmermans. VR-smart home. Lecture Notes in Computer Science, 7469:107–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32686-8_10/.

Rocha:2012:SSG

[379] Oscar Rodríguez Rocha, Luis Javier Suarez-Meza, and Boris Moltchanov. Smart space governing through service mashups. Lecture Notes in Computer Science, 7469:119–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32686-8_11/.

Korolev:2012:SSA

[380] Yury Korolev, Dmitry Korzun, and Ivan Galov. Smart space applications integration: a mediation formalism and design for smart-M3. Lecture Notes in Computer Science, 7469:128–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32686-8_12/.

Smirnov:2012:SLS

[381] Alexander Smirnov, Nikolay Shilov, Alexey Kashevnik, and Nikolay Teslya. Smart logistic service for dynamic ridesharing. Lecture Notes in Computer Science, 7469:140–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32686-8_13/.

Vanrompay:2012:MAQ

[382] Yves Vanrompay and Yolande Berbers. A methodological approach to quality of future context for proactive smart systems. Lecture Notes in Computer Science, 7469:152–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32686-8_14/.

Blaunstein:2012:IAL

[383] Natan Blaunstein and Michael Borisovich Sergeev. Integration of advanced LTE technology and MIMO network based on adaptive multibeam antennas. Lecture Notes in Computer Science, 7469:164–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32686-8_15/.

Khoryaev:2012:FAD

[384] Alexey Khoryaev, Mikhail Shilov, Sergey Panteleev, Andrey Chervyakov, and Artyom Lomayev. Feasibility analysis of dynamic adjustment of TDD configurations in macro-femto heterogeneous LTE networks. Lecture Notes in Computer Science, 7469:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32686-8_16/.

Gerasimenko:2012:PCS

[385] Mikhail Gerasimenko, Sergey Andreev, Yevgeni Koucheryavy, and Alexey Trushanin. Performance comparison of system level simulators for 3GPP LTE uplink. Lecture Notes in Computer Science, 7469:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32686-8_17/.

Petrov:2012:PMA

[386] Dmitry Petrov, Ilmari Repo, and Marko Lampinen. Performance of multiflow aggregation scheme for HSDPA with joint intra-site scheduling and in presence of CQI imperfections. Lecture Notes in Computer Science, 7469:

198–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32686-8_18/.

Gudkova:2012:MRA

[387] Irina A. Gudkova and Konstantin E. Samouylov. Modelling a radio admission control scheme for video telephony service in wireless networks. Lecture Notes in Computer Science, 7469:208–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32686-8_19/.

Liu:2012:MPC

[388] Wei Liu, Yueyun Chen, and Yudong Yao. Multi-point cooperative fountain codes multicast for LTE cellular system. Lecture Notes in Computer Science, 7469:216–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32686-8_20/.

Anonymous:2012:FMt

[389] Anonymous. Front matter. Lecture Notes in Computer Science, 7469: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32686-8/1.

Foster:2012:TCA

[390] Nate Foster, Kazutaka Matsuda, and Janis Voigtländer. Three complementary approaches to bidirectional programming. Lecture Notes in Computer Science, 7470:1–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32202-0_1/.

Hinze:2012:GPA

[391] Ralf Hinze. Generic programming with adjunctions. Lecture Notes in Computer Science, 7470:47–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32202-0_2/.

Kiselyov:2012:TTF

[392] Oleg Kiselyov. Typed tagless final interpreters. Lecture Notes in Computer Science, 7470:130–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32202-0_3/.

Siek:2012:CCE

[393] Jeremy G. Siek. The C++0x “concepts” effort. Lecture Notes in Computer Science, 7470:175–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32202-0_4/.

Weirich:2012:GPD

[394] Stephanie Weirich and Chris Casinghino. Generic programming with dependent types. Lecture Notes in Computer Science, 7470:217–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32202-0_5/.

Anonymous:2012:BMf

[395] Anonymous. Back matter. Lecture Notes in Computer Science, 7470: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32202-0/1.

Anonymous:2012:FMu

[396] Anonymous. Front matter. Lecture Notes in Computer Science, 7470: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32202-0/1.

Ranathunga:2012:ICA

[397] Surangika Ranathunga and Stephen Cranefield. Interfacing a cognitive agent platform with second life. Lecture Notes in Computer Science, 7471:1–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32326-3_1/.

vanOijen:2012:CMI

[398] Joost van Oijen, Loïs Vanhée, and Frank Dignum. CIGA: a middleware for intelligent agents in virtual environments. Lecture Notes in Computer Science, 7471:22–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32326-3_2/.

Gemrot:2012:HCU

[399] Jakub Gemrot, Cyril Brom, and Joanna Bryson. How to compare usability of techniques for the specification of virtual agents’ behavior? An experimental pilot study with human subjects. Lecture Notes in Computer Science, 7471:38–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32326-3_3/.

Kao:2012:DDV

[400] Edward Chao-Chun Kao and Von-Wun Soo. Dialog designs in virtual drama: Balancing agency and scripted dialogs. Lecture Notes in Computer Science, 7471:63–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32326-3_4/.

Cai:2012:LPA

[401] Yundong Cai and Zhiqi Shen. Learning by playing in agent-oriented virtual learning environment. Lecture Notes in Computer Science, 7471: 79–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32326-3_5/.

Tsukamoto:2012:CAM

[402] Takeo Tsukamoto and Yumi Muroya. Collection and analysis of multimodal interaction in direction-giving dialogues: Towards an automatic gesture selection mechanism for metaverse avatars. Lecture Notes in Computer Science, 7471:94–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32326-3_6/.

Westra:2012:OSA

[403] Joost Westra, Frank Dignum, and Virginia Dignum. Organizing scalable adaptation in serious games. Lecture Notes in Computer Science, 7471: 106–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32326-3_7/.

Chien:2012:IPD

[404] Alex Yu-Hung Chien and Von-Wun Soo. Inferring pragmatics from dialogue contexts in simulated virtual agent games. Lecture Notes in Computer Science, 7471:123–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32326-3_8/.

Alvarez-Napagao:2012:SAE

[405] Sergio Alvarez-Napagao. Socially-aware emergent narrative. Lecture Notes in Computer Science, 7471:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32326-3_9/.

Wissner:2012:ILM

[406] Michael Wißner, Wouter Beek, and Esther Lozano. Increasing learners’ motivation through pedagogical agents: The cast of virtual characters in the DynaLearn ILE. Lecture Notes in Computer Science, 7471: 151–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32326-3_10/.

Hadad:2012:AAH

[407] Meirav Hadad and Avi Rosenfeld. ADAPT: Abstraction hierarchies to better simulate teamwork under dynamics. Lecture Notes in Computer Science, 7471:166–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32326-3_11/.

Campano:2012:AAB

[408] Sabrina Campano and Etienne de Sevin. An architecture for affective behaviour based on the conservation of resources. Lecture Notes in Computer Science, 7471:183–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32326-3_12/.

Anonymous:2012:BMg

[409] Anonymous. Back matter. Lecture Notes in Computer Science, 7471: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32326-3/1.

Anonymous:2012:FMv

[410] Anonymous. Front matter. Lecture Notes in Computer Science, 7471: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32326-3/1.

Tang:2012:ISM

[411] Jiliang Tang, Xufei Wang, and Huan Liu. Integrating social media data for community detection. Lecture Notes in Computer Science, 7472: 1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33684-3_1/.

Atzmueller:2012:FFC

[412] Martin Atzmueller and Stephan Doerfel. Face-to-face contacts at a conference: Dynamics of communities and roles. Lecture Notes in Computer Science, 7472:21–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33684-3_2/.

Singer:2012:FIC

[413] Philipp Singer and Claudia Wagner. Factors influencing the co-evolution of social and content networks in online social media. Lecture Notes in Computer Science, 7472:40–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33684-3_3/.

Redmond:2012:MDS

[414] Ursula Redmond and Martin Harrigan. Mining dense structures to uncover anomalous behaviour in financial network data. Lecture Notes in Computer Science, 7472:60–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33684-3_4/.

Lemmerich:2012:DLU

[415] Florian Lemmerich and Martin Atzmueller. Describing locations using tags and images: Explorative pattern mining in social media. Lecture Notes in Computer Science, 7472:77–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33684-3_5/.

Appice:2012:LTG

[416] Annalisa Appice and Michelangelo Ceci. Learning and transferring geographically weighted regression trees across time. Lecture Notes in Computer Science, 7472:97–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33684-3_6/.

Guccione:2012:TCB

[417] Pietro Guccione and Annalisa Appice. Trend cluster based kriging interpolation in sensor data networks. Lecture Notes in Computer Science, 7472: 118–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33684-3_7/.

Nygren:2012:SUP

[418] Else Nygren. Simulation of user participation and interaction in online discussion groups. Lecture Notes in Computer Science, 7472:138– 157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33684-3_8/.

Tesoriero:2012:MDP

[419] Ricardo Tesoriero and Mohamed Bourimi. Model-driven privacy and security in multi-modal social media UIs. Lecture Notes in Computer Science, 7472:158–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33684-3_9/.

Anonymous:2012:BMh

[420] Anonymous. Back matter. Lecture Notes in Computer Science, 7472: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33684-3/1.

Anonymous:2012:FMw

[421] Anonymous. Front matter. Lecture Notes in Computer Science, 7472: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33684-3/1.

Jianfang:2012:NEN

[422] Shan Jianfang and Liu Zongtian. A novel event network matching algorithm. Lecture Notes in Computer Science, 7473:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34062-8_1/.

Jiang:2012:SPM

[423] Kunpeng Jiang, Huifang Guo, Shengping Zhu, and Julong Lan. Static patterns matching for high speed networks. Lecture Notes in Computer Science, 7473:15–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34062-8_2/.

Li:2012:IPO

[424] Niu Li and Han Xiaoting. Information propagation in online social networks based on user behavior. Lecture Notes in Computer Science, 7473:

23–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_3/.

Li:2012:ROM

[425] Xin Li and Yue fang Yang. Research on operation and management of railway transport of dangerous goods in third-party logistics enterprises. Lecture Notes in Computer Science, 7473:31–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34062-8_4/.

Shen:2012:PAS

[426] Jing Shen and Yusong Du. A password authentication scheme against Smart Card security breach. Lecture Notes in Computer Science, 7473: 37–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_5/.

Han:2012:VAG

[427] Bo Han, Qing Wang, Fajiang Yu, and Xianda Zhang. A vulnerability attack graph generation method based on scripts. Lecture Notes in Computer Science, 7473:45–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34062-8_6/.

Abdullah:2012:DGE

[428] Zailani Abdullah, Tutut Herawan, A. Noraziah, and Mustafa Mat Deris. DFP-growth: An efficient algorithm for mining frequent patterns in dynamic database. Lecture Notes in Computer Science, 7473:51– 58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34062-8_7/.

Zhang:2012:AKN

[429] Xizhe Zhang, Guolong Zhao, Tianyang Lv, Ying Yin, and Bin Zhang. Analysis on key nodes behavior for complex software network. Lecture Notes in Computer Science, 7473:59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34062-8_8/.

Ren:2012:WIH

[430] Junling Ren and Li Zhang. Webpage information hiding algorithm based on integration of tags and data. Lecture Notes in Computer Science, 7473:67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_9/.

Zhu:2012:LMU

[431] Wei Zhu, Ping Sun, Ying nan Ma, Rui Song, Shi wei He, and Ke Hui Liu. Location method of underground pipeline monitoring point based on cavity detection. Lecture Notes in Computer Science, 7473:75–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34062-8_10/.

Zhong:2012:RPI

[432] Qichun Zhong and Jinghong Hu. Research on purified Internet environment for College students. Lecture Notes in Computer Science, 7473: 82–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_11/.

Li:2012:MCS

[433] Qianmu Li, Qiugan Shi, Jun Hou, Yong Qi, and Hong Zhang. A mobilecertificate security method of satellite-Earth integration networks. Lecture Notes in Computer Science, 7473:88–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34062-8_12/.

Lai:2012:DME

[434] Wen-Hsing Lai and Siou-Lin Wang. Duration modeling for emotional speech. Lecture Notes in Computer Science, 7473:98–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34062-8_13/.

Zhao:2012:RIR

[435] Hongwei Zhao, Xiao Chen, Wei Huang, Pingping Liu, and Lingjiao Ma. Research on image retrieval based on color and shape features. Lecture Notes in Computer Science, 7473:104–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34062-8_14/.

Wang:2012:ESP

[436] Kaihua Wang, Wenxiang Zhang, and Zhanji Gui. Existence and simulations of periodic solution for impulsive predator-prey system with stage structure for the predator. Lecture Notes in Computer Science, 7473:

112–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_15/.

Yan:2012:DSM

[437] Yan Yan, Kaihua Wang, and Zhanji Gui. Dynamics and simulations of multi-species competition-predator system with impulsive. Lecture Notes in Computer Science, 7473:120–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34062-8_16/.

Chen:2012:RDS

[438] Ying Chen and Shaobai Zhang. Research on the distal supervised learning model of speech inversion. Lecture Notes in Computer Science, 7473: 128–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_17/.

Chen:2012:LPP

[439] Ruei-Chang Chen, Shih-Fong Lee, and Yeong-Chau Kuo. Low power pulse width modulation design for Class D audio amplifier systems. Lecture Notes in Computer Science, 7473:136–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34062-8_18/.

Wei:2012:NLM

[440] Guan Wei and Lv Yuanhai. A new logic method for education resource software guarantee. Lecture Notes in Computer Science, 7473: 144–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34062-8_19/.

Hou:2012:ROR

[441] Zhenjie Hou, Junsheng Huang, and Jianhua Zhang. Research on 3D object rounding photography systems and technology. Lecture Notes in Computer Science, 7473:152–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34062-8_20/.

Anonymous:2012:FMx

[442] Anonymous. Front matter. Lecture Notes in Computer Science, 7473: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34062-8/1.

Leal-Taixe:2012:EPI

[443] Laura Leal-Taixé and Gerard Pons-Moll. Exploiting pedestrian interaction via global optimization and social behaviors. Lecture Notes in Computer Science, 7474:1–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34091-8_1/.

Schneider:2012:EFS

[444] Nicolai Schneider and Stefan Gehrig. An evaluation framework for stereo-based driver assistance. Lecture Notes in Computer Science, 7474: 27–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_2/.

Morales:2012:RWS

[445] Sandino Morales and Simon Hermann. Real-world stereo-analysis evaluation. Lecture Notes in Computer Science, 7474:52–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34091-8_3/.

Mochizuki:2012:PTS

[446] Yoshihiko Mochizuki and Atsushi Imiya. Pyramid transform and scalespace analysis in image analysis. Lecture Notes in Computer Science, 7474:78–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_4/.

Dragon:2012:TFB

[447] Ralf Dragon, Michele Fenzi, and Wolf Siberski. Towards feature-based situation assessment for airport apron video surveillance. Lecture Notes in Computer Science, 7474:110–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34091-8_5/.

Jian:2012:GSP

[448] Yong-Dian Jian and Doru C. Balcan. Generalized subgraph preconditioners for large-scale bundle adjustment. Lecture Notes in Computer Science, 7474:131–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34091-8_6/.

Brilakis:2012:ACR

[449] Ioannis Brilakis and Fei Dai. Achievements and challenges in recognizing and reconstructing civil infrastructure. Lecture Notes in Computer Science, 7474:151–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34091-8_7/.

Raviv:2012:EAI

[450] Dan Raviv and Alexander M. Bronstein. Equi-affine invariant geometries of articulated objects. Lecture Notes in Computer Science, 7474: 177–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_8/.

Sattler:2012:TFI

[451] Torsten Sattler, Bastian Leibe, and Leif Kobbelt. Towards fast imagebased localization on a city-scale. Lecture Notes in Computer Science, 7474:191–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34091-8_9/.

Sedlazeck:2012:PNP

[452] Anne Sedlazeck and Reinhard Koch. Perspective and non-perspective camera models in underwater imaging — overview and error analysis. Lecture Notes in Computer Science, 7474:212–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34091-8_10/.

Gall:2012:IRF

[453] Juergen Gall and Nima Razavi. An introduction to random forests for multi-class object detection. Lecture Notes in Computer Science, 7474: 243–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_11/.

Wegner:2012:SCO

[454] Jan D. Wegner, Bodo Rosenhahn, and Uwe Sörgel. Segmentation and classification of objects with implicit scene context. Lecture Notes in Computer Science, 7474:264–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34091-8_12/.

Mayer:2012:DRW

[455] Helmut Mayer and Jan Bartelsen. Dense 3D reconstruction from wide baseline image sets. Lecture Notes in Computer Science, 7474:285–304, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34091-8_13/.

Pons-Moll:2012:DDM

[456] Gerard Pons-Moll and Laura Leal-Taixé. Data-driven manifolds for outdoor motion capture. Lecture Notes in Computer Science, 7474: 305–328, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_14/.

Kondermann:2012:PAO

[457] Daniel Kondermann and Steffen Abraham. On performance analysis of optical flow algorithms. Lecture Notes in Computer Science, 7474: 329–355, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_15/.

Debard:2012:CBF

[458] Glen Debard and Peter Karsmakers. Camera-based fall detection on real world data. Lecture Notes in Computer Science, 7474:356–375, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34091-8_16/.

Bao:2012:SSM

[459] Sid Yingze Bao and Silvio Savarese. Semantic structure from motion: a novel framework for joint object recognition and 3D reconstruction. Lecture Notes in Computer Science, 7474:376–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34091-8_17/.

Klowsky:2012:HSR

[460] Ronny Klowsky and Patrick Mücke. Hierarchical surface reconstruction from multi-resolution point samples. Lecture Notes in Computer Science, 7474:398–418, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34091-8_18/.

Reulke:2012:TOS

[461] Ralf Reulke, Dominik Rueß, and Kristian Manthey. Traffic observation and situation assessment. Lecture Notes in Computer Science, 7474: 419–441, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34091-8_19/.

Anonymous:2012:BMi

[462] Anonymous. Back matter. Lecture Notes in Computer Science, 7474: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34091-8/1.

Anonymous:2012:FMy

[463] Anonymous. Front matter. Lecture Notes in Computer Science, 7474: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34091-8/1.

Frohlich:2012:TGA

[464] Björn Fröhlich, Erik Rodner, and Joachim Denzler. As time goes by— anytime semantic segmentation with iterative context forests. Lecture Notes in Computer Science, 7476:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32717-9_1/.

Zankl:2012:ILI

[465] Georg Zankl, Yll Haxhimusa, and Adrian Ion. Interactive labeling of image segmentation hierarchies. Lecture Notes in Computer Science, 7476: 11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32717-9_2/.

Nagaraja:2012:HLR

[466] Naveen Shankar Nagaraja, Peter Ochs, Kun Liu, and Thomas Brox. Hierarchy of localized random forests for video annotation. Lecture Notes in Computer Science, 7476:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_3/.

Ballester:2012:TLO

[467] Coloma Ballester, Lluis Garrido, Vanel Lazcano, and Vicent Caselles. A TV-L1 optical flow method with occlusion detection. Lecture Notes in Computer Science, 7476:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_4/.

Shekhovtsov:2012:CPM

[468] Alexander Shekhovtsov, Pushmeet Kohli, and Carsten Rother. Curvature prior for MRF-based segmentation and shape inpainting. Lecture Notes in Computer Science, 7476:41–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_5/.

Schelten:2012:MFC

[469] Kevin Schelten and Stefan Roth. Mean field for continuous high-order MRFs. Lecture Notes in Computer Science, 7476:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32717-9_6/.

Gao:2012:HWD

[470] Qi Gao and Stefan Roth. How Well do filter-based MRFs model natural images? Lecture Notes in Computer Science, 7476:62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32717-9_7/.

Schroers:2012:ARI

[471] Christopher Schroers, Henning Zimmer, Levi Valgaerts, Andrés Bruhn, and Oliver Demetz. Anisotropic range image integration. Lecture Notes in Computer Science, 7476:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_8/.

Schmidt:2012:MSS

[472] Thorsten Schmidt, Margret Keuper, Taras Pasternak, Klaus Palme, and Olaf Ronneberger. Modeling of sparsely sampled tubular surfaces using coupled curves. Lecture Notes in Computer Science, 7476:83–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32717-9_9/.

Winkelbach:2012:SSS

[473] Simon Winkelbach, Jens Spehr, Dirk Buchholz, Markus Rilk, and Friedrich M. Wahl. Shape (self-)similarity and dissimilarity rating for segmentation and matching. Lecture Notes in Computer Science, 7476: 93–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32717-9_10/.

Ummenhofer:2012:DRH

[474] Benjamin Ummenhofer and Thomas Brox. Dense 3D reconstruction with a hand-held camera. Lecture Notes in Computer Science, 7476: 103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32717-9_11/.

Aldoma:2012:OCO

[475] Aitor Aldoma, Federico Tombari, Radu Bogdan Rusu, and Markus Vincze. OUR-CVFH — oriented, unique and repeatable clustered viewpoint feature histogram for object recognition and 6DOF pose estimation. Lecture Notes in Computer Science, 7476:113–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32717-9_12/.

Fenzi:2012:ORP

[476] Michele Fenzi, Ralf Dragon, Laura Leal-Taixé, Bodo Rosenhahn, and Jörn Ostermann. 3D object recognition and pose estimation for multiple objects using multi-prioritized RANSAC and model updating. Lecture Notes in Computer Science, 7476:123–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_13/.

Bilen:2012:CGL

[477] Hakan Bilen, Vinay P. Namboodiri, and Luc J. Van Gool. Classification with global, local and shared features. Lecture Notes in Computer Science, 7476:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32717-9_14/.

Franzel:2012:ODM

[478] Thorsten Franzel, Uwe Schmidt, and Stefan Roth. Object detection in multi-view X-ray images. Lecture Notes in Computer Science, 7476: 144–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32717-9_15/.

Hahmann:2012:ELU

[479] Ferdinand Hahmann, Heike Ruppertshofen, Gordon Böer, Ralf Stannarius, and Hauke Schramm. Eye localization using the discriminative generalized Hough transform. Lecture Notes in Computer Science, 7476:

155–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32717-9_16/.

Fugl:2012:SEM

[480] Andreas Rune Fugl, Andreas Jordt, Henrik Gordon Petersen, and Morten Willatzen. Simultaneous estimation of material properties and pose for deformable objects from depth and color images. Lecture Notes in Computer Science, 7476:165–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32717-9_17/.

vonEnzberg:2012:SQI

[481] Sebastian von Enzberg and Bernd Michaelis. Surface quality inspection of deformable parts with variable B-spline surfaces. Lecture Notes in Computer Science, 7476:175–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_18/.

Zach:2012:AIF

[482] Fabian Zach, Christian Riess, and Elli Angelopoulou. Automated image forgery detection through classification of JPEG ghosts. Lecture Notes in Computer Science, 7476:185–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32717-9_19/.

Kostinger:2012:SBL

[483] Martin Köstinger, Peter M. Roth, and Horst Bischof. Synergy-based learning of facial identity. Lecture Notes in Computer Science, 7476: 195–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32717-9_20/.

Anonymous:2012:FMz

[484] Anonymous. Front matter. Lecture Notes in Computer Science, 7476: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32717-9/1.

Mokbel:2012:HQC

[485] Bassam Mokbel, Sebastian Gross, Markus Lux, Niels Pinkwart, and Barbara Hammer. How to quantitatively compare data dissimilarities for unsupervised machine learning? Lecture Notes in Computer Science, 7477:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_1/.

Hofmann:2012:KRS

[486] Daniela Hofmann and Barbara Hammer. Kernel robust soft learning vector quantization. Lecture Notes in Computer Science, 7477:14–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33212-8_2/.

Maltoni:2012:ILM

[487] Davide Maltoni and Erik M. Rehn. Incremental learning by message passing in hierarchical temporal memory. Lecture Notes in Computer Science, 7477:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_3/.

Lausser:2012:RPS

[488] Ludwig Lausser, Christoph Müssel, and Hans A. Kestler. Representative prototype sets for data characterization and classification. Lecture Notes in Computer Science, 7477:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33212-8_4/.

Nagatani:2012:FSB

[489] Takashi Nagatani and Shigeo Abe. Feature selection by block addition and block deletion. Lecture Notes in Computer Science, 7477: 48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_5/.

Tokic:2012:GAE

[490] Michel Tokic and Günther Palm. Gradient algorithms for exploration/ exploitation trade-offs: Global and local variants. Lecture Notes in Computer Science, 7477:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33212-8_6/.

Trentin:2012:TNP

[491] Edmondo Trentin and Marco Bongini. Towards a novel probabilistic graphical model of sequential data: Fundamental notions and a solution to the problem of parameter learning. Lecture Notes in Computer Science, 7477:72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_7/.

Bongini:2012:TNP

[492] Marco Bongini and Edmondo Trentin. Towards a novel probabilistic graphical model of sequential data: a solution to the problem of structure learning and an empirical evaluation. Lecture Notes in Computer Science, 7477:82–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33212-8_8/.

Savchenko:2012:SRS

[493] Andrey V. Savchenko. Statistical recognition of a set of patterns using novel probability neural network. Lecture Notes in Computer Science, 7477:93–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_9/.

Schmidt:2012:GAM

[494] Miriam Schmidt, Günther Palm, and Friedhelm Schwenker. On graphassociated matrices and their eigenvalues for optical character recognition. Lecture Notes in Computer Science, 7477:104–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33212-8_10/.

Manzan:2012:IIR

[495] José Ricardo Gonçalves Manzan, Shigueo Nomura, and Keiji Yamanaka. Improving Iris recognition through new target vectors in MLP artificial neural networks. Lecture Notes in Computer Science, 7477:115– 126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_11/.

Garcia-Manso:2012:RCS

[496] Antonio García-Manso, Carlos J. García-Orellana, and Ramón Gallardo-Caballero. Robustness of a CAD system on digitized mammograms. Lecture Notes in Computer Science, 7477:127–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33212-8_12/.

Roy:2012:FER

[497] Kaushik Roy and Mohamed S. Kamel. Facial expression recognition using game theory. Lecture Notes in Computer Science, 7477:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33212-8_13/.

Zamberletti:2012:CSO

[498] Alessandro Zamberletti, Ignazio Gallo, Simone Albertini, Marco Vanetti, and Angelo Nodari. Classification of segmented objects through a multi-net approach. Lecture Notes in Computer Science, 7477:151–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33212-8_14/.

Korfiatis:2012:DSS

[499] Vasileios Korfiatis, Simone Tassani, and George K. Matsopoulos. A decision support system for the prediction of the trabecular fracture zone. Lecture Notes in Computer Science, 7477:163–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33212-8_15/.

Fernandez:2012:TPI

[500] Kelwin Fernandez and Carolina Chang. Teeth/palate and interdental segmentation using artificial neural networks. Lecture Notes in Computer Science, 7477:175–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33212-8_16/.

Meudt:2012:ISA

[501] Sascha Meudt and Friedhelm Schwenker. On instance selection in audio based emotion recognition. Lecture Notes in Computer Science, 7477: 186–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33212-8_17/.

Hillebrand:2012:TSC

[502] Matthias Hillebrand, Ulrich Kreßel, Christian Wöhler, and Franz Kummert. Traffic sign classifier adaption by semi-supervised co-training. Lecture Notes in Computer Science, 7477:193–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33212-8_18/.

Pearson:2012:USO

[503] Paul T. Pearson and Cameron I. Cooper. Using self organizing maps to analyze demographics and swing state voting in the 2008 U.S. Presidential Election. Lecture Notes in Computer Science, 7477:201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33212-8_19/.

deSouza:2012:GIR

[504] Osvaldo de Souza, Paulo César Cortez, and Francisco A. T. F. da Silva. Grayscale images and RGB video: Compression by morphological neural network. Lecture Notes in Computer Science, 7477:213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33212-8_20/.

Anonymous:2012:BMj

[505] Anonymous. Back matter. Lecture Notes in Computer Science, 7477: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33212-8/1.

Anonymous:2012:FMba

[506] Anonymous. Front matter. Lecture Notes in Computer Science, 7477: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33212-8/1.

Abdalla:2012:CTT

[507] Amin Abdalla and Andrew U. Frank. Combining trip and task planning: How to get from A to passport. Lecture Notes in Computer Science, 7478:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33024-7_1/.

Anderson-Tarver:2012:ACD

[508] Chris Anderson-Tarver, Mike Gleason, and Barbara Buttenfield. Automated centerline delineation to enrich the national hydrography dataset. Lecture Notes in Computer Science, 7478:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33024-7_2/.

Buchin:2012:ESO

[509] Kevin Buchin, Bettina Speckmann, and Sander Verdonschot. Evolution strategies for optimizing rectangular cartograms. Lecture Notes in Computer Science, 7478:29–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33024-7_3/.

Buchin:2012:CAS

[510] Maike Buchin, Somayeh Dodge, and Bettina Speckmann. Context-aware similarity of trajectories. Lecture Notes in Computer Science, 7478: 43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33024-7_4/.

Chiang:2012:GNR

[511] Yao-Yi Chiang and Craig A. Knoblock. Generating named road vector data from raster maps. Lecture Notes in Computer Science, 7478: 57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33024-7_5/.

Dube:2012:OCT

[512] Matthew P. Dube and Max J. Egenhofer. An ordering of convex topological relations. Lecture Notes in Computer Science, 7478:72–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33024-7_6/.

Gaffuri:2012:TWM

[513] Julien Gaffuri. Toward Web mapping with vector data. Lecture Notes in Computer Science, 7478:87–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33024-7_7/.

Kessler:2012:SER

[514] Carsten Keßler, Krzysztof Janowicz, and Tomi Kauppinen. spatial@linkedscience — exploring the research field of GIScience with linked data. Lecture Notes in Computer Science, 7478:102–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33024-7_8/.

Maisonneuve:2012:CSI

[515] Nicolas Maisonneuve and Bastien Chopard. Crowdsourcing satellite imagery analysis: Study of parallel and iterative models. Lecture Notes in Computer Science, 7478:116–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33024-7_9/.

Mohan:2012:QRS

[516] Pradeep Mohan, Xun Zhou, and Shashi Shekhar. Quantifying resolution sensitivity of spatial autocorrelation: a resolution correlogram approach. Lecture Notes in Computer Science, 7478:132–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33024-7_10/.

Nittel:2012:LSD

[517] Silvia Nittel, Christopher Dorr, and John C. Whittier. LocalAlert: Simulating decentralized ad-hoc collaboration in emergency situations. Lecture Notes in Computer Science, 7478:146–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33024-7_11/.

Rude:2012:HLE

[518] Avinash Rude and Kate Beard. High-level event detection in spatially distributed time series. Lecture Notes in Computer Science, 7478: 160–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33024-7_12/.

Siqueira:2012:TVG

[519] Thiago Luís Lopes Siqueira. Towards vague geographic data warehouses. Lecture Notes in Computer Science, 7478:173–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33024-7_13/.

Sun:2012:MIB

[520] Guibo Sun, Hui Lin, and Rongrong Li. Measuring the influence of built environment on walking behavior: An accessibility approach. Lecture Notes in Computer Science, 7478:187–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33024-7_14/.

Touya:2012:SWA

[521] Guillaume Touya. Social welfare to assess the global legibility of a generalized map. Lecture Notes in Computer Science, 7478:198–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33024-7_15/.

Wallgrun:2012:ICC

[522] Jan Oliver Wallgrün, Jinlong Yang, and Alexander Klippel. Investigations into the cognitive conceptualization and similarity assessment of spatial scenes. Lecture Notes in Computer Science, 7478:212–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33024-7_16/.

Walton:2012:QBM

[523] Lisa A. Walton and Michael Worboys. A qualitative bigraph model for indoor space. Lecture Notes in Computer Science, 7478:226–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33024-7_17/.

Watanabe:2012:DRC

[524] Toyohide Watanabe and Kosuke Yamamoto. Dynamic refuse collection strategy based on adjacency relationship between Euler cycles. Lecture Notes in Computer Science, 7478:241–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33024-7_18/.

Wei:2012:IIL

[525] Ting Wei and Scott Bell. Impact of indoor location information reliability on users’ trust of an indoor positioning system. Lecture Notes in Computer Science, 7478:258–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33024-7_19/.

Wiegand:2012:OEG

[526] Nancy Wiegand. Ontology for the engineering of geospatial systems. Lecture Notes in Computer Science, 7478:270–283, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33024-7_20/.

Anonymous:2012:FMbb

[527] Anonymous. Front matter. Lecture Notes in Computer Science, 7478: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33024-7/1.

Porsch:2012:TAE

[528] Marco Porsch and Thomas Bauschert. A testbed analysis of the effects of IEEE 802.11s power save on mesh link performance. Lecture Notes in Computer Science, 7479:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_1/.

Rafique:2012:PSA

[529] Muhammad Irfan Rafique and Thomas Bauschert. Path selection and adaptive selection of smart antenna transmission schemes in multi-hop wireless networks. Lecture Notes in Computer Science, 7479:12–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32808-4_2/.

Cano:2012:RIV

[530] Cristina Cano, Boris Bellalta, and Miquel Oliver. Receiver-initiated vs. short-preamble burst MAC approaches for multi-channel wireless sensor networks. Lecture Notes in Computer Science, 7479:23–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32808-4_3/.

Domingo:2012:WSU

[531] Albert Domingo, Boris Bellalta, and Miquel Oliver. White spaces in UHF band: Catalonia case study and impact of the digital dividend. Lecture Notes in Computer Science, 7479:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32808-4_4/.

Tavakoli:2012:FSC

[532] Zahra Tavakoli, Sebastian Meier, and Alexander Vensmer. A framework for security context migration in a firewall secured virtual machine environment. Lecture Notes in Computer Science, 7479:41–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32808-4_5/.

Czirkos:2012:ECI

[533] Zoltán Czirkos and Gábor Hosszú. Enhancing collaborative intrusion detection methods using a Kademlia overlay network. Lecture Notes in Computer Science, 7479:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_6/.

Celeda:2012:FBS

[534] Pavel Čeleda, Radek Krejčí, and Vojtěch Krmíček. Flow-based security issue detection in building automation and control networks. Lecture Notes in Computer Science, 7479:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_7/.

Godor:2012:HBM

[535] Győző Gódor and Sándor Imre. Hash-based mutual authentication protocol for low-cost RFID systems. Lecture Notes in Computer Science, 7479:76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32808-4_8/.

Ottow:2012:IIP

[536] Christiaan Ottow, Frank van Vliet, Pieter-Tjerk de Boer, and Aiko Pras. The impact of IPv6 on penetration testing. Lecture Notes in Computer Science, 7479:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32808-4_9/.

Alloush:2012:LTS

[537] Iyas Alloush, Vanea Chiprianov, Yvon Kermarrec, and Siegfried Rouvrais. Linking telecom service high-level abstract models to simulators based on model transformations: The IMS case study. Lecture Notes in Computer Science, 7479:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32808-4_10/.

Turk:2012:NMO

[538] Stefan Türk, Ying Liu, Rico Radeke, and Ralf Lehnert. Network migration optimization using genetic algorithms. Lecture Notes in Computer Science, 7479:112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32808-4_11/.

Thongtra:2012:OBN

[539] Patcharee Thongtra, Finn Arve Aagesen, and Kornschnok Dittawit. OWL-based node capability parameter configuration. Lecture Notes in Computer Science, 7479:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_12/.

Walczyk:2012:TEC

[540] Konrad Walczyk and Annie Gravey. Techno-economic comparison of next-generation access networks for the French market. Lecture Notes in Computer Science, 7479:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_13/.

Tsiaras:2012:DSF

[541] Christos Tsiaras, Martin Waldburger, Guilherme Sperb Machado, and Andrei Vancea. The design of a single funding point charging architecture. Lecture Notes in Computer Science, 7479:148–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32808-4_14/.

Richter:2012:TLG

[542] Volker Richter, Rico Radeke, and Ralf Lehnert. A two layer guaranteed and sustained rate based scheduler for IEEE 802.16-2009 based WiMAX networks. Lecture Notes in Computer Science, 7479:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32808-4_15/.

Usach:2012:IEC

[543] Régel González Usach and Mirja Kühlewind. Implementation and evaluation of coupled congestion control for multipath TCP. Lecture Notes in Computer Science, 7479:173–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_16/.

Villa:2012:IFA

[544] Bjørn J. Villa, Poul E. Heegaard, and Anders Instefjord. Improving fairness for adaptive HTTP video streaming. Lecture Notes in Computer Science, 7479:183–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32808-4_17/.

Marszalkowski:2012:PHP

[545] Jakub Marszałkowski. Prototype of high performance scalable advertising server with local memory storage and centralised processing. Lecture Notes in Computer Science, 7479:194–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32808-4_18/.

Orosz:2012:RSM

[546] Mihály Orosz and Gábor Hosszú. Responsible source multicasting. Lecture Notes in Computer Science, 7479:204–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32808-4_19/.

Noguero:2012:MSR

[547] Adrian Noguero, Isidro Calvo, Luis Almeida, and Unai Gangoiti. A model for system resources in flexible time-triggered middleware architectures. Lecture Notes in Computer Science, 7479:215–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32808-4_20/.

Anonymous:2012:FMbc

[548] Anonymous. Front matter. Lecture Notes in Computer Science, 7479: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32808-4/1.

vanderAalst:2012:DBP

[549] Wil M. P. van der Aalst. A decade of business process management conferences: Personal reflections on a developing discipline. Lecture Notes in Computer Science, 7481:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32885-5_1/.

Buchmann:2012:CEA

[550] Alejandro Buchmann, Stefan Appel, Tobias Freudenreich, and Sebastian Frischbier. From calls to events: Architecting future BPM systems. Lecture Notes in Computer Science, 7481:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32885-5_2/.

Claes:2012:TPM

[551] Jan Claes, Irene Vanderfeesten, Hajo A. Reijers, Jakob Pinggera, and Matthias Weidlich. Tying process model quality to the modeling process: The impact of structuring, movement, and speed. Lecture Notes in Computer Science, 7481:33–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32885-5_3/.

Koehler:2012:CLM

[552] Jana Koehler, Joerg Hofstetter, and Roland Woodtly. Capabilities and levels of maturity in IT-based case management. Lecture Notes in Computer Science, 7481:49–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32885-5_4/.

Eid-Sabbagh:2012:BPA

[553] Rami-Habib Eid-Sabbagh, Remco Dijkman, and Mathias Weske. Business process architecture: Use and correctness. Lecture Notes in Computer Science, 7481:65–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32885-5_5/.

deLeoni:2012:AEL

[554] Massimiliano de Leoni, Fabrizio Maria Maggi, and Wil M. P. van der Aalst. Aligning event logs and declarative process models for conformance checking. Lecture Notes in Computer Science, 7481:82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32885-5_6/.

vanderWerf:2012:CAC

[555] Jan Martijn E. M. van der Werf, H. M. W. Verbeek, and Wil M. P. van der Aalst. Context-aware compliance checking. Lecture Notes in Computer Science, 7481:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32885-5_7/.

Banescu:2012:MPC

[556] Sebastian Banescu, Milan Petković, and Nicola Zannone. Measuring privacy compliance using fitness metrics. Lecture Notes in Computer Science, 7481:114–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32885-5_8/.

Estruch:2012:EDM

[557] Antonio Estruch and José Antonio Heredia Álvaro. Event-driven manufacturing process management approach. Lecture Notes in Computer Science, 7481:120–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32885-5_9/.

Tranquillini:2012:PBD

[558] Stefano Tranquillini, Patrik Spieß, Florian Daniel, Stamatis Karnouskos, and Fabio Casati. Process-based design and integration of wireless sensor network applications. Lecture Notes in Computer Science, 7481:

134–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32885-5_10/.

Scekic:2012:MRI

[559] Ognjen Scekic, Hong-Linh Truong, and Schahram Dustdar. Modeling rewards and incentive mechanisms for social BPM. Lecture Notes in Computer Science, 7481:150–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32885-5_11/.

Favre:2012:DRI

[560] Cédric Favre and Hagen Völzer. The difficulty of replacing an inclusive OR-join. Lecture Notes in Computer Science, 7481:156–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32885-5_12/.

Accorsi:2012:AIF

[561] Rafael Accorsi and Andreas Lehmann. Automatic information flow analysis of business process models. Lecture Notes in Computer Science, 7481: 172–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32885-5_13/.

Benner:2012:MTT

[562] Marian Benner, Matthias Book, Tobias Brückmann, Volker Gruhn, and Thomas Richter. Managing and tracing the traversal of process clouds with templates, agendas and artifacts. Lecture Notes in Computer Science, 7481:188–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32885-5_14/.

Amziani:2012:GFS

[563] Mourad Amziani, Tarek Melliti, and Samir Tata. A generic framework for service-based business process elasticity in the cloud. Lecture Notes in Computer Science, 7481:194–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32885-5_15/.

Li:2012:BPD

[564] Juan Li, Ross Jeffery, Kam Hay Fung, Liming Zhu, Qing Wang, He Zhang, and Xiwei Xu. A business process-driven approach for requirements dependency analysis. Lecture Notes in Computer Science, 7481:200–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32885-5_16/.

Martinho:2012:RAC

[565] David Martinho and António Rito Silva. A recommendation algorithm to capture end-users’ tacit knowledge. Lecture Notes in Computer Science, 7481:216–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32885-5_17/.

del-Rio-Ortega:2012:DPP

[566] Adela del Río-Ortega, Manuel Resinas Arias de Reyna, and Amador Durán Toro. Defining process performance indicators by using templates and patterns. Lecture Notes in Computer Science, 7481:223–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32885-5_18/.

Fahland:2012:RPM

[567] Dirk Fahland and Wil M. P. van der Aalst. Repairing process models to reflect reality. Lecture Notes in Computer Science, 7481:229– 245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32885-5_19/.

Yan:2012:FIA

[568] Zhiqiang Yan, Remco Dijkman, and Paul Grefen. FNet: An index for advanced business process querying. Lecture Notes in Computer Science, 7481:246–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32885-5_20/.

Anonymous:2012:FMbd

[569] Anonymous. Front matter. Lecture Notes in Computer Science, 7481: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32885-5/1.

Mahmud:2012:TAR

[570] Md. Sadim Mahmud and Sylvia L. Osborn. Tradeoff analysis of relational database storage of privacy preferences. Lecture Notes in Computer Science, 7482:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32873-2_1/.

Shoaran:2012:DPP

[571] Maryam Shoaran, Alex Thomo, and Jens Weber. Differential privacy in practice. Lecture Notes in Computer Science, 7482:14–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32873-2_2/.

Whang:2012:MQI

[572] Steven Euijong Whang and Hector Garcia-Molina. A model for quantifying information leakage. Lecture Notes in Computer Science, 7482: 25–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32873-2_3/.

Janusz:2012:PPL

[573] Daniel Janusz, Martin Kost, and Johann-Christoph Freytag. Privacy protocol for linking distributed medical data. Lecture Notes in Computer Science, 7482:45–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32873-2_4/.

Wang:2012:CCF

[574] Wen Qiang Wang, Dinh Tien Tuan Anh, Hock Beng Lim, and Anwitaman Datta. Cloud and the city: Facilitating flexible access control over datastreams. Lecture Notes in Computer Science, 7482:58–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32873-2_5/.

Carminati:2012:PCE

[575] Barbara Carminati, Elena Ferrari, and Michele Guglielmi. Policies for composed emergencies in support of disaster management. Lecture Notes in Computer Science, 7482:75–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32873-2_6/.

Hore:2012:IED

[576] Bijit Hore, Ee-Chien Chang, Mamadou H. Diallo, and Sharad Mehrotra. Indexing encrypted documents for supporting efficient keyword search. Lecture Notes in Computer Science, 7482:93–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32873-2_7/.

Zhang:2012:SQR

[577] Chunwang Zhang, Junjie Jin, Ee-Chien Chang, and Sharad Mehrotra. Secure quasi-realtime collaborative editing over low-cost storage services. Lecture Notes in Computer Science, 7482:111–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32873-2_8/.

Kozak:2012:SMB

[578] Stepan Kozak, David Novak, and Pavel Zezula. Secure metric-based index for similarity cloud. Lecture Notes in Computer Science, 7482: 130–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32873-2_9/.

Bienvenu:2012:PWD

[579] Meghyn Bienvenu, Daniel Deutch, and Fabian M. Suchanek. Provenance for Web 2.0 data. Lecture Notes in Computer Science, 7482: 148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32873-2_10/.

Thorpe:2012:TEB

[580] Sean Thorpe, Tyrone Grandison, Indrajit Ray, and Abbie Barbir. Towards enabling behavioral trust among participating cloud forensic data center agencies. Lecture Notes in Computer Science, 7482:156–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32873-2_11/.

Deng:2012:TTH

[581] Mina Deng, Marco Nalin, Milan Petković, Ilaria Baroni, and Abitabile Marco. Towards trustworthy health platform cloud. Lecture Notes in Computer Science, 7482:162–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32873-2_12/.

Anonymous:2012:BMk

[582] Anonymous. Back matter. Lecture Notes in Computer Science, 7482: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32873-2/1.

Anonymous:2012:FMbe

[583] Anonymous. Front matter. Lecture Notes in Computer Science, 7482: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32873-2/1.

Pathak:2012:PPS

[584] Manas Pathak, Jose Portelo, Bhiksha Raj, and Isabel Trancoso. Privacypreserving speaker authentication. Lecture Notes in Computer Science, 7483:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33383-5_1/.

Mendel:2012:DAR

[585] Florian Mendel, Tomislav Nad, Stefan Scherz, and Martin Schläffer. Differential attacks on reduced RIPEMD-160. Lecture Notes in Computer Science, 7483:23–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33383-5_2/.

Groza:2012:RDN

[586] Bogdan Groza and Bogdan Warinschi. Revisiting difficulty notions for client puzzles and DoS resilience. Lecture Notes in Computer Science, 7483:39–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33383-5_3/.

Kunihiro:2012:OBS

[587] Noboru Kunihiro. On optimal bounds of small inverse problems and approximate GCD problems with higher degree. Lecture Notes in Computer Science, 7483:55–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33383-5_4/.

Suoranta:2012:SAM

[588] Sanna Suoranta, André Andrade, and Tuomas Aura. Strong authentication with mobile phone. Lecture Notes in Computer Science, 7483: 70–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33383-5_5/.

Amrutkar:2012:MSI

[589] Chaitrali Amrutkar, Patrick Traynor, and Paul C. van Oorschot. Measuring SSL indicators on mobile browsers: Extended life, or end of the road? Lecture Notes in Computer Science, 7483:86–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33383-5_6/.

Bender:2012:DSP

[590] Jens Bender, Özgür Dagdelen, Marc Fischlin, and Dennis Kügler. Domainspecific pseudonymous signatures for the German identity card. Lecture Notes in Computer Science, 7483:104–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33383-5_7/.

Strenzke:2012:SSP

[591] Falko Strenzke. Solutions for the storage problem of McEliece public and private keys on memory-constrained platforms. Lecture Notes in Computer Science, 7483:120–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33383-5_8/.

Bag:2012:CLA

[592] Samiran Bag, Aritra Dhar, and Pinaki Sarkar. 100% connectivity for location aware code based KPD in clustered WSN: Merging blocks. Lecture Notes in Computer Science, 7483:136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33383-5_9/.

Zhao:2012:LFG

[593] Lei Zhao, Debin Gao, and Lina Wang. Learning fine-grained structured input for memory corruption detection. Lecture Notes in Computer Science, 7483:151–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33383-5_10/.

Alsouri:2012:DAD

[594] Sami Alsouri, Jan Sinschek, Andreas Sewe, Eric Bodden, and Mira Mezini. Dynamic anomaly detection for more trustworthy outsourced computation. Lecture Notes in Computer Science, 7483:168–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33383-5_11/.

Wang:2012:ESD

[595] Jiangang Wang, Xiaohong Li, Xuhui Liu, Xinshu Dong, Junjie Wang, and Zhenkai Liang. An empirical study of dangerous behaviors in Firefox extensions. Lecture Notes in Computer Science, 7483:188–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33383-5_12/.

Clear:2012:CPA

[596] Michael Clear, Karl Reid, Desmond Ennis, Arthur Hughes, and Hitesh Tewari. Collaboration-preserving authenticated encryption for operational transformation systems. Lecture Notes in Computer Science, 7483: 204–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33383-5_13/.

Bosch:2012:SDR

[597] Christoph Bösch, Qiang Tang, Pieter Hartel, and Willem Jonker. Selective document retrieval from encrypted database. Lecture Notes in Computer Science, 7483:224–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33383-5_14/.

Peter:2012:AHE

[598] Andreas Peter, Max Kronberg, Wilke Trei, and Stefan Katzenbeisser. Additively homomorphic encryption with a double decryption mechanism, revisited. Lecture Notes in Computer Science, 7483:242–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33383-5_15/.

Fujioka:2012:SHI

[599] Atsushi Fujioka, Taiichi Saito, and Keita Xagawa. Secure hierarchical identity-based identification without random oracles. Lecture Notes in Computer Science, 7483:258–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33383-5_16/.

Ghadafi:2012:ETM

[600] E. Ghadafi and N. P. Smart. Efficient two-move blind signatures in the common reference string model. Lecture Notes in Computer Science, 7483: 274–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33383-5_17/.

Hu:2012:CCU

[601] Jinwei Hu, Khaled M. Khan, Yun Bai, and Yan Zhang. Compliance checking for usage-constrained credentials in trust negotiation systems. Lecture Notes in Computer Science, 7483:290–305, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33383-5_18/.

Drabik:2012:QAI

[602] Peter Drábik, Fabio Martinelli, and Charles Morisset. A quantitative approach for inexact enforcement of security policies. Lecture Notes in Computer Science, 7483:306–321, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33383-5_19/.

Nassr:2012:OOS

[603] Nezar Nassr, Nidal Aboudagga, and Eric Steegmans. OSDM: An organizational supervised delegation model for RBAC. Lecture Notes in Computer Science, 7483:322–337, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33383-5_20/.

Anonymous:2012:FMbf

[604] Anonymous. Front matter. Lecture Notes in Computer Science, 7483: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33383-5/1.

Monien:2012:SDO

[605] Burkhard Monien and Christian Scheideler. Selfish distributed optimization. Lecture Notes in Computer Science, 7484:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32820-6_1.

Rana:2012:TST

[606] Omer Rana, Marios Dikaiakos, Daniel S. Katz, and Christine Morin. Topic 1: Support tools and environments. Lecture Notes in Computer Science, 7484:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-32820-6_2.

Wong:2012:TVF

[607] Yi Wen Wong, Tomasz Dubrownik, Wai Teng Tang, Wen Jun Tan, and Rubing Duan. Tulipse: a visualization framework for user-guided parallelization. Lecture Notes in Computer Science, 7484:4–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32820-6_3/.

Marozzo:2012:ECI

[608] Fabrizio Marozzo, Francesc Lordan, Roger Rafanell, Daniele Lezzi, and Domenico Talia. Enabling cloud interoperability with COMPSs. Lecture Notes in Computer Science, 7484:16–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32820-6_4/.

Calotoiu:2012:PID

[609] Alexandru Calotoiu, Christian Siebert, and Felix Wolf. Patternindependent detection of manual collectives in MPI programs. Lecture Notes in Computer Science, 7484:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32820-6_5/.

Hulette:2012:TBA

[610] Geoffrey C. Hulette, Matthew J. Sottile, and Allen D. Malony. A typebased approach to separating protocol from application logic. Lecture Notes in Computer Science, 7484:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32820-6_6/.

Malony:2012:TPP

[611] Allen D. Malony, Helen Karatza, William Knottenbelt, and Sally Mc-Kee. Topic 2: Performance prediction and evaluation. Lecture Notes in Computer Science, 7484:52–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32820-6_7.

Marowka:2012:ECM

[612] Ami Marowka. Energy consumption modeling for hybrid computing. Lecture Notes in Computer Science, 7484:54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32820-6_8/.

Aguilera:2012:HFS

[613] Alvaro Aguilera, Michael Kluge, Thomas William, and Wolfgang E. Nagel. HPC file systems in wide area networks: Understanding the performance of Lustre over WAN. Lecture Notes in Computer Science, 7484: 65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32820-6_9/.

Logan:2012:UPU

[614] Jeremy Logan, Scott Klasky, Hasan Abbasi, Qing Liu, George Ostrouchov, and Manish Parashar. Understanding I/O performance using I/O skeletal applications. Lecture Notes in Computer Science, 7484: 77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32820-6_10/.

Castro:2012:AAS

[615] Pablo de Oliveira Castro, Eric Petit, Jean Christophe Beyler, and William Jalby. ASK: Adaptive sampling kit for performance characterization. Lecture Notes in Computer Science, 7484:89–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32820-6_11/.

Jiao:2012:CPW

[616] Shuai Jiao, Paolo Ienne, Xiaochun Ye, Da Wang, Dongrui Fan, and Ninghui Sun. CRAW/P: a workload partition method for the efficient parallel simulation of manycores. Lecture Notes in Computer Science, 7484: 102–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32820-6_12/.

Trystram:2012:TSL

[617] Denis Trystram, Ioannis Milis, Zhihui Du, and Uwe Schwiegelshohn. Topic 3: Scheduling and load balancing. Lecture Notes in Computer Science, 7484:115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-32820-6_13.

Chretien:2012:JSU

[618] Stephane Chretien, Jean-Marc Nicod, Laurent Philippe, Veronika Rehn-Sonigo, and Lamiel Toch. Job scheduling using successive linear programming approximations of a sparse model. Lecture Notes in Computer Science, 7484:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32820-6_14/.

Angel:2012:SSP

[619] Eric Angel, Evripidis Bampis, Fadi Kacem, and Dimitrios Letsios. Speed scaling on parallel processors with migration. Lecture Notes in Computer Science, 7484:128–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32820-6_15/.

Narang:2012:DDS

[620] Ankur Narang, Abhinav Srivastava, Ramnik Jain, and R. K. Shyamasundar. Dynamic distributed scheduling algorithm for state space search. Lecture Notes in Computer Science, 7484:141–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32820-6_16/.

Janjic:2012:ULI

[621] Vladimir Janjic and Kevin Hammond. Using load information in workstealing on distributed systems with non-uniform communication latencies. Lecture Notes in Computer Science, 7484:155–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32820-6_17/.

Sanders:2012:EEF

[622] Peter Sanders and Jochen Speck. Energy efficient frequency scaling and scheduling for malleable tasks. Lecture Notes in Computer Science, 7484: 167–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32820-6_18/.

VeigaNeves:2012:SMJ

[623] Marcelo Veiga Neves, Tiago Ferreto, and César De Rose. Scheduling MapReduce jobs in HPC clusters. Lecture Notes in Computer Science, 7484:179–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32820-6_19/.

Utrera:2012:JSA

[624] Gladys Utrera, Siham Tabik, Julita Corbalan, and Jesús Labarta. A job scheduling approach for multi-core clusters based on virtual malleability. Lecture Notes in Computer Science, 7484:191–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32820-6_20/.

Anonymous:2012:FMbg

[625] Anonymous. Front matter. Lecture Notes in Computer Science, 7484: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32820-6/1.

Xie:2012:DPK

[626] Xiang Xie, Rui Xue, and Rui Zhang. Deterministic public key encryption and identity-based encryption from lattices in the auxiliary-input setting. Lecture Notes in Computer Science, 7485:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32928-9_1/.

Gentry:2012:RSB

[627] Craig Gentry, Shai Halevi, Chris Peikert, and Nigel P. Smart. Ring switching in BGV-style homomorphic encryption. Lecture Notes in Computer Science, 7485:19–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32928-9_2/.

Damgaard:2012:ZKP

[628] Ivan Damgård and Adriana López-Alt. Zero-knowledge proofs with low amortized communication from lattice assumptions. Lecture Notes in Computer Science, 7485:38–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32928-9_3/.

Camenisch:2012:FAA

[629] Jan Camenisch, Gregory Neven, and Markus Rückert. Fully anonymous attribute tokens from lattices. Lecture Notes in Computer Science, 7485: 57–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32928-9_4/.

Camenisch:2012:ESP

[630] Jan Camenisch, Maria Dubovitskaya, and Kristiyan Haralambiev. Efficient structure-preserving signature scheme from standard assumptions. Lecture Notes in Computer Science, 7485:76–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32928-9_5/.

Blazy:2012:CRO

[631] Olivier Blazy, David Pointcheval, and Damien Vergnaud. Compact roundoptimal partially-blind signatures. Lecture Notes in Computer Science, 7485:95–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32928-9_6/.

Fischlin:2012:HFS

[632] Marc Fischlin, Anja Lehmann, and Dominique Schröder. History-free sequential aggregate signatures. Lecture Notes in Computer Science, 7485: 113–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32928-9_7/.

Chase:2012:NHS

[633] Melissa Chase and Markulf Kohlweiss. A new hash-and-sign approach and structure-preserving signatures from DLIN. Lecture Notes in Computer Science, 7485:131–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32928-9_8/.

Myers:2012:BCM

[634] Steven Myers, Mona Sergi, and abhi shelat. Blackbox construction of a more than non-malleable CCA 1 encryption scheme from plaintext awareness. Lecture Notes in Computer Science, 7485:149–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32928-9_9/.

Phan:2012:DDB

[635] Duong Hieu Phan, David Pointcheval, and Mario Strefler. Decentralized dynamic broadcast encryption. Lecture Notes in Computer Science, 7485: 166–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32928-9_10/.

Kasamatsu:2012:TSE

[636] Kohei Kasamatsu, Takahiro Matsuda, Keita Emura, Nuttapong Attrapadung, and Goichiro Hanaoka. Time-specific encryption from forwardsecure encryption. Lecture Notes in Computer Science, 7485:184–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32928-9_11/.

Kolesnikov:2012:IST

[637] Vladimir Kolesnikov and Ranjit Kumaresan. Improved secure two-party computation via information-theoretic garbled circuits. Lecture Notes in Computer Science, 7485:205–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32928-9_12/.

Baron:2012:SPM

[638] Joshua Baron, Karim El Defrawy, Kirill Minkovich, Rafail Ostrovsky, and Eric Tressler. 5PM: Secure pattern matching. Lecture Notes in Computer Science, 7485:222–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32928-9_13/.

Damgaard:2012:IAA

[639] Ivan Damgård, Marcel Keller, Enrique Larraia, Christian Miles, and Nigel P. Smart. Implementing AES via an actively/covertly secure dishonest-majority MPC protocol. Lecture Notes in Computer Science, 7485:241–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32928-9_14/.

Choi:2012:CLC

[640] Seung Geol Choi, Dana Dachman-Soled, and Moti Yung. On the centrality of off-line E-cash to concrete partial information games. Lecture Notes in Computer Science, 7485:264–280, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32928-9_15/.

Canetti:2012:UCS

[641] Ran Canetti and Margarita Vald. Universally composable security with local adversaries. Lecture Notes in Computer Science, 7485:281–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32928-9_16/.

Yasuda:2012:SCE

[642] Masaya Yasuda, Takeshi Shimoyama, Jun Kogure, and Tetsuya Izu. On the strength comparison of the ECDLP and the IFP. Lecture Notes in Computer Science, 7485:302–325, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32928-9_17/.

Herold:2012:NAK

[643] Gottfried Herold and Alexander Meurer. New attacks for knapsack based cryptosystems. Lecture Notes in Computer Science, 7485:326– 342, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32928-9_18/.

Blondeau:2012:MDC

[644] Céline Blondeau, Benoît Gérard, and Kaisa Nyberg. Multiple differential cryptanalysis using LLR and χ^{2} statistics. Lecture Notes in Computer Science, 7485:343–360, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32928-9_19/.

Thomae:2012:QVQ

[645] Enrico Thomae. Quo vadis quaternion? Cryptanalysis of Rainbow over non-commutative rings. Lecture Notes in Computer Science, 7485: 361–373, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32928-9_20/.

Anonymous:2012:FMbh

[646] Anonymous. Front matter. Lecture Notes in Computer Science, 7485: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32928-9/1.

Gabbay:2012:BAF

[647] Dov Gabbay. Bipolar argumentation frames and contrary to duty obligations, preliminary report. Lecture Notes in Computer Science, 7486: 1–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32897-8_1/.

Lakemeyer:2012:MAO

[648] Gerhard Lakemeyer. Multi-agent only-knowing. Lecture Notes in Computer Science, 7486:25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-32897-8_2.

Lorini:2012:LRA

[649] Emiliano Lorini. Logics for reasoning about agents’ attitudes in strategic contexts. Lecture Notes in Computer Science, 7486:26, 2012.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32897-8_3.

Pinyol:2012:TSM

[650] Isaac Pinyol. A time-situated meta-logic for characterizing goal-processing bounded agents. Lecture Notes in Computer Science, 7486:27–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32897-8_4/.

Lam:2012:DDS

[651] Ho-Pun Lam, Guido Governatori, and Ken Satoh. Distributed defeasible speculative reasoning in ambient environment. Lecture Notes in Computer Science, 7486:43–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32897-8_5/.

Dignum:2012:FSA

[652] Frank Dignum and Virginia Dignum. A formal semantics for agent (re)organization. Lecture Notes in Computer Science, 7486:61–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32897-8_6/.

Guelev:2012:EAP

[653] Dimitar P. Guelev and Catalin Dima. Epistemic ATL with perfect recall, past and strategy contexts. Lecture Notes in Computer Science, 7486:77–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32897-8_7/.

Popovici:2012:UEG

[654] Matei Popovici. Using evolution graphs for describing topology-aware prediction models in large clusters. Lecture Notes in Computer Science, 7486:94–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32897-8_8/.

Mirbel:2012:EGB

[655] Isabelle Mirbel and Serena Villata. Enhancing goal-based requirements consistency: An argumentation-based approach. Lecture Notes in Computer Science, 7486:110–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32897-8_9/.

Bulling:2012:GTA

[656] Nils Bulling, Michael Köster, and Matei Popovici. A game theoretic approach for optimal network topologies in opportunistic networks. Lecture Notes in Computer Science, 7486:128–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32897-8_10/.

Homola:2012:MKB

[657] Martin Homola, Matthias Knorr, and João Leite. MKNF knowledge bases in multi-context systems. Lecture Notes in Computer Science, 7486: 146–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32897-8_11/.

Wright:2012:IRP

[658] Ben Wright, Enrico Pontelli, and Tran Cao Son. Implementing reversible processes in multi-agent action languages using answer set planning. Lecture Notes in Computer Science, 7486:163–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32897-8_12/.

Gratie:2012:FHC

[659] Cristian Gratie and Adina Magda Florea. Full hybrid μ-calculus, its bisimulation invariance and application to argumentation. Lecture Notes in Computer Science, 7486:181–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32897-8_13/.

Gabbay:2012:NAM

[660] Dov Gabbay and Odinaldo Rodrigues. A numerical approach to the merging of argumentation networks. Lecture Notes in Computer Science, 7486: 195–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32897-8_14/.

Anonymous:2012:BMl

[661] Anonymous. Back matter. Lecture Notes in Computer Science, 7486: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32897-8/1.

Anonymous:2012:FMbi

[662] Anonymous. Front matter. Lecture Notes in Computer Science, 7486: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32897-8/1.

Ortiz:2012:RQA

[663] Magdalena Ortiz and Mantas Šimkus. Reasoning and query answering in description logics. Lecture Notes in Computer Science, 7487: 1–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33158-9_1/.

Gottlob:2012:DES

[664] Georg Gottlob, Giorgio Orsi, and Andreas Pieris. Datalog and its extensions for semantic Web databases. Lecture Notes in Computer Science, 7487:54–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33158-9_2/.

Arenas:2012:FNS

[665] Marcelo Arenas and Jorge Pérez. Federation and navigation in SPARQL 1.1. Lecture Notes in Computer Science, 7487:78–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33158-9_3/.

Krotzsch:2012:OPI

[666] Markus Krötzsch. OWL 2 profiles: An introduction to lightweight ontology languages. Lecture Notes in Computer Science, 7487:112–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33158-9_4/.

Flesca:2012:ROD

[667] Sergio Flesca, Tim Furche, and Linda Oro. Reasoning and ontologies in data extraction. Lecture Notes in Computer Science, 7487: 184–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33158-9_5/.

Qi:2012:RUI

[668] Guilin Qi and Jianfeng Du. Reasoning with uncertain and inconsistent OWL ontologies. Lecture Notes in Computer Science, 7487:211– 244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33158-9_6/.

Le-Phuoc:2012:LSD

[669] Danh Le-Phuoc and Josiane Xavier Parreira. Linked stream data processing. Lecture Notes in Computer Science, 7487:245–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33158-9_7/.

Koubarakis:2012:DMQ

[670] Manolis Koubarakis and Manos Karpathiotakis. Data models and query languages for linked geospatial data. Lecture Notes in Computer Science, 7487:290–328, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33158-9_8/.

Bry:2012:SWA

[671] François Bry and Sebastian Schaffert. Semantic wikis: Approaches, applications, and perspectives. Lecture Notes in Computer Science, 7487: 329–369, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33158-9_9/.

Toni:2012:RWA

[672] Francesca Toni. Reasoning on the Web with assumption-based argumentation. Lecture Notes in Computer Science, 7487:370–386, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33158-9_10/.

Anonymous:2012:BMm

[673] Anonymous. Back matter. Lecture Notes in Computer Science, 7487: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33158-9/1.

Anonymous:2012:FMbj

[674] Anonymous. Front matter. Lecture Notes in Computer Science, 7487: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33158-9/1.

Cassidy:2012:ARC

[675] Taylor Cassidy, Heng Ji, Hongbo Deng, Jing Zheng, and Jiawei Han. Analysis and refinement of cross-lingual entity linking. Lecture Notes in Computer Science, 7488:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33247-0_1/.

Nordlie:2012:SYI

[676] Ragnar Nordlie and Nils Pharo. Seven years of INEX interactive retrieval experiments — lessons and challenges. Lecture Notes in Computer Science, 7488:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33247-0_2/.

Hanbury:2012:BAD

[677] Allan Hanbury, Henning Müller, Georg Langs, Marc André Weber, and Bjoern H. Menze. Bringing the algorithms to the data: Cloud–based benchmarking for medical image analysis. Lecture Notes in Computer Science, 7488:24–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33247-0_3/.

Jurgens:2012:GBC

[678] Julia J. Jürgens, Preben Hansen, and Christa Womser-Hacker. Going beyond CLEF-IP: The ‘reality’ for patent searchers? Lecture Notes in Computer Science, 7488:30–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33247-0_4/.

Orio:2012:MMM

[679] Nicola Orio, Cynthia C. S. Liem, Geoffroy Peeters, and Markus Schedl. MusiClef: Multimodal music tagging task. Lecture Notes in Computer Science, 7488:36–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33247-0_5/.

Berendsen:2012:GPT

[680] Richard Berendsen, Manos Tsagkias, Maarten de Rijke, and Edgar Meij. Generating pseudo test collections for learning to rank scientific articles. Lecture Notes in Computer Science, 7488:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33247-0_6/.

Piroi:2012:ELT

[681] Florina Piroi, Mihai Lupu, and Allan Hanbury. Effects of language and topic size in patent IR: An empirical study. Lecture Notes in Computer Science, 7488:54–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33247-0_7/.

Gupta:2012:CLH

[682] Parth Gupta, Alberto Barrón-Cedeño, and Paolo Rosso. Cross-language high similarity search using a conceptual thesaurus. Lecture Notes in Computer Science, 7488:67–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33247-0_8/.

Keszler:2012:AGC

[683] Anita Keszler, Levente Kovács, and Tamás Szirányi. The appearance of the giant component in descriptor graphs and its application for descriptor selection. Lecture Notes in Computer Science, 7488:76–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33247-0_9/.

Yan:2012:HMM

[684] Xueliang Yan, Guanglai Gao, Xiangdong Su, Hongxi Wei, Xueliang Zhang, and Qianqian Lu. Hidden Markov model for term weighting in verbose queries. Lecture Notes in Computer Science, 7488:82–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33247-0_10/.

Agosti:2012:DDS

[685] Maristella Agosti, Emanuele Di Buccio, Nicola Ferro, Ivano Masiero, and Simone Peruzzo. DIRECTions: Design and specification of an IR evaluation infrastructure. Lecture Notes in Computer Science, 7488: 88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33247-0_11/.

Galuscakova:2012:PFE

[686] Petra Galuščáková, Pavel Pecina, and Jan Hajič. Penalty functions for evaluation measures of unsegmented speech retrieval. Lecture Notes in Computer Science, 7488:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33247-0_12/.

Angelini:2012:CRP

[687] Marco Angelini, Nicola Ferro, Kalervo Järvelin, Heikki Keskustalo, and Ari Pirkola. Cumulated relative position: a metric for ranking evaluation. Lecture Notes in Computer Science, 7488:112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33247-0_13/.

Schaer:2012:BTT

[688] Philipp Schaer. Better than their reputation? On the reliability of relevance assessments with students. Lecture Notes in Computer Science, 7488:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33247-0_14/.

Kursten:2012:CIS

[689] Jens Kürsten and Maximilian Eibl. Comparing IR system components using beanplots. Lecture Notes in Computer Science, 7488:136–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33247-0_15.

Bhaskar:2012:LIQ

[690] Pinaki Bhaskar and Sivaji Bandyopadhyay. Language independent query focused snippet generation. Lecture Notes in Computer Science, 7488: 138–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33247-0_16/.

Pertile:2012:TCE

[691] Solange de L. Pertile and Viviane P. Moreira. A test collection to evaluate plagiarism by missing or incorrect references. Lecture Notes in Computer Science, 7488:141–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33247-0_17/.

Anonymous:2012:BMn

[692] Anonymous. Back matter. Lecture Notes in Computer Science, 7488: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33247-0/1.

Anonymous:2012:FMbk

[693] Anonymous. Front matter. Lecture Notes in Computer Science, 7488: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33247-0/1.

Khoo:2012:WWG

[694] Michael Khoo and Catherine Hall. What would ‘Google’ do? Users’ mental models of a digital library search engine. Lecture Notes in Computer Science, 7489:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33290-6_1/.

McKay:2012:EES

[695] Dana McKay, Annika Hinze, Ralf Heese, Nicholas Vanderschantz, and Claire Timpany. An exploration of ebook selection behavior in academic library collections. Lecture Notes in Computer Science, 7489: 13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33290-6_2/.

Loizides:2012:ISV

[696] Fernando Loizides. Information seekers’ visual focus during time constraint document triage. Lecture Notes in Computer Science, 7489: 25–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33290-6_3/.

Tahmasebi:2012:WWD

[697] Nina Tahmasebi, Gerhard Gossen, and Thomas Risse. Which words do you remember? Temporal properties of language use in digital archives. Lecture Notes in Computer Science, 7489:32–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33290-6_4/.

Lee:2012:TMF

[698] Dongwon Lee. Toward mobile-friendly libraries: The status quo. Lecture Notes in Computer Science, 7489:38–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33290-6_5/.

Hinze:2012:LTC

[699] Annika Hinze and David Bainbridge. Listen to tipple: Creating a mobile digital library with location-triggered audio books. Lecture Notes in Computer Science, 7489:51–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_6/.

Hinze:2012:RFP

[700] Annika Hinze and Amay Dighe. Re-finding physical documents: Extending a digital library into a human-centred workplace. Lecture Notes in Computer Science, 7489:57–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_7/.

Sweetnam:2012:UNE

[701] Mark S. Sweetnam, Maristella Agosti, Nicola Orio, Chiara Ponchia, and Christina M. Steiner. User needs for enhanced engagement with cultural heritage collections. Lecture Notes in Computer Science, 7489: 64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33290-6_8/.

Adams:2012:DLS

[702] Anne Adams and Pauline Ngimwa. Digital library sustainability and design processes. Lecture Notes in Computer Science, 7489:76–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33290-6_9/.

Dudczak:2012:CTV

[703] Adam Dudczak, Miłosz Kmieciak, and Marcin Werla. Creation of textual versions of historical documents from Polish digital libraries. Lecture Notes in Computer Science, 7489:89–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_10/.

Buchler:2012:IRT

[704] Marco Büchler, Gregory Crane, Maria Moritz, and Alison Babeu. Increasing recall for text re-use in historical documents to support research in the humanities. Lecture Notes in Computer Science, 7489: 95–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33290-6_11/.

Cui:2012:PPE

[705] Anqi Cui, Liner Yang, Dejun Hou, Min-Yen Kan, Yiqun Liu, Min Zhang, and Shaoping Ma. PrEV: Preservation Explorer and Vault for Web 2.0 user-generated content. Lecture Notes in Computer Science, 7489: 101–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33290-6_12/.

Mayer:2012:PSP

[706] Rudolf Mayer, Andreas Rauber, Martin Alexander Neumann, John Thomson, and Gonçalo Antunes. Preserving scientific processes from design to publications. Lecture Notes in Computer Science, 7489:113– 124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33290-6_13/.

SalahEldeen:2012:LMR

[707] Hany M. SalahEldeen and Michael L. Nelson. Losing my revolution: How many resources shared on social media have been lost? Lecture Notes in Computer Science, 7489:125–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_14/.

Sumbana:2012:AVD

[708] Maria Sumbana, Marcos André Gonçalves, Rodrigo Silva, Jussara Almeida, and Adriano Veloso. Automatic vandalism detection in Wikipedia with active associative classification. Lecture Notes in Computer Science, 7489:138–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_15/.

Sutton:2012:ADL

[709] Electra Sutton, Chloe Reynolds, Fredric C. Gey, and Ray R. Larson. Applying digital library technologies to nuclear forensics. Lecture Notes in Computer Science, 7489:144–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_16/.

Boland:2012:IRD

[710] Katarina Boland, Dominique Ritze, Kai Eckert, and Brigitte Mathiak. Identifying references to datasets in publications. Lecture Notes in Computer Science, 7489:150–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_17/.

Mahoui:2012:CTA

[711] M. Mahoui, C. Boston-Clay, R. Stein, and N. Tirupattur. Collaborative tagging of art digital libraries: Who should be tagging? Lecture Notes in Computer Science, 7489:162–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_18/.

DiBuccio:2012:SEL

[712] Emanuele Di Buccio, Giorgio Maria Di Nunzio, and Gianmaria Silvello. A system for exposing linguistic linked open data. Lecture Notes in Computer Science, 7489:173–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33290-6_19/.

Gartner:2012:LPR

[713] Richard Gartner. Linking the parliamentary record: a new approach to metadata for legislative proceedings. Lecture Notes in Computer Science, 7489:179–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33290-6_20/.

Anonymous:2012:FMbl

[714] Anonymous. Front matter. Lecture Notes in Computer Science, 7489: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33290-6/1.

Gropp:2012:MBW

[715] William Gropp. MPI 3 and beyond: Why MPI is successful and what challenges it faces. Lecture Notes in Computer Science, 7490:1– 9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33518-1_1/.

Bruning:2012:MFT

[716] Ulrich Brüning. MPI functions and their impact on interconnect hardware. Lecture Notes in Computer Science, 7490:10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-33518-1_2.

Sumimoto:2012:MCL

[717] Shinji Sumimoto. The MPI Communication Library for the K computer: Its design and implementation. Lecture Notes in Computer Science, 7490:11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-33518-1_3.

Danalis:2012:MCT

[718] Anthony Danalis. MPI and compiler technology: a love-hate relationship. Lecture Notes in Computer Science, 7490:12–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-33518-1_4.

Gropp:2012:AMI

[719] William Gropp, Ewing Lusk, and Rajeev Thakur. Advanced MPI including new MPI-3 features. Lecture Notes in Computer Science, 7490:14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33518-1_5.

Geimer:2012:HPH

[720] Markus Geimer, Michael Gerndt, Sameer Shende, Bert Wesarg, and Brian Wylie. Hands-on practical hybrid parallel application performance engineering. Lecture Notes in Computer Science, 7490:15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33518-1_6.

Zhao:2012:ASO

[721] Xin Zhao, Gopalakrishnan Santhanaraman, and William Gropp. Adaptive strategy for one-sided communication in MPICH2. Lecture Notes in Computer Science, 7490:16–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33518-1_7/.

Barrett:2012:LIF

[722] Brian W. Barrett, Ron Brightwell, and Keith D. Underwood. A low impact flow control implementation for offload communication interfaces. Lecture Notes in Computer Science, 7490:27–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33518-1_8/.

Didelot:2012:IMC

[723] Sylvain Didelot, Patrick Carribault, Marc Pérache, and William Jalby. Improving MPI communication overlap with collaborative polling. Lecture Notes in Computer Science, 7490:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33518-1_9/.

Yoshinaga:2012:DBM

[724] Kazumi Yoshinaga, Yuichi Tsujita, Atsushi Hori, Mikiko Sato, and Mitaro Namiki. Delegation-based MPI communications for a hybrid parallel computer with many-core architecture. Lecture Notes in Computer Science, 7490:47–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33518-1_10/.

Dinan:2012:EMC

[725] James Dinan, David Goodell, William Gropp, Rajeev Thakur, and Pavan Balaji. Efficient multithreaded context ID allocation in MPI. Lecture Notes in Computer Science, 7490:57–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33518-1_11/.

Jain:2012:CTT

[726] Nikhil Jain, JohnMark Lau, and Laxmikant Kale. Collectives on twotier direct networks. Lecture Notes in Computer Science, 7490:67–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33518-1_12/.

Venkata:2012:EAO

[727] Manjunath Gorentla Venkata, Richard L. Graham, Joshua S. Ladd, and Pavel Shamis. Exploiting atomic operations for barrier on Cray XE/XK systems. Lecture Notes in Computer Science, 7490:78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33518-1_13/.

Pellegrini:2012:EDA

[728] Simone Pellegrini, Torsten Hoefler, and Thomas Fahringer. Exact dependence analysis for increased communication overlap. Lecture Notes in Computer Science, 7490:89–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33518-1_14/.

Traff:2012:MTM

[729] Jesper Larsson Träff. mpicroscope: Towards an MPI benchmark tool for performance guideline verification. Lecture Notes in Computer Science, 7490:100–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33518-1_15/.

Bureddy:2012:OGM

[730] D. Bureddy, H. Wang, A. Venkatesh, S. Potluri, and D. K. Panda. OMB-GPU: a micro-benchmark suite for evaluating MPI libraries on GPU clusters. Lecture Notes in Computer Science, 7490:110–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33518-1_16/.

Schneider:2012:MAC

[731] Timo Schneider, Robert Gerstenberger, and Torsten Hoefler. Microapplications for communication data access patterns and MPI datatypes. Lecture Notes in Computer Science, 7490:121–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33518-1_17/.

Hoefler:2012:LMO

[732] Torsten Hoefler, James Dinan, Darius Buntinas, Pavan Balaji, and Brian W. Barrett. Leveraging MPI’s one-sided communication interface for shared-memory programming. Lecture Notes in Computer Science, 7490:132–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33518-1_18/.

Urena:2012:WFM

[733] Isaías A. Comprés Ureña, Michael Gerndt, and Carsten Trinitis. Waitfree message passing protocol for non-coherent shared memory architectures. Lecture Notes in Computer Science, 7490:142–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33518-1_19/.

Hori:2012:EKL

[734] Atsushi Hori, Toyohisa Kameyama, Yuichi Tsujita, Mitaro Namiki, and Yutaka Ishikawa. An efficient kernel-level blocking MPI implementation. Lecture Notes in Computer Science, 7490:153–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33518-1_20/.

Anonymous:2012:FMbm

[735] Anonymous. Front matter. Lecture Notes in Computer Science, 7490: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33518-1/1.

Glasmachers:2012:CIF

[736] Tobias Glasmachers. Convergence of the IGO-flow of isotropic Gaussian distributions on convex quadratic problems. Lecture Notes in Computer Science, 7491:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32937-1_1/.

Mambrini:2012:HHI

[737] Andrea Mambrini, Dirk Sudholt, and Xin Yao. Homogeneous and heterogeneous Island models for the set cover problem. Lecture Notes in Computer Science, 7491:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32937-1_2/.

Moraglio:2012:GSG

[738] Alberto Moraglio, Krzysztof Krawiec, and Colin G. Johnson. Geometric semantic genetic programming. Lecture Notes in Computer Science, 7491:21–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_3/.

Textor:2012:ENS

[739] Johannes Textor. Efficient negative selection algorithms by sampling and approximate counting. Lecture Notes in Computer Science, 7491: 32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_4/.

Akimoto:2012:CCT

[740] Youhei Akimoto, Anne Auger, and Nikolaus Hansen. Convergence of the continuous time trajectories of isotropic evolution strategies on monotonic C^{∈}-composite functions. Lecture Notes in Computer Science, 7491: 42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_5/.

Sutton:2012:PRA

[741] Andrew M. Sutton and Frank Neumann. A parameterized runtime analysis of simple evolutionary algorithms for makespan scheduling. Lecture Notes in Computer Science, 7491:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32937-1_6/.

Qian:2012:ADB

[742] Chao Qian, Yang Yu, and Zhi-Hua Zhou. On algorithm-dependent boundary case identification for problem classes. Lecture Notes in Computer Science, 7491:62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32937-1_7/.

Chotard:2012:CSS

[743] Alexandre Chotard, Anne Auger, and Nikolaus Hansen. Cumulative stepsize adaptation on linear functions. Lecture Notes in Computer Science, 7491:72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_8/.

Arnold:2012:BSC

[744] Dirk V. Arnold. On the behaviour of the (1, λ) − σ SA-ES for a constrained linear problem. Lecture Notes in Computer Science, 7491: 82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_9/.

Whitley:2012:EES

[745] Darrell Whitley, Wenxiang Chen, and Adele Howe. An empirical evaluation of O(1) steepest descent for NK-landscapes. Lecture Notes in Computer Science, 7491:92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32937-1_10/.

Urli:2012:ESC

[746] Tommaso Urli, Markus Wagner, and Frank Neumann. Experimental supplements to the computational complexity analysis of genetic programming for problems modelling isolated program semantics. Lecture Notes in Computer Science, 7491:102–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32937-1_11/.

Kotzing:2012:ABE

[747] Timo Kötzing and Hendrik Molter. ACO beats EA on a dynamic pseudo-Boolean function. Lecture Notes in Computer Science, 7491: 113–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_12/.

Brockhoff:2012:RAS

[748] Dimo Brockhoff, Manuel López-Ibáñez, Boris Naujoks, and Günter Rudolph. Runtime analysis of simple interactive evolutionary biobjective optimization algorithms. Lecture Notes in Computer Science, 7491: 123–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_13/.

Wagner:2012:PPV

[749] Markus Wagner and Frank Neumann. Parsimony pressure versus multiobjective optimization for variable length representations. Lecture Notes in Computer Science, 7491:133–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32937-1_14/.

Amelio:2012:EGB

[750] Alessia Amelio and Clara Pizzuti. An evolutionary and graph-based method for image segmentation. Lecture Notes in Computer Science, 7491: 143–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32937-1_15/.

Ugolotti:2012:RTG

[751] Roberto Ugolotti, Youssef S. G. Nashed, and Stefano Cagnoni. Realtime GPU based road sign detection and classification. Lecture Notes in Computer Science, 7491:153–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32937-1_16/.

Sikulova:2012:AEI

[752] Michaela Sikulova and Lukas Sekanina. Acceleration of evolutionary image filter design using coevolution in Cartesian GP. Lecture Notes in Computer Science, 7491:163–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32937-1_17/.

Pelikan:2012:TLS

[753] Martin Pelikan, Mark W. Hauschild, and Pier Luca Lanzi. Transfer learning, soft distance-based bias, and the hierarchical BOA. Lecture Notes in Computer Science, 7491:173–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32937-1_18/.

Thill:2012:RLT

[754] Markus Thill, Patrick Koch, and Wolfgang Konen. Reinforcement learning with N-tuples on the game Connect-4. Lecture Notes in Computer Science, 7491:184–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32937-1_19/.

Koch:2012:ESH

[755] Patrick Koch and Wolfgang Konen. Efficient sampling and handling of variance in tuning data mining models. Lecture Notes in Computer Science, 7491:195–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32937-1_20/.

Anonymous:2012:FMbn

[756] Anonymous. Front matter. Lecture Notes in Computer Science, 7491: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32937-1/1.

Deb:2012:TED

[757] Kalyanmoy Deb, Sunith Bandaru, and Cem Celal Tutum. Temporal evolution of design principles in engineering systems: Analogies with human evolution. Lecture Notes in Computer Science, 7492:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32964-7_1/.

Waldock:2012:EPI

[758] Antony Waldock and David W. Corne. Exploiting prior information in multi-objective route planning. Lecture Notes in Computer Science, 7492:11–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32964-7_2/.

Kowatari:2012:APS

[759] Naoya Kowatari, Akira Oyama, Hernán Aguirre, and Kiyoshi Tanaka. Analysis on population size and neighborhood recombination on manyobjective optimization. Lecture Notes in Computer Science, 7492:22– 31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32964-7_3/.

Handl:2012:CCM

[760] Julia Handl and Joshua Knowles. Clustering criteria in multiobjective data clustering. Lecture Notes in Computer Science, 7492:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32964-7_4/.

Ghandar:2012:EPT

[761] Adam Ghandar, Zbigniew Michalewicz, and Ralf Zurbruegg. Enhancing profitability through interpretability in algorithmic trading with a multiobjective evolutionary fuzzy system. Lecture Notes in Computer Science, 7492:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32964-7_5/.

Israel:2012:BAF

[762] Shlomo Israel and Amiram Moshaiov. Bootstrapping aggregate fitness selection with evolutionary multi-objective optimization. Lecture Notes in Computer Science, 7492:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32964-7_6/.

Peng:2012:NTP

[763] Wei Peng and Qingfu Zhang. Network topology planning using MOEA/ D with objective-guided operators. Lecture Notes in Computer Science, 7492:62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32964-7_7/.

Luong:2012:EAM

[764] Hoang N. Luong and Peter A. N. Bosman. Elitist archiving for multiobjective evolutionary algorithms: To adapt or not to adapt. Lecture Notes in Computer Science, 7492:72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32964-7_8/.

Garza-Fabre:2012:IMS

[765] Mario Garza-Fabre, Eduardo Rodriguez-Tello, and Gregorio Toscano-Pulido. An improved multiobjectivization strategy for HP model-based protein structure prediction. Lecture Notes in Computer Science, 7492: 82–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32964-7_9/.

Li:2012:MDI

[766] Hui Li, Xiaolei Su, Zongben Xu, and Qingfu Zhang. MOEA/D with iterative thresholding algorithm for sparse optimization problems. Lecture Notes in Computer Science, 7492:93–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32964-7_10/.

DiNuovo:2012:SEM

[767] Alessandro G. Di Nuovo, Giuseppe Ascia, and Vincenzo Catania. A study on evolutionary multi-objective optimization with fuzzy approximation for computational expensive problems. Lecture Notes in Computer Science, 7492:102–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32964-7_11/.

Tangpattanakul:2012:MOO

[768] Panwadee Tangpattanakul, Nicolas Jozefowiez, and Pierre Lopez. Multiobjective optimization for selecting and scheduling observations by agile Earth observing satellites. Lecture Notes in Computer Science, 7492: 112–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32964-7_12/.

Moshaiov:2012:TMC

[769] Amiram Moshaiov and Yafit Snir. Tailoring ϵ-MOEA to conceptbased problems. Lecture Notes in Computer Science, 7492:122–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32964-7_13/.

Ishibuchi:2012:RSP

[770] Hisao Ishibuchi, Naoya Akedo, and Yusuke Nojima. Recombination of similar parents in SMS-EMOA on many-objective 0/1 knapsack problems. Lecture Notes in Computer Science, 7492:132–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32964-7_14/.

Rodriguez:2012:ABC

[771] Francisco J. Rodriguez, Carlos García-Martínez, Christian Blum, and Manuel Lozano. An artificial bee colony algorithm for the unrelated parallel machines scheduling problem. Lecture Notes in Computer Science, 7492:143–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32964-7_15/.

Fernandes:2012:CPP

[772] Carlos M. Fernandes, Juan J. Merelo, and Agostinho C. Rosa. Controlling the parameters of the particle swarm optimization with a selforganized criticality model. Lecture Notes in Computer Science, 7492: 153–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32964-7_16/.

McNabb:2012:ATE

[773] Andrew McNabb and Kevin Seppi. The apiary topology: Emergent behavior in communities of particle swarms. Lecture Notes in Computer Science, 7492:164–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32964-7_17/.

Tsutsui:2012:AMG

[774] Shigeyoshi Tsutsui. ACO on multiple GPUs with CUDA for faster solution of QAPs. Lecture Notes in Computer Science, 7492:174–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32964-7_18/.

Bim:2012:FSO

[775] Jan Bim, Giorgos Karafotias, S. K. Smit, A. E. Eiben, and Evert Haasdijk. It’s fate: a self-organising evolutionary algorithm. Lecture Notes in Computer Science, 7492:185–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32964-7_19/.

Reynolds:2012:GOA

[776] Alan P. Reynolds, Asaad Abdollahzadeh, David W. Corne, Mike Christie, and Brian Davies. Guide objective assisted particle swarm optimization and its application to history matching. Lecture Notes in Computer Science, 7492:195–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32964-7_20/.

Anonymous:2012:FMbo

[777] Anonymous. Front matter. Lecture Notes in Computer Science, 7492: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32964-7/1.

Bollen:2012:CSC

[778] Lars Bollen, Hannie Gijlers, and Wouter van Joolingen. Computersupported collaborative drawing in primary school education — technical realization and empirical findings. Lecture Notes in Computer Science, 7493:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_1/.

Emmerich:2012:TCM

[779] Katharina Emmerich, Katja Neuwald, Julia Othlinghaus, Sabrina Ziebarth, and H. Ulrich Hoppe. Training conflict management in a collaborative virtual environment. Lecture Notes in Computer Science, 7493:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_2/.

Bordies:2012:RDF

[780] Osmel Bordiés, Eloy Villasclaras, Yannis Dimitriadis, and Adolfo Ruiz-Calleja. Reusability of data flow designs in complex CSCL scripts: Evaluation results from a case study. Lecture Notes in Computer Science, 7493:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_3/.

Cruz:2012:TOC

[781] Armando Cruz, António Correia, Hugo Paredes, Benjamim Fonseca, and Leonel Morgado. Towards an overarching classification model of CSCW and groupware: a socio-technical perspective. Lecture Notes in Computer Science, 7493:41–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33284-5_4/.

Nolte:2012:NUC

[782] Alexander Nolte and Michael Prilla. Normal users cooperating on process models: Is it possible at all? Lecture Notes in Computer Science, 7493:57–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_5/.

Gutierrez:2012:DSS

[783] Francisco Gutierrez, Nelson Baloian, Sergio F. Ochoa, and Gustavo Zurita. Designing the software support for partially virtual communities. Lecture Notes in Computer Science, 7493:73–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33284-5_6/.

Prilla:2012:SST

[784] Michael Prilla. Supporting social tasks of individuals: a matter of access to cooperation systems. Lecture Notes in Computer Science, 7493: 89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_7/.

Oliva:2012:CKD

[785] Gustavo A. Oliva, Francisco W. Santana, and Kleverton C. M. de Oliveira. Characterizing key developers: a case study with Apache ant. Lecture Notes in Computer Science, 7493:97–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33284-5_8/.

Magdaleno:2012:ESC

[786] Andréa M. Magdaleno, Renata M. Araujo, and Cláudia M. L. Werner. An exploratory study on collaboration understanding in software development social networks. Lecture Notes in Computer Science, 7493: 113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_9/.

Shi:2012:KFU

[787] Shi Shi, Thomas Largillier, and Julita Vassileva. Keeping up with friends’ updates on facebook. Lecture Notes in Computer Science, 7493: 121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_10/.

Weyers:2012:FMM

[788] Benjamin Weyers, Wolfram Luther, Nelson Baloian, and Jose A. Pino. Formal modeling of multi-user interfaces in cooperative work. Lecture Notes in Computer Science, 7493:129–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33284-5_11/.

Largillier:2012:UCT

[789] Thomas Largillier and Julita Vassileva. Using collective trust for group formation. Lecture Notes in Computer Science, 7493:137–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33284-5_12/.

Chounta:2012:TSA

[790] Irene-Angelica Chounta and Nikolaos Avouris. Time series analysis of collaborative activities. Lecture Notes in Computer Science, 7493: 145–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_13/.

Harrer:2012:SOC

[791] Andreas Harrer, Thomas Irgang, Norbert Sattes, and Kerstin Pfahler. SoCCR — optimistic concurrency control for the Web-based collaborative framework metafora. Lecture Notes in Computer Science, 7493: 153–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_14/.

Bollen:2012:EEM

[792] Lars Bollen, Sabrina C. Eimler, Marc Jansen, and Jan Engler. Enabling and evaluating mobile learning scenarios with multiple input channels. Lecture Notes in Computer Science, 7493:161–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33284-5_15/.

GildeLaIglesia:2012:SRS

[793] Didac Gil de La Iglesia, Marcelo Milrad, and Jesper Andersson. Software requirements to support QoS in collaborative M-learning activities. Lecture Notes in Computer Science, 7493:176–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33284-5_16/.

Kohen-Vacs:2012:SIC

[794] Dan Kohen-Vacs, Arianit Kurti, Marcelo Milrad, and Miky Ronen. Systems integration challenges for supporting cross context collaborative pedagogical scenarios. Lecture Notes in Computer Science, 7493: 184–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33284-5_17/.

Cernea:2012:TWU

[795] Daniel Cernea, Simone Mora, Alfredo Perez, Achim Ebert, Andreas Kerren, and Monica Divitini. Tangible and wearable user interfaces for supporting collaboration among emergency workers. Lecture Notes in Computer Science, 7493:192–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33284-5_18/.

daSilva:2012:CAV

[796] Sírius Thadeu Ferreira da Silva, Jonice Oliveira, and Marcos R. S. Borges. Contextual analysis of the victims’ social network for people recommendation on the emergency scenario. Lecture Notes in Computer Science, 7493:200–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33284-5_19/.

Ziebarth:2012:MMA

[797] Sabrina Ziebarth, Nils Malzahn, and H. Ulrich Hoppe. Matchballs — a multi-agent-system for ontology-based collaborative learning games. Lecture Notes in Computer Science, 7493:208–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33284-5_20/.

Anonymous:2012:BMo

[798] Anonymous. Back matter. Lecture Notes in Computer Science, 7493: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33284-5/1.

Anonymous:2012:FMbp

[799] Anonymous. Front matter. Lecture Notes in Computer Science, 7493: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33284-5/1.

Green:2012:LPL

[800] Todd J. Green, Molham Aref, and Grigoris Karvounarakis. LogicBlox, platform and language: a tutorial. Lecture Notes in Computer Science, 7494:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32925-8_1/.

Gurevich:2012:DPP

[801] Yuri Gurevich. Datalog: a perspective and the potential. Lecture Notes in Computer Science, 7494:9–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32925-8_2/.

Mugnier:2012:ERG

[802] Marie-Laure Mugnier. Existential rules: a graph-based view. Lecture Notes in Computer Science, 7494:21–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32925-8_3/.

Polleres:2012:HWD

[803] Axel Polleres. How (well) do datalog, SPARQL and RIF interplay? Lecture Notes in Computer Science, 7494:27–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32925-8_4/.

Alviano:2012:MSD

[804] Mario Alviano, Nicola Leone, Marco Manna, Giorgio Terracina, and Pierfrancesco Veltri. Magic-sets for datalog with existential quantifiers. Lecture Notes in Computer Science, 7494:31–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32925-8_5/.

Ameloot:2012:CC

[805] Tom J. Ameloot and Jan Van den Bussche. On the CRON conjecture. Lecture Notes in Computer Science, 7494:44–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32925-8_6/.

Brass:2012:ODA

[806] Stefan Brass. Order in datalog with applications to declarative output. Lecture Notes in Computer Science, 7494:56–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32925-8_7/.

Civili:2012:BCF

[807] Cristina Civili and Riccardo Rosati. A broad class of first-order rewritable tuple-generating dependencies. Lecture Notes in Computer Science, 7494: 68–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32925-8_8/.

Febbraro:2012:DDT

[808] Onofrio Febbraro, Giovanni Grasso, Nicola Leone, Kristian Reale, and Francesco Ricca. Datalog development tools. Lecture Notes in Computer Science, 7494:81–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32925-8_9/.

Gardezi:2012:QRU

[809] Jaffer Gardezi and Leopoldo Bertossi. Query rewriting using datalog for duplicate resolution. Lecture Notes in Computer Science, 7494: 86–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32925-8_10/.

Interlandi:2012:RAK

[810] Matteo Interlandi. Reasoning about knowledge in distributed systems using datalog. Lecture Notes in Computer Science, 7494:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32925-8_11/.

Kohler:2012:DDD

[811] Sven Köhler, Bertram Ludäscher, and Yannis Smaragdakis. Declarative datalog debugging for mere mortals. Lecture Notes in Computer Science, 7494:111–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32925-8_12/.

Lukasiewicz:2012:ITQ

[812] Thomas Lukasiewicz, Maria Vanina Martinez, and Gerardo I. Simari. Inconsistency-tolerant query rewriting for linear Datalog±. Lecture Notes in Computer Science, 7494:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32925-8_13/.

Marczak:2012:CAD

[813] William R. Marczak, Peter Alvaro, Neil Conway, Joseph M. Hellerstein, and David Maier. Confluence analysis for distributed programs: a model-theoretic approach. Lecture Notes in Computer Science, 7494: 135–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32925-8_14/.

Ritter:2012:BNR

[814] Daniel Ritter and Till Westmann. Business network reconstruction using datalog. Lecture Notes in Computer Science, 7494:148–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32925-8_15/.

Sacca:2012:DED

[815] Domenico Saccà and Edoardo Serra. Data exchange in datalog is mainly a matter of choice. Lecture Notes in Computer Science, 7494:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32925-8_16/.

Shaw:2012:OLS

[816] Marianne Shaw, Paraschos Koutris, Bill Howe, and Dan Suciu. Optimizing large-scale semi-naïve Datalog evaluation in Hadoop. Lecture Notes in Computer Science, 7494:165–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32925-8_17/.

Zaniolo:2012:LFC

[817] Carlo Zaniolo. Logical foundations of continuous query languages for data streams. Lecture Notes in Computer Science, 7494:177–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32925-8_18/.

Anonymous:2012:BMp

[818] Anonymous. Back matter. Lecture Notes in Computer Science, 7494: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32925-8/1.

Anonymous:2012:FMbq

[819] Anonymous. Front matter. Lecture Notes in Computer Science, 7494: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32925-8/1.

Baetens:2012:TPT

[820] Jan M. Baetens and Bernard De Baets. Topological perturbations and their effect on the dynamics of totalistic cellular automata. Lecture Notes in Computer Science, 7495:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33350-7_1/.

Das:2012:CCR

[821] Sukanta Das, Avik Chakraborty, and Biplab K. Sikdar. Counting cycles in reversible cellular automata. Lecture Notes in Computer Science, 7495:11–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_2/.

Deserable:2012:PML

[822] Dominique Désérable. Propagative mode in a lattice-grain CA: Time evolution and timestep synchronization. Lecture Notes in Computer Science, 7495:20–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_3/.

Ishida:2012:LCC

[823] Toshikazu Ishida and Shuichi Inokuchi. Limit cycle for composited cellar automata. Lecture Notes in Computer Science, 7495:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33350-7_4/.

Kutrib:2012:IAL

[824] Martin Kutrib and Andreas Malcher. Iterative arrays: Little resources big size impact. Lecture Notes in Computer Science, 7495: 42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_5/.

Mukhopadhyay:2012:GEG

[825] Debdeep Mukhopadhyay. Generating expander graphs using cellular automata. Lecture Notes in Computer Science, 7495:52–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33350-7_6/.

Naskar:2012:ART

[826] Nazma Naskar, Avik Chakraborty, Pradipta Maji, and Sukanta Das. Analysis of reachability tree for identification of cyclic and acyclic CA states. Lecture Notes in Computer Science, 7495:63–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33350-7_7/.

Ohi:2012:CLD

[827] Fumio Ohi and Takanori Ichikawa. Confliction-like dynamics of Rule 20 ECA of Wolfram Class II. Lecture Notes in Computer Science, 7495: 73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_8/.

Adachi:2012:DCT

[828] Susumu Adachi, Jia Lee, Ferdinand Peper, Teijiro Isokawa, and Katsunobu Imai. Determining the critical temperature of the continuousstate game of life. Lecture Notes in Computer Science, 7495:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33350-7_9/.

Gwizdalla:2012:DDI

[829] Tomasz M. Gwizdałła. The dynamics of disproportionality index for cellular automata based sociophysical models. Lecture Notes in Computer Science, 7495:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33350-7_10/.

Maignan:2012:STA

[830] Luidnel Maignan and Jean-Baptiste Yunès. A spatio-temporal algorithmic point of view on firing squad synchronisation problem. Lecture Notes in Computer Science, 7495:101–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33350-7_11/.

Oliveira:2012:CAC

[831] Gina M. B. Oliveira and Paulo M. Vidica. A coevolutionary approach to cellular automata-based task scheduling. Lecture Notes in Computer Science, 7495:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33350-7_12/.

Piwonska:2012:SCA

[832] Anna Piwonska, Franciszek Seredynski, and Miroslaw Szaban. Searching cellular automata rules for solving two-dimensional binary classification problem. Lecture Notes in Computer Science, 7495:121–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33350-7_13/.

Sidiropoulos:2012:MOC

[833] Epaminondas Sidiropoulos. Multi-objective cellular automata optimization. Lecture Notes in Computer Science, 7495:131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33350-7_14/.

Scarlatos:2012:BSD

[834] Stylianos Scarlatos. Behavior of social dynamical models I: Fixation in the symmetric cyclic system. Lecture Notes in Computer Science, 7495: 141–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_15/.

Adamopoulos:2012:BSD

[835] Adam Adamopoulos and Stylianos Scarlatos. Behavior of social dynamical models II: Clustering for some multitype particle systems with confidence threshold. Lecture Notes in Computer Science, 7495:151– 160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_16/.

Sharifulina:2012:ISP

[836] Anastasiya Sharifulina. Investigation of stable patterns formed by totalistic cellular automata evolution. Lecture Notes in Computer Science, 7495: 161–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_17/.

Umeo:2012:RDC

[837] Hiroshi Umeo and Keisuke Kubo. Recent developments in constructing square synchronizers. Lecture Notes in Computer Science, 7495: 171–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33350-7_18/.

TranconyWidemann:2012:SOS

[838] Baltasar Trancón y Widemann. Structural operational semantics for Cellular Automata. Lecture Notes in Computer Science, 7495:184–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33350-7_19/.

Zawidzki:2012:COB

[839] Machi Zawidzki and Katsuhiro Nishinari. Controlling the opacity of a building envelope by a triangular two-color two-dimensional cellular automaton. Lecture Notes in Computer Science, 7495:194–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33350-7_20/.

Anonymous:2012:FMbr

[840] Anonymous. Front matter. Lecture Notes in Computer Science, 7495: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33350-7/1.

Abe:2012:TBG

[841] Masayuki Abe. Tools over bilinear groups for modular design of cryptographic tasks. Lecture Notes in Computer Science, 7496:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33272-2_1.

Liu:2012:OMC

[842] Dennis Y. W. Liu and Duncan S. Wong. One-move convertible nominative signature in the standard model. Lecture Notes in Computer Science, 7496:2–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33272-2_2/.

Zeng:2012:ERO

[843] Shengke Zeng, Zhiguang Qin, Qing Lu, and Qinyi Li. Efficient and random oracle-free conditionally anonymous ring signature. Lecture Notes in Computer Science, 7496:21–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_3/.

Selvi:2012:IBS

[844] S. Sharmila Deva Selvi, S. Sree Vivek, Dhinakaran Vinayagamurthy, and C. Pandu Rangan. ID based signcryption scheme in standard model. Lecture Notes in Computer Science, 7496:35–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33272-2_4/.

Chen:2012:CPK

[845] Cheng Chen, Jie Chen, Hoon Wei Lim, Zhenfeng Zhang, and Dengguo Feng. Combined public-key schemes: The case of ABE and ABS. Lecture Notes in Computer Science, 7496:53–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33272-2_5/.

Boureanu:2012:SWB

[846] Ioana Boureanu and Serge Vaudenay. Several weak bit-commitments using seal-once tamper-evident devices. Lecture Notes in Computer Science, 7496:70–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33272-2_6/.

Niitsoo:2012:DRO

[847] Margus Niitsoo. Deterministic random oracles. Lecture Notes in Computer Science, 7496:88–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33272-2_7/.

Ciobotaru:2012:NEU

[848] Oana Ciobotaru. On the (non-)equivalence of UC security notions. Lecture Notes in Computer Science, 7496:104–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33272-2_8/.

Yang:2012:LUC

[849] Bo Yang and Mingwu Zhang. LR-UESDE: a continual-leakage resilient encryption with unbounded extensible set delegation. Lecture Notes in Computer Science, 7496:125–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_9/.

Chen:2012:AIB

[850] Yu Chen, Zongyang Zhang, Dongdai Lin, and Zhenfu Cao. Anonymous identity-based hash proof system and its applications. Lecture Notes in Computer Science, 7496:143–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_10/.

Zhang:2012:EEF

[851] Yunmei Zhang, Joseph K. Liu, Xinyi Huang, Man Ho Au, and Willy Susilo. Efficient escrow-free identity-based signature. Lecture Notes in Computer Science, 7496:161–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_11/.

Nishioka:2012:PKP

[852] Mototsugu Nishioka. Perfect keyword privacy in PEKS systems. Lecture Notes in Computer Science, 7496:175–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33272-2_12/.

Pandit:2012:EFS

[853] Tapas Pandit and Rana Barua. Efficient fully secure attribute-based encryption schemes for general access structures. Lecture Notes in Computer Science, 7496:193–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33272-2_13/.

Yoshino:2012:SIP

[854] Masayuki Yoshino, Noboru Kunihiro, Ken Naganuma, and Hisayoshi Sato. Symmetric inner-product predicate encryption based on three groups. Lecture Notes in Computer Science, 7496:215–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33272-2_14/.

Suga:2012:SKS

[855] Takanori Suga, Takashi Nishide, and Kouichi Sakurai. Secure keyword search using Bloom filter with specified character positions. Lecture Notes in Computer Science, 7496:235–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_15/.

Chen:2012:FSD

[856] Cheng Chen, Zhenfeng Zhang, and Dengguo Feng. Fully secure doublyspatial encryption under simple assumptions. Lecture Notes in Computer Science, 7496:253–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33272-2_16/.

Yang:2012:SAK

[857] Zheng Yang and Jörg Schwenk. Strongly authenticated key exchange protocol from bilinear groups without random oracles. Lecture Notes in Computer Science, 7496:264–275, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_17/.

Guo:2012:AKE

[858] Yanfei Guo and Zhenfeng Zhang. Authenticated key exchange with entities from different settings and varied groups. Lecture Notes in Computer Science, 7496:276–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33272-2_18/.

Reyhanitabar:2012:CHD

[859] Mohammad Reza Reyhanitabar and Willy Susilo. On capabilities of hash domain extenders to preserve enhanced security properties. Lecture Notes in Computer Science, 7496:288–299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33272-2_19/.

Tang:2012:RSS

[860] Zhaohui Tang, Hoon Wei Lim, and Huaxiong Wang. Revisiting a secret sharing approach to network codes. Lecture Notes in Computer Science, 7496:300–317, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33272-2_20/.

Anonymous:2012:BMq

[861] Anonymous. Back matter. Lecture Notes in Computer Science, 7496: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33272-2/1.

Anonymous:2012:FMbs

[862] Anonymous. Front matter. Lecture Notes in Computer Science, 7496: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33272-2/1.

Brewka:2012:MCS

[863] Gerhard Brewka. Multi-context systems: Specifying the interaction of knowledge bases declaratively. Lecture Notes in Computer Science, 7497:1–4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33203-6_1/.

DiNoia:2012:SMR

[864] Tommaso Di Noia. Semantic matchmaking and ranking: Beyond deduction in retrieval scenarios. Lecture Notes in Computer Science, 7497:5– 8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33203-6_2/.

Borgwardt:2012:TAF

[865] Stefan Borgwardt and Rafael Peñaloza. A tableau algorithm for fuzzy description logics over residuated De Morgan lattices. Lecture Notes in Computer Science, 7497:9–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33203-6_3/.

Calvanese:2012:OBG

[866] Diego Calvanese, Giuseppe De Giacomo, Domenico Lembo, Marco Montali, and Ario Santoso. Ontology-based governance of data-aware processes. Lecture Notes in Computer Science, 7497:25–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33203-6_4/.

Civili:2012:QPE

[867] Cristina Civili and Riccardo Rosati. Query patterns for existential rules. Lecture Notes in Computer Science, 7497:42–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33203-6_5/.

Eiter:2012:SID

[868] Thomas Eiter, Michael Fink, and Daria Stepanova. Semantic independence in DL-programs. Lecture Notes in Computer Science, 7497: 58–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33203-6_6/.

Gutierrez-Basulto:2012:UQA

[869] Víctor Gutiérrez-Basulto, Yazmín Angélica Ibáñez-García, and Roman Kontchakov. An update on query answering with restricted forms of negation. Lecture Notes in Computer Science, 7497:75–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33203-6_7/.

Gutierrez-Basulto:2012:TUA

[870] Víctor Gutiérrez-Basulto and Szymon Klarman. Towards a unifying approach to representing and querying temporal data in description logics. Lecture Notes in Computer Science, 7497:90–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33203-6_8/.

Janhunen:2012:MPA

[871] Tomi Janhunen and Vesa Luukkala. Meta programming with answer sets for smart spaces. Lecture Notes in Computer Science, 7497:106– 121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33203-6_9/.

Konig:2012:SCB

[872] Mélanie König, Michel Leclère, Marie-Laure Mugnier, and Michaël Thomazo. A sound and complete backward chaining algorithm for existential rules. Lecture Notes in Computer Science, 7497:122–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33203-6_10/.

Liang:2012:DPS

[873] Senlin Liang and Michael Kifer. Deriving predicate statistics for logic rules. Lecture Notes in Computer Science, 7497:139–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33203-6_11/.

Lukasiewicz:2012:CAP

[874] Thomas Lukasiewicz, Maria Vanina Martinez, and Gerardo I. Simari. Consistent answers in probabilistic Datalog± ontologies. Lecture Notes in Computer Science, 7497:156–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33203-6_12/.

Tsalapati:2012:QRU

[875] Eleni Tsalapati, Giorgos Stoilos, Giorgos Stamou, and George Koletsos. Query rewriting under ontology contraction. Lecture Notes in Computer Science, 7497:172–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33203-6_13/.

Umbrich:2012:IRL

[876] Jürgen Umbrich, Aidan Hogan, Axel Polleres, and Stefan Decker. Improving the recall of live linked data querying through reasoning. Lecture Notes in Computer Science, 7497:188–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33203-6_14/.

Barilaro:2012:DOB

[877] Rosamaria Barilaro, Nicola Leone, Francesco Ricca, and Giorgio Terracina. Distributed ontology based data access via logic programming. Lecture Notes in Computer Science, 7497:205–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33203-6_15/.

Evripidou:2012:AVI

[878] Valentinos Evripidou and Francesca Toni. Argumentation and voting for an intelligent user empowering business directory on the Web. Lecture Notes in Computer Science, 7497:209–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33203-6_16/.

Faber:2012:PRA

[879] Wolfgang Faber. A polynomial reduction from ASPDA to ASP. Lecture Notes in Computer Science, 7497:213–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33203-6_17/.

Franconi:2012:DAP

[880] Enrico Franconi, Nhung Ngo, and Evgeny Sherkhonov. The definability abduction problem for data exchange. Lecture Notes in Computer Science, 7497:217–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33203-6_18/.

Hu:2012:RBC

[881] Beibei Hu, Theodore Patkos, Abdelghani Chibani, and Yacine Amirat. Rule-based context assessment in smart cities. Lecture Notes in Computer Science, 7497:221–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33203-6_19/.

Knorr:2012:RAI

[882] Matthias Knorr, David Carral Martínez, Pascal Hitzler, and Adila A. Krisnadhi. Recent advances in integrating OWL and rules (technical communication). Lecture Notes in Computer Science, 7497:225–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33203-6_20/.

Anonymous:2012:FMbt

[883] Anonymous. Front matter. Lecture Notes in Computer Science, 7497: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33203-6/1.

Rushby:2012:VSO

[884] John Rushby. The versatile synchronous observer. Lecture Notes in Computer Science, 7498:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33296-8_1.

Schulte:2012:TYA

[885] Wolfram Schulte. Thirteen years of automated code analysis at Microsoft. Lecture Notes in Computer Science, 7498:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-33296-8_2.

Acosta:2012:MCP

[886] Araceli Acosta and Cecilia Kilmurray. Model checking propositional Deontic temporal logic via a μ-calculus characterization. Lecture Notes in Computer Science, 7498:3–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33296-8_3/.

Barbosa:2012:AUB

[887] Haniel Barbosa and David Déharbe. An approach using the B method to formal verification of PLC programs in an industrial setting. Lecture Notes in Computer Science, 7498:19–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33296-8_4/.

Braz:2012:PIS

[888] Fernando A. F. Braz and Jader S. Cruz. Palytoxin inhibits the sodiumpotassium pump — an investigation of an electrophysiological model using probabilistic model checking. Lecture Notes in Computer Science, 7498:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33296-8_5/.

deMatos:2012:BBB

[889] Ernesto C. B. de Matos. BETA: a B based testing approach. Lecture Notes in Computer Science, 7498:51–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33296-8_6/.

Carvalho:2012:PAB

[890] Gustavo Carvalho and Diogo Falcão. A process algebra based strategy for generating test vectors from SCR specifications. Lecture Notes in Computer Science, 7498:67–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33296-8_7/.

Cavalheiro:2012:SPP

[891] Simone André da Costa Cavalheiro. Specification patterns for properties over reachable states of graph grammars. Lecture Notes in Computer Science, 7498:83–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33296-8_8/.

Davies:2012:CRM

[892] Jim Davies, Jeremy Gibbons, and David Milward. Compositionality and refinement in model-driven engineering. Lecture Notes in Computer Science, 7498:99–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33296-8_9/.

Didier:2012:IHF

[893] André Didier and Alexandre Mota. Identifying hardware failures systematically. Lecture Notes in Computer Science, 7498:115–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33296-8_10/.

Huang:2012:ITP

[894] Yanhong Huang, Yongxin Zhao, and Jianqi Shi. Investigating time properties of interrupt-driven programs. Lecture Notes in Computer Science, 7498:131–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33296-8_11/.

Regis:2012:SVD

[895] Germán Regis and Nicolás Ricci. Specifying and verifying declarative fluent temporal logic properties of workflows. Lecture Notes in Computer Science, 7498:147–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33296-8_12/.

Schulz:2012:CMT

[896] Christoph Schulz and Michael Löwe. Composition of model transformations: A Categorical framework. Lecture Notes in Computer Science, 7498: 163–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33296-8_13/.

Sekerinski:2012:VRE

[897] Emil Sekerinski and Tian Zhang. Verification rules for exception handling in Eiffel. Lecture Notes in Computer Science, 7498:179–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33296-8_14/.

Sharma:2012:SRP

[898] Subodh Sharma and Ganesh Gopalakrishnan. A sound reduction of persistent-sets for deadlock detection in MPI applications. Lecture Notes in Computer Science, 7498:194–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33296-8_15/.

Zanarini:2012:ATT

[899] Dante Zanarini, Carlos Luna, and Luis Sierra. Alternating-time temporal logic in the calculus of (co)inductive constructions. Lecture Notes in Computer Science, 7498:210–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33296-8_16/.

Anonymous:2012:BMr

[900] Anonymous. Back matter. Lecture Notes in Computer Science, 7498: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33296-8/1.

Anonymous:2012:FMbu

[901] Anonymous. Front matter. Lecture Notes in Computer Science, 7498: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33296-8/1.

Anonymous:2012:FMbv

[902] Anonymous. Front matter. Lecture Notes in Computer Science, 7499: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32790-2/1/1.

Kilgarriff:2012:GKY

[903] Adam Kilgarriff. Getting to know your corpus. Lecture Notes in Computer Science, 7499:3–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32790-2_1/.

Mitkov:2012:CRW

[904] Ruslan Mitkov, Richard Evans, Constantin Orăsan, Iustin Dornescu, and Miguel Rios. Coreference resolution: To What Extent Does It Help NLP Applications? Lecture Notes in Computer Science, 7499: 16–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32790-2_2/.

Anonymous:2012:FMbw

[905] Anonymous. Front matter. Lecture Notes in Computer Science, 7499: 29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32790-2/2/1.

Kobylinski:2012:SSF

[906] Łukasz Kobyliński and Mateusz Kopeć. Semantic similarity functions in Word Sense Disambiguation. Lecture Notes in Computer Science, 7499: 31–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32790-2_3/.

Scholz:2012:OMG

[907] Thomas Scholz, Stefan Conrad, and Lutz Hillekamps. Opinion mining on a German corpus of a media response analysis. Lecture Notes in Computer Science, 7499:39–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32790-2_4/.

Pinto:2012:SPA

[908] David Pinto, Darnes Vilariño, Yuridiana Alemán, Helena Gómez, and Nahun Loya. The soundex phonetic algorithm revisited for SMS Text Representation. Lecture Notes in Computer Science, 7499:47–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32790-2_5/.

Sevcikova:2012:SMA

[909] Magda Ševčíková and Jiří Mírovský. Sentence modality assignment in the Prague Dependency Treebank. Lecture Notes in Computer Science, 7499:56–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32790-2_6/.

Kokkinakis:2012:LDI

[910] Dimitrios Kokkinakis, Markus Forsberg, Sofie Johansson Kokkinakis, and Frida Smith. Literacy demands and information to cancer patients. Lecture Notes in Computer Science, 7499:64–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32790-2_7/.

Wawer:2012:EOA

[911] Aleksander Wawer and Konrad Gołuchowski. Expanding opinion attribute lexicons. Lecture Notes in Computer Science, 7499:72–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32790-2_8/.

Radziszewski:2012:TGT

[912] Adam Radziszewski and Szymon Acedański. Taggers gonna tag: An Argument against Evaluating disambiguation capacities of morphosyntactic taggers. Lecture Notes in Computer Science, 7499:81–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32790-2_9/.

Wolinski:2012:AAT

[913] Marcin Woliński and Andrzej Zaborowski. An ambiguity aware Treebank search tool. Lecture Notes in Computer Science, 7499:88–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32790-2_10/.

Petasis:2012:NAT

[914] Georgios Petasis and Mara Tsoumari. A new annotation tool for aligned bilingual corpora. Lecture Notes in Computer Science, 7499:95–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32790-2_11/.

Saric:2012:OSB

[915] Frane Šarić, Jan Šnajder, and Bojana Dalbelo Bašić. Optimizing sentence boundary detection for croatian. Lecture Notes in Computer Science, 7499:105–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32790-2_12/.

Gralinski:2012:MWI

[916] Filip Graliński. Mining the Web for idiomatic expressions Using Metalinguistic Markers. Lecture Notes in Computer Science, 7499:112–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32790-2_13/.

Krasnowska:2012:UTT

[917] Katarzyna Krasnowska, Witold Kieraś, Marcin Woliński, and Adam Przepiórkowski. Using tree transducers for detecting errors in a Treebank of Polish. Lecture Notes in Computer Science, 7499:119–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32790-2_14/.

Jelinek:2012:CMA

[918] Tomáš Jelínek, Barbora Štindlová, Alexandr Rosen, and Jirka Hana. Combining manual and automatic annotation of a Learner Corpus. Lecture Notes in Computer Science, 7499:127–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32790-2_15/.

Kiss:2012:MAC

[919] Márton Kiss, Ágoston Nagy, Veronika Vincze, Attila Almási, Zoltán Alexin, and János Csirik. A manually annotated corpus of pharmaceutical patents. Lecture Notes in Computer Science, 7499:135–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32790-2_16/.

Radziszewski:2012:LSE

[920] Adam Radziszewski and Adam Pawlaczek. Large-scale experiments with NP chunking of Polish. Lecture Notes in Computer Science, 7499: 143–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32790-2_17/.

Slavcheva:2012:MLS

[921] Milena Slavcheva. Mapping a lexical semantic resource to a common framework of computational lexicons. Lecture Notes in Computer Science, 7499:150–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32790-2_18/.

Kettnerova:2012:RBA

[922] Václava Kettnerová, Markéta Lopatková, and Zdeňka Urešová. The rulebased approach to Czech Grammaticalized Alternations. Lecture Notes in Computer Science, 7499:158–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32790-2_19/.

Glavas:2012:SSA

[923] Goran Glavaš, Jan Šnajder, and Bojana Dalbelo Bašić. Semi-supervised acquisition of Croatian Sentiment Lexicon. Lecture Notes in Computer Science, 7499:166–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32790-2_20/.

Anonymous:2012:FMbx

[924] Anonymous. Front matter. Lecture Notes in Computer Science, 7499: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32790-2/1.

Hameurlain:2012:DMC

[925] Abdelkader Hameurlain, Farookh Khadeer Hussain, Franck Morvan, and A. Min Tjoa, editors. Data Management in Cloud, Grid and P2P Systems: 5th International Conference, Globe 2012, Vienna, Austria, September 5– 6, 2012. Proceedings, volume 7450 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32343-X (print), 3-642-32344-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32344-7.

Bohm:2012:ITB

[926] Christian Böhm, Sami Khuri, Lenka Lhotská, and M. Elena Renda, editors. Information Technology in Bio- and Medical Informatics: Third International Conference, ITBAM 2012, Vienna, Austria, September 4– 5, 2012. Proceedings, volume 7451 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32394-4 (print), 3-642-32395-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32395-9.

Ko:2012:ADG

[927] Andrea Kő, Christine Leitner, Herbert Leitold, and Alexander Prosser, editors. Advancing Democracy, Government and Governance: Joint International Conference on Electronic Government and the Information Systems Perspective, and Electronic Democracy, EGOVIS/EDEM 2012, Vienna, Austria, September 3–6, 2012. Proceedings, volume 7452 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32700-1 (print), 3-642-32701-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32701-8.

Auweter:2012:IKT

[928] Axel Auweter, Dieter Kranzlmüller, Amirreza Tahamtan, and A. Min Tjoa, editors. ICT as Key Technology against Global Warming: Second International Conference, ICT-GLOW 2012, Vienna, Austria, September 6, 2012. Proceedings, volume 7453 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32605-6 (print), 3-642-32606-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32606-6.

Koutny:2012:CCT

[929] Maciej Koutny and Irek Ulidowski, editors. CONCUR 2012 — Concurrency Theory: 23rd International Conference, CONCUR 2012, Newcastle upon Tyne, UK, September 4–7, 2012. Proceedings, volume 7454 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32939-X (print), 3-642-32940-3 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32940-1.

Rahwan:2012:PPP

[930] Iyad Rahwan, Wayne Wobcke, Sandip Sen, and Toshiharu Sugawara, editors. PRIMA 2012: Principles and Practice of Multi-Agent Systems: 15th International Conference, Kuching, Sarawak, Malaysia, September 3–7, 2012. Proceedings, volume 7455 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32728-1 (print), 3-642-32729-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32729-2.

Ong:2012:LLI

[931] Luke Ong and Ruy de Queiroz, editors. Logic, Language, Information and Computation: 19th International Workshop, WoLLIC 2012, Buenos Aires, Argentina, September 3–6, 2012. Proceedings, volume 7456 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32620-X (print), 3-642-32621-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32621-9.

Richards:2012:KMA

[932] Deborah Richards and Byeong Ho Kang, editors. Knowledge Management and Acquisition for Intelligent Systems: 12th Pacific Rim Knowledge Acquisition Workshop, PKAW 2012, Kuching, Malaysia, September 5–6, 2012. Proceedings, volume 7457 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32540-8 (print), 3-642-32541-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32541-0.

Anthony:2012:PTA

[933] Patricia Anthony, Mitsuru Ishizuka, and Dickson Lukose, editors. PRICAI 2012: Trends in Artificial Intelligence: 12th Pacific Rim International Conference on Artificial Intelligence, Kuching, Malaysia, September 3– 7, 2012. Proceedings, volume 7458 of Lecture Notes in Computer Science.

Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32694-3 (print), 3-642-32695-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32695-0.

Foresti:2012:CSE

[934] Sara Foresti, Moti Yung, and Fabio Martinelli, editors. Computer Security — ESORICS 2012: 17th European Symposium on Research in Computer Security, Pisa, Italy, September 10–12, 2012. Proceedings, volume 7459 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33166-1 (print), 3-642-33167-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33167-1.

Mine:2012:SAI

[935] Antoine Miné and David Schmidt, editors. Static Analysis: 19th International Symposium, SAS 2012, Deauville, France, September 11–13, 2012. Proceedings, volume 7460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33124-6 (print), 3-642-33125-4 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33125-1.

Dorigo:2012:SII

[936] Marco Dorigo, Mauro Birattari, Christian Blum, Anders Lyhne Christensen, Andries P. Engelbrecht, Roderich Groß, and Thomas Stützle, editors. Swarm Intelligence: 8th International Conference, ANTS 2012, Brussels, Belgium, September 12–14, 2012. Proceedings, volume 7461 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32649-8 (print), 3-642-32650-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32650-9.

Balzarotti:2012:RAI

[937] Davide Balzarotti, Salvatore J. Stolfo, and Marco Cova, editors. Research in Attacks, Intrusions, and Defenses: 15th International Symposium, RAID 2012, Amsterdam, The Netherlands, September 12–14, 2012. Proceedings, volume 7462 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33337-0 (print), 3-642-33338-9 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33338-5.

Stachniss:2012:SCV

[938] Cyrill Stachniss, Kerstin Schill, and David Uttal, editors. Spatial Cognition VIII: International Conference, Spatial Cognition 2012, Kloster Seeon, Germany, August 31–September 3, 2012. Proceedings, volume 7463 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32731-1 (print), 3-642-32732-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32732-2.

Rovan:2012:MFC

[939] Branislav Rovan, Vladimiro Sassone, and Peter Widmayer, editors. Mathematical Foundations of Computer Science 2012: 37th International Symposium, MFCS 2012, Bratislava, Slovakia, August 27–31, 2012. Proceedings, volume 7464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32588-2 (print), 3-642-32589-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-32589-2.

Quirchmayr:2012:MRP

[940] Gerald Quirchmayr, Josef Basl, Ilsun You, Lida Xu, and Edgar Weippl, editors. Multidisciplinary Research and Practice for Information Systems: IFIP WG 8.4, 8.9/TC 5 International Cross-Domain Conference and Workshop on Availability, Reliability, and Security, CD-ARES 2012, Prague, Czech Republic, August 20–24, 2012. Proceedings, volume 7465 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32497-5 (print), 3-642-32498-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32498-7.

Agudo:2012:CBR

[941] Belén Díaz Agudo and Ian Watson, editors. Case-Based Reasoning Research and Development: 20th International Conference, ICCBR 2012, Lyon, France, September 3–6, 2012. Proceedings, volume 7466 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32985-3 (print), 3-642-32986-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32986-9.

Luo:2012:CDV

[942] Yuhua Luo, editor. Cooperative Design, Visualization, and Engineering: 9th International Conference, CDVE 2012, Osaka, Japan, September 2– 5, 2012. Proceedings, volume 7467 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32608-0 (print), 3-642-32609-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32609-7.

Magnusson:2012:HAI

[943] Charlotte Magnusson, Delphine Szymczak, and Stephen Brewster, editors. Haptic and Audio Interaction Design: 7th International Conference, HAID 2012, Lund, Sweden, August 23–24, 2012. Proceedings, volume 7468 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32795-8 (print), 3-642-32796-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32796-4.

Andreev:2012:ITS

[944] Sergey Andreev, Sergey Balandin, and Yevgeni Koucheryavy, editors. Internet of Things, Smart Spaces, and Next Generation Networking: 12th International Conference, NEW2AN 2012, and 5th Conference, ruS-MART 2012, St. Petersburg, Russia, August 27–29, 2012. Proceedings, volume 7469 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32685-4 (print), 3-642-32686-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-32686-8.

Gibbons:2012:GIP

[945] Jeremy Gibbons, editor. Generic and Indexed Programming: International Spring School, SSGIP 2010, Oxford, UK, March 22–26, 2010, Revised Lectures, volume 7470 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32201-8 (print), 3-642-32202-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32202-0.

Beer:2012:AEG

[946] Martin Beer, Cyril Brom, Frank Dignum, and Von-Wun Soo, editors. Agents for Educational Games and Simulations: International Workshop, AEGS 2011, Taipei, Taiwan, May 2, 2011. Revised Papers, volume 7471 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32325-1 (print), 3-642-32326-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32326-3.

Atzmueller:2012:MMU

[947] Martin Atzmueller, Alvin Chin, Denis Helic, and Andreas Hotho, editors. Modeling and Mining Ubiquitous Social Media: International Workshops MSM 2011, Boston, MA, USA, October 9, 2011, and MUSE 2011, Athens, Greece, September 5, 2011, Revised Selected Papers, volume 7472 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33683-3 (print), 3-642-33684-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33684-3.

Liu:2012:ICA

[948] Baoxiang Liu, Maode Ma, and Jincai Chang, editors. Information Computing and Applications: Third International Conference, ICICA 2012, Chengde, China, September 14–16, 2012. Proceedings, volume 7473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34061-X (print), 3-642-34062-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34062-8.

Dellaert:2012:OLS

[949] Frank Dellaert, Jan-Michael Frahm, Marc Pollefeys, Laura Leal-Taixé, and Bodo Rosenhahn, editors. Outdoor and Large-Scale Real-World Scene Analysis: 15th International Workshop on Theoretical Foundations of Computer Vision, Dagstuhl Castle, Germany, June 26 — July 1, 2011. Revised Selected Papers, volume 7474 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34090-3 (print), 3-642-34091-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34091-8.

Pinz:2012:PRJ

[950] Axel Pinz, Thomas Pock, Horst Bischof, and Franz Leberl, editors. Pattern Recognition: Joint 34th DAGM and 36th OAGM Symposium, Graz, Austria, August 28–31, 2012. Proceedings, volume 7476 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32716-8 (print), 3-642-32717-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32717-9.

Mana:2012:ANN

[951] Nadia Mana, Friedhelm Schwenker, and Edmondo Trentin, editors. Artificial Neural Networks in Pattern Recognition: 5th INNS IAPR TC 3 GIRPR Workshop, ANNPR 2012, Trento, Italy, September 17–19, 2012. Proceedings, volume 7477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33211-0 (print), 3-642-33212-9 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33212-8.

Xiao:2012:GIS

[952] Ningchuan Xiao, Mei-Po Kwan, Michael F. Goodchild, and Shashi Shekhar, editors. Geographic Information Science: 7th International Conference, GIScience 2012, Columbus, OH, USA, September 18–21, 2012. Proceedings, volume 7478 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33023-1 (print), 3-642-33024-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33024-7.

Szabo:2012:ICT

[953] Róbert Szabó and Attila Vidács, editors. Information and Communication Technologies: 18th EUNICE/IFIP WG 6.2, 6.6 International Conference, EUNICE 2012, Budapest, Hungary, August 29–31, 2012. Proceedings, volume 7479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32807-5 (print), 3-642-32808-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32808-4.

Barros:2012:BPM

[954] Alistair Barros, Avigdor Gal, and Ekkart Kindler, editors. Business Process Management: 10th International Conference, BPM 2012, Tallinn, Estonia, September 3–6, 2012. Proceedings, volume 7481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32884-9 (print), 3-642-32885-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32885-5.

Jonker:2012:SDM

[955] Willem Jonker and Milan Petković, editors. Secure Data Management: 9th VLDB Workshop, SDM 2012, Istanbul, Turkey, August 27, 2012. Proceedings, volume 7482 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32872-5 (print), 3-642-32873-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-32873-2.

Gollmann:2012:ISI

[956] Dieter Gollmann and Felix C. Freiling, editors. Information Security: 15th International Conference, ISC 2012, Passau, Germany, September 19–21, 2012. Proceedings, volume 7483 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33382-6 (print), 3-642-33383-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33383-5.

Kaklamanis:2012:EPP

[957] Christos Kaklamanis, Theodore Papatheodorou, and Paul G. Spirakis, editors. Euro-Par 2012 Parallel Processing: 18th International Conference, Euro-Par 2012, Rhodes Island, Greece, August 27–31, 2012. Proceedings, volume 7484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32819-9 (print), 3-642-32820-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-32820-6.

Visconti:2012:SCN

[958] Ivan Visconti and Roberto De Prisco, editors. Security and Cryptography for Networks: 8th International Conference, SCN 2012, Amalfi, Italy, September 5–7, 2012. Proceedings, volume 7485 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32927-6 (print), 3-642-32928-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32928-9.

Fisher:2012:CLM

[959] Michael Fisher, Leon van der Torre, Mehdi Dastani, and Guido Governatori, editors. Computational Logic in Multi-Agent Systems: 13th International Workshop, CLIMA XIII, Montpellier, France, August 27–28, 2012. Proceedings, volume 7486 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32896-2 (print), 3-642-32897-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32897-8.

Eiter:2012:RWS

[960] Thomas Eiter and Thomas Krennwallner, editors. Reasoning Web. Semantic Technologies for Advanced Query Answering: 8th International Summer School 2012, Vienna, Austria, September 3–8, 2012. Proceedings, volume 7487 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33157-2 (print), 3-642-33158-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33158-9.

Catarci:2012:IAE

[961] Tiziana Catarci, Pamela Forner, Djoerd Hiemstra, Anselmo Peñas, and Giuseppe Santucci, editors. Information Access Evaluation. Multilinguality, Multimodality, and Visual Analytics: Third International Conference of the CLEF Initiative, CLEF 2012, Rome, Italy, September 17–20, 2012. Proceedings, volume 7488 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33246-3 (print), 3-642-33247-1 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33247-0.

Zaphiris:2012:TPD

[962] Panayiotis Zaphiris, George Buchanan, Edie Rasmussen, and Fernando Loizides, editors. Theory and Practice of Digital Libraries: Second International Conference, TPDL 2012, Paphos, Cyprus, September 23–27, 2012. Proceedings, volume 7489 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33289-7 (print), 3-642-33290-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33290-6.

Traff:2012:RAM

[963] Jesper Larsson Träff, Siegfried Benkner, and Jack J. Dongarra, editors. Recent Advances in the Message Passing Interface: 19th European MPI Users’ Group Meeting, EuroMPI 2012, Vienna, Austria, September 23– 26, 2012. Proceedings, volume 7490 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33517-9 (print), 3-642-33518-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33518-1.

Coello:2012:PPSa

[964] Carlos A. Coello Coello, Vincenzo Cutello, Kalyanmoy Deb, Stephanie Forrest, Giuseppe Nicosia, and Mario Pavone, editors. Parallel Problem Solving from Nature — PPSN XII: 12th International Conference, Taormina, Italy, September 1–5, 2012, Proceedings, Part I, volume 7491 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32936-5 (print), 3-642-32937-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32937-1.

Coello:2012:PPSb

[965] Carlos A. Coello Coello, Vincenzo Cutello, Kalyanmoy Deb, Stephanie Forrest, Giuseppe Nicosia, and Mario Pavone, editors. Parallel Problem Solving from Nature — PPSN XII: 12th International Conference, Taormina, Italy, September 1–5, 2012, Proceedings, Part II, volume 7492 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32963-2 (print), 3-642-32964-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32964-7.

Herskovic:2012:CTI

[966] Valeria Herskovic, H. Ulrich Hoppe, Marc Jansen, and Jürgen Ziegler, editors. Collaboration and Technology: 18th International Conference, CRIWG 2012 Raesfeld, Germany, September 16–19, 2012 Proceedings, volume 7493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33283-8 (print), 3-642-33284-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33284-5.

Barcelo:2012:DAI

[967] Pablo Barceló and Reinhard Pichler, editors. Datalog in Academia and Industry: Second International Workshop, Datalog 2.0, Vienna, Austria, September 11–13, 2012. Proceedings, volume 7494 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32924-1 (print), 3-642-32925-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32925-8.

Sirakoulis:2012:CAI

[968] Georgios Ch. Sirakoulis and Stefania Bandini, editors. Cellular Automata: 10th International Conference on Cellular Automata for Research and Industry, ACRI 2012, Santorini Island, Greece, September 24–27, 2012. Proceedings, volume 7495 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33349-4 (print), 3-642-33350-8 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33350-7.

Takagi:2012:PSI

[969] Tsuyoshi Takagi, Guilin Wang, Zhiguang Qin, Shaoquan Jiang, and Yong Yu, editors. Provable Security: 6th International Conference, ProvSec 2012, Chengdu, China, September 26–28, 2012. Proceedings, volume 7496 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33271-4 (print), 3-642-33272-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33272-2.

Krotzsch:2012:WRR

[970] Markus Krötzsch and Umberto Straccia, editors. Web Reasoning and Rule Systems: 6th International Conference, RR 2012, Vienna, Austria, September 10–12, 2012. Proceedings, volume 7497 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33202-1 (print), 3-642-33203-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33203-6.

Gheyi:2012:FMF

[971] Rohit Gheyi and David Naumann, editors. Formal Methods: Foundations and Applications: 15th Brazilian Symposium, SBMF 2012, Natal, Brazil, September 23–28, 2012. Proceedings, volume 7498 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33295-1 (print), 3-642-33296-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33296-8.

Sojka:2012:TSD

[972] Petr Sojka, Aleš Horák, Ivan Kopeček, and Karel Pala, editors. Text, Speech and Dialogue: 15th International Conference, TSD 2012, Brno, Czech Republic, September 3–7, 2012. Proceedings, volume 7499 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32789-3 (print), 3-642-32790-7 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32790-2.